### v1

# Hyperparameter Optimization in Machine Learning

Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determine the effectiveness of systems based on these technologies. Manual hyperparameter search is often unsatisfactory and becomes unfeasible when the number of hyperparameters is large. Automating the search is an important step towards automating machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples and insights into the state-of-the-art. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model- and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
[…]
Chapter 1 Introduction
Over the last decade, machine learning has established itself as a key driver of innovation in science and industry. The digitalization of human interactions has led to the generation of data at an unprecedented scale. These data could be stored ever more cheaply thanks to advances in micro-electronics, as well as the advent of cloud computing. Progress in hardware and low-power chip design also led to an exponential increase of computing capabilities of respectively cloud infrastructure, such as graphical processing units (GPUs), and devices on the edge, such as mobile phones. These two trends, cheaper and better hardware for storage and compute, lay the foundation for the successful application of statistical machine learning in practice and its adoption in artificial intelligence. It enabled learning algorithms to sift through vast amounts of data to discover and extract patterns that can then be used to make data-driven decisions without or with limited human intervention.
Perhaps the most notable illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed by machine learning experts at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of configuration parameters, called hyperparameters, were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization, an algorithm that sequentially predicts and then evaluates the performance of algorithms such as AlphaGo when it uses specific hyperparameters. Until a certain cost or time budget is exhausted, Bayesian optimization repeatedly computes an exploration-exploitation trade-off to determine what is the next most interesting set of hyperparameters to evaluate based on its performance observed so far.
In practice, the generalization capabilities of any machine learning algorithm also depends on hyperparameters. Whereas conventional machine learning algorithms, such as support vector machines [shawe-taylor:2004:kernel_methods_pattern_analysis], need hand-engineered features to transform the raw data in a suitable format, deep neural networks learn in an “end-to-end” fashion by ingesting the raw data directly [lecun2015nature]. However, in both cases, the composition of the model (e.g., the choice of feature extractors or the neural network architecture) and the optimization algorithm depend on quantities that will determine how effective the algorithm can learn a specific natural language understanding, computer vision or speech recognition task.
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as is commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression trained with stochastic gradient descent to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Table 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $\ell_{1}$ or $\ell_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Table 1.1(b). Interestingly, they showed that tuned logistic regression performed on par with a convolutional neural network using a bag of words representation of the text and only slightly worse that a sequential convolutional neural network. We will come back to this example in more details in the next chapter.
(a) Hyperparameters.
(b) Accuracy.
Figure 1.1: We reproduce the results obtained by yogotama2015-emnlp15. Table 1.1(a) are the hyperparameters that were optimized. The feature extraction depends on the range of $n$-grams considered through $n_{min}$ and $n_{max}$. Stop words could be removed or not and words weighted according different weighting schemes. Two choices of regularization were considered and the amount of regularization was optimized. The convergence tolerance was also optimized. The binary classifiers compared are either SVM (support vector machine), LR (logistic regression), NN (multi-layer perceptron) or CNN (convolutional neural network). The authors found that when optimizing all these hyperparameters, simple logistic regression trained by stochastic gradient descent was performing as well as convolutional neural networks, which were the state of the art at the time of publication.
Hence, hyperparameters play a critical role in machine learning as they do not only determine the generalization capabilities of trained models, but they may actually determine what constitutes the state-of-the-art. Indeed, results reported in an empirical study might look very differently for another choice of hyperparameters and so would be the conclusions drawn. Unfortunately, it is not uncommon that published results do not report the specific hyperparameters that were used to run experiments, for instance, to demonstrate that a proposed method is superior to previously published ones, and are thus not reproducible, putting in question what actually is the state-of-the-art in machine learning today [haibe-kains:2020:transparency].
The subject of this monograph is hyperparameter optimization (HPO). The availability of algorithmic tools to automate this task and their adoption will help machine learning researchers dealing with hyperparameters in a principled way. More importantly, these tools will guarantee that new published advances are reproducible and thus unequivocally contribute to the field. Along with the experimental set-up, we argue that researchers would need to publish the HPO algorithm they used, the amount of resources it was allowed to consume [dodge:2019:show_your_work], and the hyperparameter search space that was considered [gundersen:2018:state_art]. Our objective is to provide a comprehensive overview of established techniques and recent advances in HPO. By doing so we hope we will help establish the practice of systematically using automatic HPO algorithms in research and industry.
1.1 What is hyperparameter optimization?
Let $A(\lambda)$ denote a learning algorithm configured by the hyperparameter vector $\lambda$ and consider a function $\mathscr{M}$ that evaluates its performance on a learning task. We call the composition $f(\lambda)=\mathscr{M}\circ A(\lambda)$ the response function. HPO consists in optimizing $f$ with respect to $\lambda$.
What distinguishes the HPO problem from conventional optimization problems arising for instance in convex or combinatorial optimization, is its nested nature: to evaluate the response function $f$ at a specific point $\lambda$, it is required to run the learning algorithm $A$. In practice, running $A$ can be time-consuming. For example, training ResNet-50 from scratch on CIFAR-10 on an Amazon Elastic Cloud Compute (EC2) g4dn.xlarge instance takes more than 2 hours. [^1] A ResNet is a deep residual neural network commonly used in computer vision. ResNet-50 has 50 convolutional layers with skip connections (one for every stack of three layers). We refer the interested reader to the original paper by he2016deep.
1.2 Why is hyperparameter optimization difficult?
In general, computing $f$ might not only be time-consuming, but also expensive in terms of amount of resources used or cost. Moreover, $f$ is required to satisfy minimal regularity constraints in practice. Its analytical form might not be known, it might have multiple local optima, it might exhibit wide flat regions, or it might be even non-smooth or non-continuous as a function of $\lambda$. We only require that $f$ is observable. Returning to the example of AlphaGo, this means that it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) and the probability of winning a game is known; what matters is that, for our specific choice of $\lambda$, we can observe if AlphaGo won or lost the game. Therefore, in general, one cannot apply conventional gradient-based optimization algorithms to solve HPO problems.
In addition, evaluations of $f$ can be further corrupted by noise, and the search space for $\lambda$ can itself be quite complex. Some hyperparameters could be continuous (e.g., learning rate), but others could be integral (e.g., number of layers in a deep neural network) or categorical to encode choices. Even more structure can emerge in the search space: conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units of the $l$-th layer of a neural network is only defined if the network has at least $l$ layers. If the domain to search over would be a small discrete set, one could carry out an exhaustive search by evaluating all the possible hyperparameter configurations in this set. Unfortunately, this is rarely the case in reality as the number of hyperparameter configurations to consider can be large or even uncountable.
All these challenges mean that, in practice, the focus of HPO is more often about finding “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than finding (global) optima. Hence, HPO often consists in determining a small set of good hyperparameters where to evaluate $f$.
1.3 Historical remarks
Modern machine learning often considers a hierarchy of model families, indexed by model parameters at the lower level and hyperparameters at the higher level. Traditional model selection, such as Akaike’s Information Criterion (AIC) and the Bayesian Information Criterion (BIC), focuses on choosing the “model size” (i.e., a single integer-valued hyperparameter that determines the number of model parameters) that results in a good generalization [akaike:1974:new_look_statistical, schwarz1978estimating]. Interestingly, the origins of some recent HPO methods can be traced back to at least two decades before the introduction of AIC and BIC: the HPO problem in machine learning can be seen as a special case of optimal experimental design, an area that has a long tradition in statistics since the introduction of the response surface methodology [box:1951:experimental_attainment_optimum]; see myers16:_respon for a recent account.
Many ideas useful for tuning hyperparameters originated from a different class of problems: the configuration of heuristics and meta-heuristics in problem solving in Computer Science. The goal in these cases is not generalization but finding a solution with a small amount of computational resources. For example the LEX system [mitchell:1983:learning_experimentation_acquiringa] already embodied some fundamental ideas behind model-based approaches to HPO. LEX ran the to-be-configured algorithm to produce examples from which a generalizer learned to produce better heuristics. Model-based approaches have been subsequently applied to global optimization and propositional satisfiability [hutter:2009:experimental_investigation_modelbased], and to supervised learning [thornton2013auto]. These model-based approaches are based on or closely related to Bayesian optimization [mockus:1975:bayesian_methods_seeking, sacks:1989:design_analysis_computer, jones1998efficient]; see also the review by shahriari:2016:taking_human_out. Bayesian optimization is a technique at the core of current model-based HPO approaches in machine learning (see Section 4.1). Yet another class of approaches is based on the idea of racing, where candidate configurations are run in parallel and, as soon as sufficient statistical evidence is collected against a candidate, the run is stopped and the candidate replaced by a new one. This idea was initially formulated for model selection in machine learning [maron:1994:hoeffding_races_accelerating], and then applied to configuration of meta-heuristics [birattari:2002:racing_algorithm_configuring]. Finally, population-based algorithms also originated in the area of problem solving with seminal works of rechenberg65 and fogel:1965:intelligent_decisionmaking_simulation, boosted by the introduction of genetic inspired operators by holland75. While their application to optimization in machine learning has a long tradition, the use of population-based algorithms in HPO is more recent and preceded by the idea of gradient-based approaches [bengio_gradient-based_2000].
1.4 Outline of the monograph
In this chapter, we gave a broad overview of the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function $f(\lambda)$ based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require a much smaller number of hyperparameters to be evaluated to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, which are genetic algorithms designed to solve global optimization problems. When applied to HPO, techniques in this family maintain a set of predictors whose hyperparameters can be updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how HPO is amenable to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency).
1.5 Notation
As usual, we denote by $\mathbb{R}$ and $\mathbb{N}$ the set of real and natural numbers, respectively. For $K\in\mathbb{N}^{+}$, we indicate by $[K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We denote by $\mathcal{N}(\mu,\Sigma)$ the Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
Chapter 2 Hyperparameter Optimization
In this chapter we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We will make use of examples of HPO problems to fix ideas and showcase different application scenarios. We conclude the chapter with an overview of the practical desiderata for HPO that we will return to when discussing the various techniques covered in this monograph.
2.1 Learning algorithms with hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm relies on in order to generalize outside of the training data. We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h$, capable of accomplishing a given task:
$A:\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda),$ (2.1)
where:
  • $\mathscr{D}$ is the collection of all finite datasets that the algorithm can process;
  • $\Lambda$ is the hyperparameter space, i.e., a subset of the Cartesian product of the domains of individual hyperparameters;
  • $\mathcal{H}$ is the model space, typically indexed by a parameter vector $w$, that represents the class of models that $A$ may output.
We call $\lambda\in\Lambda$ the hyperparameter vector or configuration. Note that in Eq. (2.1) we emphasize that the model returned by the algorithm depends on the hyperparameter vector $\lambda$ by writing $h_{\lambda}=A(\mathcal{D},\lambda)$. Likewise, in the following, we may omit that $A$ depends on the training data $\mathcal{D}$ when the context is clear. Indeed, typically, $\mathcal{D}$ is considered fixed in HPO.
2.1.1 Logistic regression for sentiment analysis
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Let us denote the data by $\mathscr{D}:=\bigcup_{N\in\mathbb{N}^{+}}(\mathcal{X}\times\mathcal{Y})^{N}$, where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda_{1}\times\dots\times\Lambda_{7}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}:\mathcal{X}\to\mathbb{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
=$\mathcal{H}=\left\{h_{\lambda}:\mathcal{X}\to\mathcal{Y}\,:\,h_{\lambda}(x)=\textbf{1}\{w^{\intercal}\Phi_{\lambda}(x)\geq 0\},\;w\in\mathbb{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
where $\mathbf{1}$ is the indicator function. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ vary in their respective spaces $\Lambda_{1}$, $\Lambda_{2}$, $\Lambda_{3}$ and $\Lambda_{4}$.
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the n-grams considered (controlled by the hyperparameters $\lambda_{1}$ and $\lambda_{2}$), the more input features and, consequently, parameters the model will have. Unigrams alone ($\lambda_{1}=\lambda_{2}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer n-grams typically requires increasingly larger corpora to avoid overfitting as the number of possible features (sequences of words of length n) grows exponentially with n.
The remaining hyperparameters, $\lambda_{5},\lambda_{6}$ and $\lambda_{7}$, do not directly affect the model space, but rather determine the way in which the optimization is carried out. Specifically, the algorithm $A$ outputs the model whose parameter vector $w$ minimizes the training objective $\mathcal{J}_{\lambda}:\mathbb{R}^{n_{\lambda}}\times\mathscr{D}\to\mathbb{R}$ defined as follows:
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log P\left(y|\Phi_{\lambda}(x);w\right)+\lambda_{6}\Omega_{\lambda_{5}}(w),$ (2.3)
where $\Omega_{\lambda_{5}}(w)$ is either the $L^{1}$ or the $L^{2}$ norm depending on the value of $\lambda_{5}$, and the probability that the model assigns a document to express a positive sentiment, encoded by $y=1$, is given by
$P\left(1|\Phi_{\lambda}(x);w\right)=\frac{1}{1+e^{-w^{\intercal}\Phi_{\lambda}(x)}}.$
Hyperparameter $\lambda_{5}$ exposes a binary choices between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values (see Figure 2.1 for a geometric interpretation). The hyperparameter $\lambda_{6}$ controls the strength of such regularization; if, for instance, $\lambda_{5}=L^{1}$, one can expect the number of 00’s in the solution vector to be proportional to the value of $\lambda_{6}$.
Figure 2.1: Geometrical representations of the $L^{2}$ (left) and $L^{1}$ (right) regularizers. $w^{*}_{0}$ represents the minimizer of the training loss $\mathcal{J}_{\lambda}$ when the regularization coefficient $\lambda_{6}$ is 00 (i.e. no regularization). $w^{*}_{\rho_{1}}$ and $w^{*}_{\rho_{2}}$ are, instead, the minimizers of the regularized losses for two different values of $\lambda_{6}=\rho_{i}$ for $i\in\{1,2\}$, with $0<\rho_{1}<\rho_{2}$. These corresponds to the projections of $w^{*}_{0}$ onto $L^{2}$ and $L^{1}$ balls with varying radii that are inversely proportional to $\lambda_{6}$.
Finally, $\lambda_{7}$ controls the convergence tolerance of the numerical solver used to minimize the objective (2.3).
Although logistic regression is a relatively simple algorithm, its practical application already exposes several hyperparameters. Other algorithms may require many more hyperparameters. This happens for example in deep learning, where feature extraction is internalized in the underlying models, and the optimization problems become more challenging.
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda_{6}$ in our example could in principle take any non-negative real value and it does not need to be limited to the interval $[10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in (2.3) instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
Hence, by exposing a (relatively) larger number of hyperparameters of the learning algorithm, such as parameters associated with the feature extractors or parameters associated with the optimizer or the regularizer, one can substantially expand the algorithm’s model space $\mathcal{H}$. These additional degrees of freedom may in turn raise the odds of finding a better model for the task at hand, at the price of having to search over a possibly much larger hyperparameter space.
2.2 The hyperparameter optimization problem
Hyperparameter configurations need to be evaluated according to a given performance metric, that we denote by $\mathscr{M}$. This metric takes as input a trained model and a validation set, and returns a scalar value, such as the validation error.  [^1] In Section 8.1 we discuss the case where $\mathscr{M}$ returns multiple scalars. The objective of an HPO problem, the response function, is then constructed from the performance metric. Specifically, given a learning algorithm $A:\mathscr{D}\times\Lambda\to\mathcal{H}$, a metric :×$\mathscr{M}:\mathcal{H}\times\mathscr{D}\to\mathds{R}$, a training set $\mathcal{D}$ and a validation set $\mathcal{V}$, the hyperparameter optimization problem is defined as follows:
$\min_{\lambda\in\Lambda}f(\lambda)$ $=$ $\mathscr{M}\left(h_{\lambda},\mathcal{V}\right)$ (2.4)
such $\mbox{such that}\quad h_{\lambda}$ $=$ $A(\mathcal{D},\lambda).$ (2.5)
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until convergence. Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of continuous algorithms out-of-the-box for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 when we will discuss gradient-based HPO techniques that explicitly leverage this structure. Here, we note only that we cannot collapse the optimization problem into one level (i.e., joint minimization of both model parameters and hyperparameters). For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda_{6}=10^{-5}$ and $\lambda_{7}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.
2.2.1 Estimating generalization
Most of the material and techniques for HPO we covered in this monograph may be used regardless of the specific nature and meaning of $\mathscr{M}$ and, hence, of $f$. However, since many hyperparameters control the inductive bias of their respective learning algorithms, the archetypal setting consists in a response function that estimates the generalization of the models outputby the learning algorithm.
Figure 2.2: Three examples of response functions. From left to right the underlying algorithms are: ridge regression, two-layers neural network regression, Monte Carlo tree search [chen_learning_2017].
Informally, a hypothesis generalizes if it “behaves sensibly” on data outside $\mathcal{D}$.
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g. in generative modelling), but a typical setting is what follows. We assume to have access to a task loss :×$\mathcal{L}:\mathcal{H}\times\mathcal{Z}\to\mathds{R}$, where $\mathcal{Z}$ is an appropriate task-dependent data space and define the generalization error of $h_{\lambda}$ as the expected loss:
$\mathcal{E}(h_{\lambda})\doteq\mathbb{E}_{z\sim p(z)}\left[\mathcal{L}(h_{\lambda},z)\right],$ (2.6)
where $p$ is the unknown task distribution from which the training dataset $\mathcal{D}$ was drawn independently. We may then define the validation error as an empirical estimate of (2.6):
$\mathscr{M}(h_{\lambda},\mathcal{V})=\frac{1}{|\mathcal{V}|}\sum_{z\in\mathcal{V}}\mathcal{L}(h_{\lambda},z),$ (2.7)
where $\mathcal{V}=\{z_{j}\}_{j=1}^{N^{\prime}}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. In the sentiment analysis problem, an instance of supervised binary classification, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. yogotama2015-emnlp15 use the 00-$1$ error (accuracy), namely $\mathcal{L}(h_{\lambda},(x,y))=\mathbf{1}\{h_{\lambda}(x)\neq y\}$.
Figure 2.2 shows some examples of other response functions. The leftmost plot concerns a simple ridge regression algorithm to regress median housing prices using the Boston dataset [harrison1978hedonic]. The scalar hyperparameter controls the $L^{2}$ regularization and $\mathcal{L}(h_{\lambda},(x,y))=\|h_{\lambda}(x)-y\|^{2}$. The plot shows the mean loss computed on the training and validation data, respectively. The center plot represent the response function of a two-layer neural network on the same learning scenario. Here the two hyperparameters are an $L^{2}$ regularization coefficient (real-valued, vertical axis) and the number of units in the hidden layer (integer, horizontal axis). The rightmost plot, reproduced from [chen:2018:bayesian_optimization_alphago] represents instead the winning probability of a Go agent introduced in Chapter 1, and somewhat deviates from the standard setting we just discussed.
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of (2.6); e.g., underestimation when the model is under-regularized as illustrated in Figure 2.2, left. This is because in order to obtain an unbiased estimate of the expected loss (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$.
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$. 222 Note that a third test set would be needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of (2.6) at the expense of increased computational cost.
2.2.2 Other types of response functions
One may also be interested in optimizing hyperparameters with respect to performance measures not necessarily linked to generalization. For instance, $\mathscr{M}$ could measure the runtime of $A$ when, e.g., tuning optimization hyperparameters [xu2008satzilla, hutter2009paramils], or the inference latency or the memory requirements of $h_{\lambda}$ when looking to deploy models for devices on the edge [benmeziane2021comprehensive]. chen:2018:bayesian_optimization_alphago take $\mathscr{M}$ to be an estimate of the winning rate of the AlphaGo agent that plays with hyperparameters $\lambda$ versus another AlphaGo agent that plays with a fixed baseline hyperparameter configuration $\lambda_{0}$. Figure 2.2 (right) plots the response function (i.e. the winning rate) along two hyperparameter axes.
Furthermore, one may be interested in optimizing the hyperparameters of a learning algorithm simultaneously with respect to multiple measures (e.g., accuracy, inference latency and memory footprint) or in ensuring that $h_{\lambda}$ satisfies some constraints regarding one or multiple measures. These scenarios give rise to multi-objective and constrained hyperparameter optimization, advanced topics which we will discuss in Chapter 8.
2.3 Hyperparameter optimization in the literature
Here below we report a few notable milestones and examples of “success stories” of HPO in the literature.
  • bergstra_random_2012, later extended in [bergstra_making_2013], brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, where widespread approaches to HPO. We cover grid, random and quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They improved the manual found hyperparameter by a human expert which marked state-of-the-art at this point by reducing the test error by $3\%$. This papers sparked interest in Bayesian optimization for HPO in the literature. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
  • As we mentioned in the introduction, chen:2018:bayesian_optimization_alphago used Gaussian-process-based Bayesian optimization to optimize the hyperparameter of the Monte-Carlo Tree-Search of Alpha GO. This improved the win-rate of Alpha GO from $50\%$ to $66.5\%$ in self-play games.
  • zoph2016neural parameterized the architecture of a convolutional and recurrent neural networks to discover new architectures. To sample new architectures they trained a recurrent neural network controller based on reinforcement learning. This work marked the beginning of a sub-field of research dubbed neural architecture search (NAS), which we discuss in Section 8.3.
  • jamieson2016non and later li2017hyperband introduced bandit-based approaches to hyperparameter optimization. These methods draw inspiration from the multi-armed bandit literature, particularly the successive halving algorithm [karnin-icml13]. They adaptively allocate resources to promising configurations while terminating unpromising ones early, emphasizing the concepts of budget and efficient compute allocation. These works have been particularly influential in the development of multi-fidelity optimization methods, which we explore in Chapter 5.
2.4 Practical desiderata for HPO algorithms
The nested optimization formulation (2.4) abstracts away some aspects which are practically relevant and, together, they set aside HPO from other optimization scenarios. For this reason one should weigh an HPO algorithm alongside multiple directions. We summarize these desiderata in Table 2.1.
Characteristic Brief Description
Sample efficiency How many evaluations of $f$ does the algorithm need to find satisfactory configurations?
Runtime efficiency How long does the algorithm take to find satisfactory configurations?
Parallelizability: How efficiently does the algorithm exploit parallel computational resources if:
   | Synchronous •  all parallel evaluations need to start at the same time?
   | Asynchronous •  evaluations can be started any time a resource becomes available?
Multi-Fidelity Support Can the algorithm take advantage of cheaper approximate evaluations of $f$?
Applicability Does the algorithm have restriction on the type of search space $\Lambda$?
Accessibility Which type of access does the method need to the underlying learning algorithm and/or training data?
Ease of use How difficult is to implement and tune the algorithm? What expertise is required to use it effectively?
Table 2.1: List of desirable properties of an HPO algorithm. [We will provide a comparison table of the various HPO techniques where these characteristics will appear as columns after the chapters reviewing the various methods. We will add a link to that table here]
Efficiency and budget.
The response function $f(\lambda)$ is typically noisy, depending on random effects in the computation of $A(\mathcal{D},\lambda)$ (e.g., weight initialization, data ordering, e.g., when performing stochastic gradient descent, etc.). It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ – expressed in time, number of function evaluations, monetary costs, etc. – rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a sub-optimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
Multi-fidelity.
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access $A_{b}(\mathcal{D},\lambda)$, $b<64$, for free. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5.
Synchronous and asynchronous parallelization.
We can trade off time efficiency against computation cost by running several evaluations of $f$ in parallel on (say) $W$ workers. For a synchronous system, all $W$ workers start new jobs at the same time. While this reduces wall-clock time, it often leads to idle time when workers need to wait for the slowest job to finish. In an asynchronous system, each worker can be assigned a new job independently from the status of the others. However, HPO decision-making for parallel execution systems comes with additional challenges. In the synchronous case, a batch of $W$ new configurations has to be suggested in one go. Configurations in a batch should be “diverse”, in that they reveal non-redundant information about the optimum [Ginsbourger:10, Desautels:14, Gonzales:16, Wilson2018]. While only a single new configuration needs to be proposed in the asynchronous case, up to $W-1$ responses of running evaluations are not yet available [snoek_practical_2012]. Not taking such pending evaluations into account runs the risk of increased redundancy in the proposed configurations. HPO can also be distributed across several computational nodes that may reside in networked computers. In this case, algorithms should also be parsimonious in terms of communication to reduce overhead as much as possible.
Applicability and accessibility.
The applicability and accessibility of an HPO algorithm refer to the requirements it may have on the type of problem or the level of access needed to the underlying learning algorithm. Some algorithms may be restricted to certain types of search spaces, e.g., continuous vs integer or categorical domains, or a mix of these. Furthermore, some algorithm can exploit conditionality in the search space, while others cannot. Some HPO methods like gradient-based techniques (Chapter 7) presuppose a white-box access to the underlying learning algorithm in terms of computational structure and/or data processing, while others assume the learning algorithm to be a black box. In between, there are a class of algorithms that require a gray-box access, like the ability to start and stop the execution of the learning algorithm and be able to observe intermediate statistics. Although HPO has been traditionally described and tackled as a black-box optimization problem, more restrictive assumptions may be met in specific cases and HPO algorithm that exploit these assumptions typically show efficiency gains.
Ease of use
The ease of use desideratum refers to the complexity involved in implementing, configuring, and effectively using an HPO algorithm. Some methods may be conceptually simple but require careful tuning of their hyperparameters or auxiliary components. Others may be more complex to implement from scratch or require deep expertise in the underlying techniques (e.g., Bayesian and multi-fidelity optimization) to use them effectively.
2.5 Tackling the search manually
In the next chapters we will introduce the main algorithms for HPO. But in this section we take a few sentences to state the obvious. Searching for the best hyperparameters by hand can be alluring: getting started requires no new frameworks or assumptions, and one may be under the impression that the search could proceed organically, driven by hunches.
Yet manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Manual search also suffers from poor reproducibility. The winding path through the hyperparameter space exists only in the searcher’s mind, impossible to retrace.
Chapter 3 Elementary Algorithms
In this chapter, we present simple non-adaptive, model-free and “memoryless” approaches to HPO, in that subsequent runs are not informed by any previous search history. These methods correspond to basic search schemes in global optimization that are conceptually simple, widely applicable, (mostly) easy to implement and straightforward to parallelize. Furthermore, these methods may work as initialization schemes for more advanced techniques like Bayesian optimization, as we shall see in later chapters.
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favour of a fixed, simple distribution over the search space, drawing configurations from it until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offers benefits over random search, although to date we are not aware of any theoretical result supporting this.
3.1 Grid Search
Perhaps the first formal description of grid search (GS) dates back to the 1930’s, appearing in a foundational work by fisher1936design on experimental design, under the name of factorial design. GS consists in evaluating $f(\lambda)$ over a predefined grid of hyperparameter configurations $\upsilon\subset\Lambda$. For each entry of $\lambda\in\Lambda=\Lambda_{1}\times\dots\Lambda_{m}$, the user defines a list of values:
$\bar{\lambda}^{i}=\{\bar{\lambda}^{i}_{j}\}_{j=1}^{n_{i}},\quad\bar{\lambda}^{i}_{j}\in\Lambda_{i},\quad\text{for }\;i\in[m].$
The search grid is then given by all the combinations on the Cartesian product generated by these $m$ sets, namely $|\upsilon|=\prod_{i=1}^{m}n_{i}$. The learning algorithm can be executed sequentially or in parallel on the resulting set of points. The hyperparameter configuration that minimizes the objective function $f$ over $\upsilon$ is returned as $\widehat{\lambda}$; that is
$\widehat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\upsilon}f(\lambda).$ (3.1)
GS is a simple memoryless method that can be used to explore low-dimensional hyperparameter spaces. It requires very little implementation effort and parallelization is straightforward. The number of total trial generated by GS, however, grows linearly in the size of the trial sets $\bar{\lambda}^{i}$ and is exponential in $m$, making the method impractical for even moderate values of $m$. In addition, while Boolean and categorical hyperparameters naturally lend themselves to an evaluation on a predefined grid, unbounded integer and real-valued hyperparameters require manually setting bounds and discretization.
Extensive experimental practice has shown that some continuous hyperparameters, such as the learning rate, benefit from discretization on a logarithmic scale [bengio2012practical]. While requiring less expert knowledge than manual search, GS is very sensitive to the definition of the grid. To reduce the computational overhead, GS can be paired with manual search: the user alternates GS steps with adjustments and refinement of the grid values, e.g., using a finer grid “centered” around the best found configuration, or moving the search bounds of continuous hyperparameters, until a satisfactory result is found.
3.2 Random Search
Perhaps conceptually the simplest procedure for HPO is random search (RS), which was already considered in the 50’ as a method for global optimization [brooks1958discussion]. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$ and it is readily parallelized. The incumbent solution after $K$ trials is given by
$\widehat{\lambda}_{K}=\operatorname*{arg\,min}_{(\lambda_{1},\dots\lambda_{k})\sim\upsilon^{K}}f(\lambda_{i}),$ (3.2)
where the trial points $\lambda_{i}$ are drawn i.i.d. from $\upsilon.$ While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or, much more frequently in HPO, after a budget is exhausted (see Section 2.4), e.g., one may run RS for 12 hours.
Even if in principle $\upsilon$ could be any distribution, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions, such as uniform or log-uniform, that factorize independently along the configuration components, i.e. $\upsilon=\prod_{j=1}^{m}\upsilon_{i}$. For instance, if $\lambda$ contains, in order, a categorical choice between $K$ different neural architectures, the learning rate for stochastic gradient descent and a global dropout factor for regularization then $\upsilon$ could be set as
$\upsilon=\mathcal{U}\{1,K\}\times\mathrm{Log}\,\mathcal{U}(10^{-5},10^{-1})\times\mathcal{U}(0.5,1).$
where $\mathcal{U}\{\cdot\}$ and $\mathcal{U}(\cdot)$ are the discrete and continuous uniform distributions and $\mathrm{Log}\,\mathcal{U}(\cdot)$ is the log-uniform (or reciprocal) distribution on the indicated intervals.
On the efficiency of random search (compared to grid search).
In an influential article, bergstra_random_2012 advocate for using RS (instead of manual search or grid search) as the “default” baseline method for HPO. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter it could be that $f(\lambda_{1},\lambda_{2})\approx g(\lambda_{1})$.
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to a positive outcome of the experiment, i.e., where $f$ is sufficiently small, and compare the volume (under random sampling from $\upsilon$) of such region with that of the full search space [brooks1958discussion]. We can model the probability of success of random search after $K$ draws as follows:
$\mathbb{P}_{\upsilon}\left(\widehat{\lambda}_{K}\in\Lambda^{*}\right)=1-\left(1-\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})\right)^{K},$ (3.3)
where $\mathbb{P}_{\upsilon}$ denotes the probability of sampling from $\upsilon$, typically uniform or log-uniform. Eq. (3.3) says that the minimum number of trials needed to assure a positive outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda_{*}$ under $\upsilon$. This can be indeed independent of the search space dimension if the response function has a low effective dimensionality, argument that may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures, where the probability of hitting a favorable configuration by random sampling is relatively high.
However, for a general search space, we should expect $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider, for instance, the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m}$ is the $m$-dimensional $1/2$ hypercube. If $\upsilon=\mathcal{U}(\Lambda)$ is the uniform distribution over $\Lambda$, then $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})=2^{-m}$. Consequently, the number of trials needed for having a positive outcome with at least $95\%$ probability is $K\geq\log(0.05)/\log(1-2^{m})$, which grows exponentially in $m$. For $m=3$ one already needs $K\geq 23$ draws to achieve a $95\%$ probability of success. On the other hand, in a low effective dimensionality setting we could have that $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m^{\prime\prime}}$ with $m^{\prime}+m^{\prime\prime}=m$. It is immediate to see that $K$ grows with $m^{\prime}$ only, illustrating the fact that RS is insensitive to to the number of inactive dimensions, covering a larger number of distinct values compared to GS.
An additional advantage of random search over grid search is that RS allows for a more natural treatment of a larger class of hyperparameters, not requiring discretization of continuous search spaces and allowing for distributions with non-compact supports. However, the randomized nature of this HPO technique means that “unlucky” runs may end up with entire regions of $\Lambda$ being under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed. Continuing with our illustrative examples, say we divide the unit hypercube in $2^{m}$ 1/2 hypercubes and we wish to have one trial per hypercube (anywhere). While GS would need exactly $2^{m}$ points to achieve this, the probability under uniform sampling that RS hits all the hypercubes in $2^{m}$ draws is $\prod_{j=1}^{2^{m}-1}(1-j/2^{m})$. For example, for $m=3$ such probability is already $3/32$, meaning that, in expectation, RS with $8$ draws covers all the hypercubes only once every ten runs.
3.3 Quasi-random search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{k\to\infty}\min_{i\in[k]}f(\lambda_{i})=f(\lambda^{*})$. RS with a uniform distribution is consistent. However, unlike GS, as we have discussed above, for a fixed budget of draws RS may generate a set of configurations that misses large portions of the search space. Intuitively, good sequences should cover evenly the search space. Also their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality; a property that GS is missing. Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks. The resulting method is conceptually very similar to both GS and RS, being highly parallelizable and easy-to-implement, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{k}$ are generated according to more complex procedures. We review a few possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. One can then derive samples for a variety of distributions with appropriate reparameterizations [devroye2006nonuniform].
Latin hypercube sampling (LHS) [mckay1979comparison].
Choose a sequence length $n>1$, which determines the number of hyperparameters generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}:[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
$\lambda_{k}\sim\mathcal{U}(H_{k}),\;\text{ where }\;H_{k}=I_{\sigma_{1}(k)}\times I_{\sigma_{2}(k)}\times\cdots\times I_{\sigma_{m}(k)},$ (3.4)
and $\mathcal{U}(H_{k})$ is the uniform distribution on $H_{k}$. This generates $n$ trial points; one may then generate further sequences of $n$ configurations by repeating the process. Latin hypercube sampling ensures that the projection of the sequence on each axis will place exactly one sample per interval.
Sobol’ sequences [sobol1979systematic].
Sobol’ sequences are a collection of separate one-dimensional sequences, one for each axis: when constructing a $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$ (the iterate) let $k=\sum_{l}a_{l}(k)2^{j}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 00’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
$[\lambda_{k}]_{j}=\sum_{l}\frac{[C_{j}\cdot a(k-1)]_{l}}{2^{l+1}}$
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., addition is modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to [Lemieux:1338331, Chapter 5] for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences by adding random shifts, i.e. $\lambda^{\prime}_{k}=\lambda_{k}+u$ (modulo the unit hypercube) with $u\sim\mathcal{U}([0,1]$ and by “scrambling” the binary representation vector $a(k)$, that is randomly permuting its entries [owen1995randomly].
Other schemes.
bousquet2017critical experimented also with Halton and Hammersley sequences, finding the latter (with scrambling and random shifting) to be overall robust and well performing on the tested deep learning scenarios. Other quasi-random low-discrepancy sequences are presented in [Lemieux:1338331]. LHS, Sobol’, Halton and Hammerley sampling are implemented in the stats.qmc module of SciPy.
Figure 3.1: Non-adaptive sampling methods on the unit square. The main plot shows a sequence of samples generated by different methods. For all methods except Latin Hypercube Sampling (LHS), the initial 4 samples are represented by darker and larger circles, followed by 5 additional samples depicted as lighter and smaller circles. For LHS, a sequence of 9 samples is directly shown, as this method requires resampling for each selected number of draws. The points on the top and right of the main plot are projections of the samples onto the respective axes. We can observe that the grid sampling method covers only two (with 4 draws) and three (with 9 draws) distinct points on the axes projections, indicating poor coverage. The first four draws of uniform random sampling illustrate the “unlucky run” phenomenon where the points concentrate in a small region of the square. Even after 9 draws, random sampling leaves the first quadrant unvisited. In contrast, Latin Hypercube Sampling and Sobol’ sequences (with random shifts and scrambling) provide more even coverage of the space.
Chapter 4 Model-based Methods
The techniques described in the previous chapter do not make any attempt to model the response function $f(\lambda)$. Except for manual search, the information retrieved on the previous runs of the learning algorithm does not influence the search. In contrast, model-based (MB) techniques construct an explicit surrogate model for $f(\lambda)$ – or for closely related quantities – based on the information collected from previous evaluations. We will denote such model by
$\hat{f}:\Lambda\times\mathcal{X}\to\mathcal{Y},$ (4.1)
where $\mathcal{X}$ may represent additional inputs (e.g., the performance of partial executions of $A$, a baseline or the value of the incumbent solution) and $\mathcal{Y}$ is an appropriate output space. For instance, $\mathcal{Y}$ could be (a distribution over) $\mathbb{R}^{+}$ if $\hat{f}$ models directly $f$, or the interval $[0,1]$ if $\hat{f}$ predicts the probability that a configuration improves over a given baseline.
The landscape of model-based methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we describe in more details in the next section. In BO the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation. Nonetheless, some recent works go in different directions. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of $f$, while hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer for performing HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, linked to gradient-based HPO may also be interpreted as an attempt to learn a surrogate model of the entire learning algorithm (applied to a specific dataset), whereby the output space of $\hat{f}$ coincides with the hypothesis space of $A$ itself, that is $\hat{\mathcal{Y}}=\mathcal{H}$ in (4.1).
Regardless of the particular implementation, one critical feature of surrogate models is that they should be much cheaper to evaluate than $f$ itself (we recall that evaluating $f$ requires executing the learning algorithm $A$, which may be very expensive). This typically allows model-based techniques to compute $\hat{f}$ on several values at each step, e.g., to seek for extrema. Another critical feature of surrogate models is the ability to provide uncertainty estimations for evaluations of $f$. This allow the optimization process to balance exploration, i.e., collecting information in areas of high uncertainty to improve the model, and exploitation, i.e., leveraging the learned model to sample hyperparameters close to the current minimizer.
The vast majority of model-based methods is, in principle, sequential: the surrogate model suggests the next point to evaluate while being simultaneously adapted online as new data from the execution of $A$ is collected and processed. We can model such iterative process introducing a two mappings $\pi_{1}$ and $\pi_{2}$, one corresponding to the iteration in the hyperparameter space, the other to the update of the surrogate model:
$\left(\begin{array}[]{c}\lambda_{k+1}\\ \hat{f}_{k+1}\end{array}\right)=\left(\begin{array}[]{c}\pi_{1}(\hat{f}_{k},\upsilon)\\ \pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\hat{f}_{k},\upsilon\right)\end{array}\right).$ (4.2)
Another possibility is to first learn $\hat{f}$ and then use it to suggest one or several hyperparameter configurations in a second stage, following a parallel scheme. Examples of this latter approach include [feurer2015initializing, fusi2018probabilistic]. The methods, however, assume the presence of multiple tasks and are closer in nature to meta or transfer learning settings. Still, even in the sequential approach, typically $\hat{f}_{0}$ is initialized by using several evaluations obtained in a “warm-up stage”, e.g., by random search or Latin hypercube sampling. While the foremost representatives of these type of methods – tied to BO – are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in Chapter 5.
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an acquisition function depending on the surrogate model, as we will see in Section 4.2. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of MB techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
4.1 Bayesian Optimization
BO has long been used as a method to globally optimize black-box functions (i.e., functions for which we only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda)$, and dates back at least to the seventies with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. While limited to small-to-medium dimensional settings, BO is an effective approach to optimize (highly) multi-modal, non-convex and possibly expensive-to-compute functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give a brief overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for an intuitive introduction of the mathematical aspects of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning.
BO is a model-based technique, whose surrogate model is defined by a stochastic process, representing a probability distribution over the objective $f$. This surrogate model is fitted to available observations by Bayesian conditioning, so that afterwards it represents the Bayesian posterior distribution over function values $f(\lambda)$ away from the observations. This posterior distribution is used to suggest the next value to probe. Typically, the suggestion rule is expressed as the maximization of an acquisition function, balancing exploration – collecting information to improve the model itself, by proposing points away from previous observations – and exploitation – sampling close to the current minimizer in order to refine the best solution. Referring to the notation of the previous section and to Equation (4.2), the first component $\pi_{1}$ constitutes the decision rule based on an acquisition function, while the second component $\pi_{2}$ maps to the Bayesian update of our surrogate model given the data observed so far. Next, we describe a “classical” instantiation of BO, whereby the surrogate model is a Gaussian process [williams2006gaussian] and the acquisition function is the expected improvement [mockus1978application]. This approach is well-suited for the global optimization of smooth multivariate real functions, and constitutes the core of the Spearmint HPO package [snoek_practical_2012].
4.1.1 Bayesian linear model
We start by introducing a simple linear model $h_{w}$ in a given feature space $\mathcal{V}$, to which we add independent Gaussian noise with variance $\sigma^{2}$:
$h_{w}(\lambda)=\varphi(\lambda)^{\intercal}w+\varepsilon;\qquad\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}),$ (4.3)
where $\varphi:\Lambda\to\mathcal{V}$ is a feature map. While using linear maps directly in the input space would result in an overly simplistic model – we are trying to model an objective, the validation error, which can be (highly) non-convex and multimodal – working in a different feature space allows us to capture much richer behaviors, but still retain most of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in (4.3) models potential errors in the observations. In HPO, this noise arises because the validation error is only an estimate of the true generalization error. Additionally, this noise may partially account for the stochastic nature of the learning algorithm itself. It is important to distinguish between aleatoric uncertainty, which derives from inherent randomness in the data, and epistemic uncertainty, which arises from lack of knowledge about the model. A heteroscedastic approach can help disentangle these types of uncertainties, as proposed by [griffiths:2021] in their work on achieving robustness to aleatoric uncertainty with heteroscedastic Bayesian optimization.
Suppose now we are at the $k$-th iteration of the algorithm, after having already collected $k$ observations. Let $Y_{k}=\left(f(\lambda_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ be the vector of observed values, and denote by $X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)$ the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
$y_{*}\sim\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k}),$
where $x_{*}=\varphi(\lambda_{*})$ is the query point in the feature space. We could follow an empirical risk minimization approach, obtaining the weights $w_{k}\in\mathbb{R}^{\mathrm{dim}(\mathcal{V})}$ of the model (4.3) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k+1}=h_{w_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w_{k})$, capable of capturing only our belief of the amount of noise present in the observations. Rather, we would like the stochasticity of $\hat{f}$ to (primarily) represent our uncertainty about the behavior of $f$ in unexplored regions.111 The following argument, mutated from jones1998efficient, may help in better grasping the difference between these two aspects. Consider the case where $f$ is a non-linear smooth deterministic function. We may then either omit the noise term $\varepsilon$ in (4.3), or leave it for capturing modeling errors. In the first scenario, the ERM approach would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously sampled. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k+1}$ would banally be inaccurate at $\lambda_{i}$ for $i\leq k$. Thus, informally speaking, both ERM modeling options would leave us with a surrogate model that is surely inaccurate for virtually all $f$’s. A way to escape this apparent modeling stalemate is to replace the independent Gaussian noise with one that depends on the evaluation point, i.e. $\varepsilon=\varepsilon(\lambda)$. Then, when fitting the model to the $k$ observations, one may have $\varepsilon_{k}(\lambda_{i})=0$ for $i\leq k$ while letting $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, by continuity ($\varepsilon_{k}$ is smooth and continuous as difference of $C^{1}$ functions), $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behavior. Note, however, that the noise term $\varepsilon$ in (4.3) remains independent and indeed represents, there, the measurement noise. To this aim, we take a Bayesian perspective on the fitting problem, and reason about all the linear models (in the feature space) that agree with the observations, weighted by their “realization probability”.
More formally, we marginalize over the weights $w$ and set
$\hat{f}_{k+1}(\lambda_{*})=\int p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)p_{w}(\cdot|X_{k},Y_{k})\,\mathrm{d}w.$ (4.4)
Given our modeling assumption (4.3) and the hypothesis that the noise is independent, the first distribution simply reads as
$p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)=p_{y_{*}}(\cdot|x_{*},w)=\mathcal{N}(h_{w}(\lambda_{*}),\sigma_{\varepsilon}^{2})=\mathcal{N}(x_{*}^{\intercal}w,\sigma_{\varepsilon}^{2}).$ (4.5)
By the Bayes’ rule, $p_{w}(\cdot|X_{k},Y_{k})$ is instead given by
$p_{w}(\cdot|X_{k},Y_{k})=\frac{p_{y}(\cdot|X_{k},w)p_{w}(\cdot|X_{k})}{p_{y}(\cdot|X_{k})}.$ (4.6)
Regarding the numerator of (4.6), we have that, for the same argument of (4.3),
$p_{y}(\cdot|X_{k},w)=\mathcal{N}(X_{k}^{\intercal}w,\sigma_{\varepsilon}^{2}I).$
This is the likelihood of the observations given the model. We must instead make a decision regarding the second term of the numerator, that is the distribution of the model parameters. We make the very natural assumption of independence from the design matrix and set
$p_{w}(\cdot|X_{k})=p_{w}=\mathcal{N}(0,\Sigma),$ (4.7)
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with (4.7), we can now also compute the denominator of (4.6) by marginalizing over $w$ [shahriari:2016:taking_human_out], obtaining
$p_{y}(\cdot|X_{k})=\int p_{y}(\cdot|X_{k},w)p_{w}\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I)$ (4.8)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the corresponding kernel
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\varphi(\lambda)^{\intercal}\Sigma\,\varphi(\lambda^{\prime})=\langle\varphi(\lambda),\varphi(\lambda^{\prime})\rangle_{\Sigma}$ (4.9)
and rewrite (4.8) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathbb{R}^{k\times k}$ as
$p_{y}(\cdot|X_{k})=\mathcal{N}(0,K+\sigma_{\varepsilon}^{2}I).$
We finally have all the terms to compute (4.4); which is, again, Gaussian [williams2006gaussian]:
$\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k})=\mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)k_{*})$ (4.10)
where $k_{*}=\{\kappa_{\Sigma}(\lambda_{*},\lambda_{i})\}_{i=1}^{k}\in\mathbb{R}^{k}$ is the vector of kernel evaluations between the query and the observations points. Equation (4.10), despite its apparent complexity, describes a rich posterior distribution over the values of $f$ that is analytically computable and whose design depends essentially only on the definition of the kernel map (4.9).
Examples of kernel functions include the $\mathcal{C}^{\infty}$ squared exponential
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\sigma_{0}e^{-r_{\Sigma}^{2}(\lambda,\lambda^{\prime})/2};\qquad r_{\Sigma}^{2}(\lambda,\lambda^{\prime})=\lambda^{\intercal}\Sigma\lambda^{\prime},$
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$ and the Matérn family, which contains kernels (defined as functions of $r_{\Sigma}$, above) with various degree of smoothness [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow to sidestep the definition of (possibly obscurer) feature maps and covariance matrices. The series of “free parameters”222 That is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel. that appears in (4.10) – usually called hyperparameters in the Bayesian literature (not to be confused with $\lambda$) – do not have to be fixed constants, but may be estimated from the observations [williams2006gaussian, shahriari:2016:taking_human_out] (automatic relevance determination). Thus, at this stage, the design choice left to the experimenter mainly boils down the choice of the kernel function.
4.1.2 Gaussian Processes
Instead of worrying about feature maps and weights, we can directly work with a kernel function, a notion which gives rise to Gaussian processes (GP) [williams2006gaussian, murphy2012machine]. In fact, the Bayesian linear model converges to a GP in the limit of infinitely many features, if the prior distribution on the weights is scaled appropriately [neal1996bayesian].
A Gaussian process is a Gaussian distribution over functions $g(\lambda)$, fully specified by a mean function $\mu(\lambda)$ and a covariance function $\kappa(\lambda,\lambda^{\prime})$, or kernel function. It has the property that for any finite set $\lambda_{1},\dots,\lambda_{n}\in\Lambda$, the joint distribution of the vector $[g(\lambda_{1}),\dots,g(\lambda_{n})]$ is multivariate Gaussian with mean $[\mu(\lambda_{i})]_{i=1,\dots,n}$ and covariance matrix $[\kappa(\lambda_{i},\lambda_{j})]_{i,j=1,\dots,n}$. Our linear model derivation above gives rise to a zero-mean Gaussian process (since the prior over weights (4.7) has mean zero) with covariance function $\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta$, where $\delta(\lambda,\lambda^{\prime})=1$ if $\lambda=\lambda^{\prime}$ and 00 otherwise. In short, we can write
$\hat{f}=\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta).$
The surrogate model update equation $\pi_{2}$ (see (4.2)) may then be written as
$\hat{f}_{k+1}=\pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\xi_{2}\right)=\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta\,|\,X_{k},Y_{k}),$ (4.11)
which expresses the update as the conditioning of the Gaussian process on the observations $X_{k}$, $Y_{k}$. While this posterior is in principle an infinite-dimensional object, we only ever need its marginal distributions over single points $\lambda_{*}$ (or over small sets, such as pairs) in order to compute acquisition functions and make sampling decisions. These marginals are Gaussian with mean and (co)variance given by simple linear algebra expressions (4.10). More details on acquisition functions and the decision step $\pi_{1}$ is provided in Section 4.2.
To summarize, Equations (4.12) and (4.11) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. The configuration parameters $\nu_{2}$ may contain choices regarding the mean and covariance functions of the GP, the method with which the GP hyperparameters are estimated, and the acquisition function to use. Apart from the simplicity of a Gaussian process setup, its uncertainty representation fares particularly well for small number of observations, and the posterior representation is updated efficiently and robustly once additional data becomes available.
4.1.3 Limitations
Gaussian processes also come with a number of drawbacks that may limit their applicability to real-world scenarios. Most importantly, computing the posterior for $k$ datapoints costs $\mathcal{O}(k^{3})$, as can be seen in (4.10). If we are able to sample $f$ a few thousand times or more, GP-based BO starts to become intractable. It is also difficult to design kernel functions for high-dimensional $\Lambda$, or if the search space contains many integer, categorical, or conditional hyperparameters. In such situations, BO with other surrogate models can perform better, as shown in a comparative study by eggensperger_towards_2013.
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions, and take on the scalability issues in snoek_scalable_2015 by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms and bergstra_making_2013 focus both on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators (see also Section 4.3). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next section) to improve performance in the first stage of HPO, when the surrogate model may still be rather unreliable.
4.2 Acquisition Functions
Suppose we fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $C=\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function (AF) $\alpha(\lambda|\hat{f}_{k})$. In other words, our suggestion mapping $\pi_{1}$ becomes
$\lambda_{k+1}=\pi_{1}(\hat{f}_{k},\nu_{2})=\arg\max_{\lambda\in\Lambda}\,\alpha(\lambda|{\hat{f}_{k}}).$ (4.12)
In this section, we restrict our attention to Gaussian process surrogate models, even though most AFs discussed here can be applied to other models just as well. As we will see, most common AFs depend on the posterior distribution of $\hat{f}_{k}$ only through its marginal mean and variance
$\mu_{k}(\sigma)=\mathbb{E}\left[\hat{f}_{k}(\lambda)\right],\quad\sigma_{k}(\lambda)^{2}=\mathbb{V}\left[\hat{f}_{k}(\lambda)\right],$
given by (4.10). Here, $\mathbb{E}$ and $\mathbb{V}$ denotes expectation and variance over the posterior distribution of $\hat{f}_{k}(\lambda)$ respectively.
A useful AF has to guide sampling so as to gain as much new information about high-performing hyperparameters as possible, given a limited sample $C$ obtained so far. In particular, it needs to balance the trade-off between exploration (i.e., sampling away from points in $C$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $C$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
$\alpha_{\text{exploit}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda).$
Maximizing $\alpha_{\text{exploit}}$ constitutes an extreme form of exploitation, which typically gets stuck in suboptimal local optima. On the other extreme, we could sample where we know least about $f$:
$\alpha_{\text{explore}}(\lambda|\hat{f}_{k})=\sigma_{k}(\lambda).$
Maximizing $\alpha_{\text{explore}}$ is entirely focused on exploration and may waste many evaluations in regions of poorly performing hyperparameters. However, a linear combination of the two,
$\alpha_{\text{LCB}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda)+\eta\sigma_{k}(\lambda),\quad\eta>0,$
is a competitive and frequently used AF, called lower-confidence bound (LCB) [Srinivas2009].
A classic acquisition function, introduced in the seminal work by mockus1978application and frequently used in HPO libraries [snoek_practical_2012] and services [AMT2020], is the expected improvement (EI), defined as
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})=\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
Here, $\beta\in\mathbb{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far, $\beta=f(\hat{\lambda}_{k})$. EI has a simple closed-form expression in terms of $\mu_{k}(\lambda)$, $\sigma_{k}(\lambda)$, and in contrast to LCB does not come with a free parameter to be set.
A variant of this acquisition function is the probability of improvement [Kushner1964], given by
$\alpha_{\text{PI}}(\lambda|\beta,\hat{f}_{k})=\Phi(\nu(\lambda)),\text{ with }\nu(\lambda)=\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}.$
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points333 Which costs $\mathcal{O}(b^{3})$ time and $\mathcal{O}(b^{2})$ memory, if $b$ is the size of the set. or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. While these have a sound theoretical foundation and have been shown to be more sample efficient, they typically come with higher computational cost and are more complex to implement than standard options such as EI or LCB. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out]. Techniques for maximizing acquisition functions include grid search, where the acquisition function is evaluated at a predefined set of points, and random search, which evaluates the acquisition function at randomly selected points. More sophisticated methods involve the use of gradient-based optimization, which leverages the gradient of the acquisition function to find local maxima more efficiently.
4.3 Tree Parzen Estimator
While Gaussian process based BO methods model the distribution $p(y|\lambda)$ of observations $y$ given configurations $\lambda$, the Tree Parzen Estimator (TPE) [bergstra2011algorithms, bergstra_making_2013] fits two conditional densities over $\Lambda$:
$l_{\beta}(\lambda)=p(\lambda|y<\beta),\quad g_{\beta}(\lambda)=p(\lambda|y\geq\beta).$
Here, the baseline $\beta$ is set to a lower $\gamma$-quantile of the observed targets $\{y_{i}\}$ (e.g., $\gamma=0.15$), giving rise to a $(\gamma,1-\gamma)$ split of the data, on which $l_{\beta}(\lambda)$, $g_{\beta}(\lambda)$ are fit respectively. While TPE uses a tree of Parzen estimators for conditional search spaces, later work made use of kernel density estimators [falkner2018bohb]. The density ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$ is a strictly increasing function of the expected improvement444 However, note that we cannot set $\beta$ to the minimum value observed so far, as in the GP-BO case, since this implies $\gamma=0$ and no data to fit $l_{\beta}(\lambda)$ with. AF with baseline $\beta$, and TPE maximizes this AF approximately by sampling a set of configurations from $l_{\beta}(\lambda)$ and suggesting the one with the largest ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$. TPE is conceptually simple, easy to run with parallel evaluations, and can naturally deal with conditional search spaces. Subsequent work has shown that density ratios can be estimated by way of probabilistic binary classification, which is simpler and better supported by ML methodology than density estimation [Tiao:21].
4.4 Decision Making Under Parallelism
In its original formulation BO is sequential, which is unsuitable for many real-world scenarios where parallel computation resources are readily available in order to increase wall-clock time efficiency. A number of extensions of BO to parallel evaluation settings have been proposed. Although more time efficient, these techniques tend to be less sample efficient than sequential BO. The extensions we discuss aim thus to find a diverse set of configurations to preserve sample efficiency as much as possible.
For simplicity, assume a fixed set of $W$ workers available for parallel evaluations. Recall from Section 2.4 that systems may either schedule evaluation jobs synchronously or asynchronously. In the synchronous case, we need to suggest batches of $W$ configurations to be evaluated next. A simple way to do this would be to pick the $W$ top-ranked configurations when optimizing our AF. However, this strategy risks suggesting very similar configurations, and their parallel evaluation may reveal little more information than any single one of them. A heuristic approach to suggest a more diverse batch is to run Thompson sampling $W$ times with independently drawn sample paths [Kandasamy:18], but this can be quite expensive. Another idea is to select batch members greedily one at a time. For each configuration placed in the batch, the AF is modified by a local penalty function, ensuring that subsequently selected points are at least some distance away [Gonzales:16].
Independent of how configurations are suggested, synchronous job scheduling tends to be substantially less efficient than asynchronous. Namely, the time to process a batch is determined by the slowest one, while up to $W-1$ workers are idle waiting. Such differences in training times are particularly pronounced when searching over parameters determining the model size. In an asynchronous system, whenever a worker becomes available, it can be assigned a new configuration. While BO needs to make single suggestions only (as opposed to batches), this has to be done in the presence of up to $W-1$ pending evaluations, i.e. jobs which have been started for previously suggested configurations, but did not return a metric value yet. Ignoring pending evaluations runs the same risk of selecting highly redundant configurations in subsequent rounds. Fortunately, dealing with pending evaluations is quite related to ensuring diversity in batch selection, and local penalty methods apply just as well [Gonzales:16, Alvi:19]. Simpler approaches start from the observation that in sequential BO, diversity is enforced by conditioning on observed points. The resulting very low posterior variance shrinks the AF to small values near such points. In the kriging believer heuristic [Ginsbourger:10, Desautels:14], we impute the response for each pending evaluation by its current posterior mean and condition on this pseudo-datapoint. A somewhat more principled approach is fantasizing [snoek_practical_2012, Wilson2018], where pending target values are integrated out by Monte Carlo: they are sampled several times from the joint posterior over pending configurations, and the AF is averaged over these samples.
More sophisticated strategies to exploit parallel compute resources, either with random search or BO, are described in the next chapter.
Chapter 5 Multi-fidelity Methods
Training and validating a machine learning algorithm every time we evaluate the objective function might make HPO overly expensive. This is particularly true for modern machine learning algorithms applied on large datasets, such as deep neural networks, which often can take hours or days to train. In this chapter, we will consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are much faster to evaluate than the full response function and often provide us with a sufficiently strong signal to guide the optimization process.
5.1 Multi-fidelity optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is fast turnaround time. To this end, practitioners will usually work on a subset of the training data in their first rough explorations. They will also not wait until training has converged, but rather check validation performance after a couple of iterations or epochs. Even though these early metrics are not entirely reliable, they are useful in order to home in on promising parts of the search space.
In multi-fidelity HPO [li2017hyperband, klein2017fast, swersky2014freeze], we assume that we have access to cheap-to-evaluate fidelities $f_{b}(\lambda)$ of the true objective function $f(\lambda)$, where the fidelity level is specified by the budget $b\in[b_{min},b_{max}]$. If we set the budget to the maximum budget $b=b_{max}$, we will recover the original objective function $f_{b_{max}}(\lambda)=f(\lambda)$, and we assume that the approximation of $f(\lambda)$ improves as $b$ increases. At the same time, we assume that the cost $c(\lambda,b)$ associated with evaluating $f_{b}(\lambda)$, for example the training time, increases with $b$ as well. While we can freely select $b$ during the optimization process, we are usually interested in the performance on the full budget $b_{max}$. More formally, we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$. In the next section, we will look at some commonly used fidelities that we can exploit to speed up the optimization process.
5.1.1 Iterations
For iterative machine learning algorithms, such as neural networks, we can interpret the number of iterations as budget and take the performance after each iteration as an approximation for the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. For neural networks we usually interpret epochs as iterations rather than single mini-batch updates, to reduce the overhead from validation error computations.
Figure 5.1 left shows the validation accuracy after each epoch of training for different hyperparameter configurations of a convolutional neural network. One can see that even after just a few epochs we are already able to visually distinguish between poor and well performing configurations. However, to identify the best configuration, we often have to evaluate the top performing configurations for the full number of epochs.
5.1.2 Training dataset subsets
Figure 5.1: Example learning curves of convolutional neural networks trained on CIFAR10.
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 right shows the validation error of a convolutional neural network trained on increasing subsets of the CIFAR10 dataset. Each color represents a different hyperparameter configuration. As one would expect, the validation error decays with the training dataset size. More interestingly, however, the ranking across hyperparameters remains relatively consistent across dataset subsets [bornschein-icml20]. This is related to recent work on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which show empirically that training loss of auto-regressive transformer [vaswani_attention_2017] models often follows a power law function with respect to the size of the training data.
In multi-fidelity optimization we can exploit this observation by interpreting the budget as the training dataset size [klein2017fast, swersky_multi-task_2013]. Thereby, to evaluate $f_{b}(\lambda)$, we only vary the training dataset size, but keep the validation set fixed, to avoid inducing additional noise. Most methods resample the training dataset uniformly at random for each evaluation of $f_{b}(\lambda)$, but also more sophisticated strategies, such as importance sampling, to select dataset subsets are possible [ariafar-aistats21]. Importantly, the class distribution needs to be preserved not to bias the performance estimation.
5.1.3 Cross-validation
Cross-validation (CV) is a popular technique to reduce the risk of overfitting on the validation dataset (see also Section 2.2.1). Rather than fixing a single training/validation fold, we use $K$ of them and average our target criterion over these. A main drawback of CV is its cost, which is $K$ times more expensive than to use a single fold. We can approximate it via multi-fidelity, by using the number of folds $b=1,\dots,K$ as budget.
Compared to the previous two fidelites, here we do not assume that the performance increases with the budget but rather that the variance estimate of the generalization performance decreases with the number of folds.
5.2 Methods based on random search
In this section, we will review algorithms which exploit multiple fidelities and suggest new configurations by random sampling. We will begin by clarifying the difference between synchronous and asynchronous decision making. Random-search based synchronous multi-fidelity methods have first been proposed in the multi-armed bandits (MAB) literature. In the following Section 5.3 we will concentrate on model-based multi-fidelity methods, where decision making is also driven by a surrogate model.
The methods discussed here assume that the evaluation costs $c(\lambda,b)$ for every fixed $\lambda$ are linear (or affine linear) in $b$.
5.2.1 Synchronous and asynchronous decision making
Multi-fidelity methods can be classified into synchronous and asynchronous. In the former, decisions about stopping or promoting trials are delayed until a certain batch of evaluations has finished. In successive halving or Hyperband, to be discussed shortly, all evaluations in a rung have to finish before a fraction of these trials are allowed to continue. In contrast, a decision whether to stop or continue a trial must be done instantaneously in asynchronous methods, based on the data available at that time. ASHA is an example for this type of scheduling.
Synchronous multi-fidelity methods are simpler to understand and analyze. On the other hand, they are harder to implement efficiently in an asynchronously parallel computation environment, and even sophisticated implementations tend to clock more idle time.
5.2.2 Successive halving
Successive Halving (SH) [karnin-icml13, jamieson2016non] is a MAB strategy that aims to identify the best hyperparameter configuration out of a finite set of configurations. Given the maximum and minimum budget $b_{max}$ and $b_{min}$ for evaluating a single hyperparameter configuration, we first define a set of rung levels
=$\mathcal{R}=\{b_{min}\eta^{k}\;|\;k=0,\dots,K\}.$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{K}$, where $K\in\mathbb{N}$. Here $\eta\geq 2$ defines the halving constant and determines how aggressively successive halving will prune configurations. We start by drawing $n$ hyperparameter configurations uniformly at random and evaluate them on the first rung level, i.e each configuration is evaluated with a budget of $b_{min}$. For convenience the number of initial configurations is often set to $n=\eta^{K}$ but this can be also arbitrary. Afterwards, successive halving promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance to the next higher rung level $b_{min}\eta$. These steps are iterated until we reach the highest rung with level $b_{max}$. See Figure 5.2 for an visualization.
While conceptually simple, successive halving showed strong performance in the past and even out-performed more sophisticated Bayesian optimization techniques that only operate on the full budget [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration on the first rung level should be representative of its performance on the highest rung level. Otherwise configurations that perform best on $b_{max}$ will be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $\frac{1}{eta}$ of the configurations to the next rung.
How is parallel computation exploited by SH? Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,k)$ jobs in parallel, where $k$ is the number of free slots remaining in the currently processed rung. Here, $W-k$ workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has a single slot only, so $W-1$ workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.
5.2.3 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while for too large $b_{min}$, our edge over standard random search may diminish. Hyperband [li2017hyperband] deals with this problem by running different instances $l=0,\dots,K$ of SH in sequence. The $l$-th instance (or bracket) has rung levels $\mathcal{R}_{l}=\{b_{min}\eta^{k}\;|\;k=l,\dots,K\}$, so $l=0$ corresponds to SH with $K+1$ rung levels, while $l=K$ is random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at modest extra expense.
What about parallel computation? Being based on SH, Hyperband inherits the issue of idle workers noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which (at the expense of some implementation complexity) eliminates idle time for asynchronous scheduling.
5.2.4 Asynchronous successive halving (ASHA)
As noted above, neither successive halving nor Hyperband fare well in the presence of parallel computation. Moreover, even if a configuration is clearly superior to others in all fidelities, it takes a long time until it is promoted to the maximum level $b_{max}$. In fact, before any configuration can be promoted towards the next level, all slots in the current rung have to be filled. These shortcomings are addressed in asynchronous successive halving (ASHA) [Li:19]. In this algorithm, rungs do not have a predefined size, but only grow over time. Also, promotion decisions are done less conservatively. For example, assume that $\eta=3$. Scanning rungs from highest to lowest, any not-yet-promoted configuration whose metric value is smaller than $2/3$ of other entries, is promoted to the next level. If no configuration can be promoted, a new one is sampled for the lowest level. In other words, configurations are promoted as early as possible: once a rung has 3 entries, one of them moves to the next level. In ASHA, any available compute resource can be used instantly, and workers are never idle. Moreover, any superior configuration is rapidly promoted towards $b_{max}$, without having to wait at a number of synchronization points. In practice, ASHA tends to outperform synchronous SH or Hyperband significantly [Klein:20].
5.3 Model-based methods
So far, the multi-fidelity methods we looked at sample hyperparameter configurations purely at random. As we have seen in Chapter 4, more sophisticated methods that use a probabilistic model to sample configurations usually require less function evaluations to approach the global optimum than random-search-based approaches. In this section, we will discuss multi-fidelity strategies that use one or multiple probabilistic models to guide the search.
5.3.1 Fabolas
Fabolas [klein2017fast] models the performance $f_{b}(\lambda)$ of a hyperparameter configuration $\lambda$ across different subsets $b\in[b_{min},b_{max}]$ of the training data. Compared to classical Bayesian optimization (see Chapter 4), in this case, we cannot simply optimize standard acquisition functions, such as expected improvement, because they tend to over-exploit in this setting and mostly evaluate configurations on the highest budget.
Instead, Fabolas is based on the entropy search method by Hennig2012 which models the probability
$p_{\min}(\lambda\mid\mathcal{D})=p(\lambda\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda))$
of $\lambda$ being the global optimum. Now, the idea of entropy search is to select in each iteration the configuration that provides the most information about $p_{min}$, which we quantify by computing the KL-Divergence $KL\big{[}p_{min}(\lambda\mid\mathcal{D})||\mathcal{U}(\lambda)\big{]}$ between $p_{min}$ and a uniform distribution $\mathcal{U(\lambda})$ which serves as a prior positing that each point $\lambda\in\Lambda$ is equally likely to be the optimum. In other words, entropy search selects in each iteration the candidate point $\lambda$ that mostly moves $p_{min}$ away from a Uniform to a highly peaked distribution around the global optimum.
For Fabolas, we model $p_{min}$ only at the highest $b=b_{max}$, since we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$ Now, to balance between information gain and the cost of evaluating $\lambda$, Fabolas maintains a second Gaussian process $p(c\mid\lambda,b)$ to model the wall-clock time $c(\lambda,b)$ of evaluating $f_{b}(\lambda)$. In each iteration, Fabolas optimizes the following acquisition function:
$a(\lambda,b)=\frac{KL\big{[}p_{min}(\lambda\mid\mathcal{D})||\mathcal{U}(\lambda)\big{]}}{\mu_{c}(\lambda,b)}$
to select both the configuration $\lambda$ and budget $b$ that provides the most information about $p_{min}$ while being not too expensive to evaluate, where $\mu_{c}(\lambda,b)=\mathbb{E}_{p(c\mid\lambda,b)}[c(\lambda,b)]$ is the mean predicted cost of evaluating $\lambda$ using a budget $b$.
5.3.2 BO-HB
To combine the sample efficiency of Bayesian optimization with the strong anytime performance of Hyperband, falkner2018bohb fits an independent TPE model (see Section 4.3) on each rung level $b_{i}\in\mathcal{R}$ to sample new configurations. More specifically, once the proposed method BO-HB observes $N_{min}$ datapoints on the rung level $b_{i}$, it trains two kernel density estimator as shown in Eq. (4.3) on all data points of this rung level. New configurations are now sampled according to density ratio $l(\lambda)/g(\lambda)$ instead of uniformly at random. This continues until we observed $N_{min}$ on the highest rung level $b_{max}$ and we keep sampling from the model on this rung level. While the original implementation of TPE by bergstra2011algorithms used uni-variant kernel density estimator, falkner2018bohb proposed to use multi-variant kernel density estimator to capture the interaction between hyperparameters. To fit such a model, we need to have at least $N_{min}=m+1$ datapoints where $m$ is the dimensionality of the search space.
Figure 5.3 shows a comparison by [falkner2018bohb] of BO-HB to other approaches for optimizing the hyperparameters of a neural network. Hyperband outperforms in the beginning, as it is able to return configurations much earlier than non-multi-fidelity methods such as random search or TPE. However, compared to model-based approaches such as TPE, Hyperband eventually converges more and more to random search, as it samples configurations uniformly at random and, in later stages, requires evaluating them on the larger budgets to improve. BO-HB, on the other hand, maintains the early speed-up of Hyperband but achieves a similar strong performance than TPE in the later stages of the optimization process.
Figure 5.3: Comparison of random search (RS), TPE, Hyperband (HB) and BO-HB overtime from falkner2018bohb. BO-HB keeps the strong anytime performance of Hyperband but because of the model samples better configurations over time.
Chapter 6 Population-based Methods
In this chapter we discuss a class of randomized methods for global optimization that evolve a population of candidate solution, drawing inspiration from the process of natural evolution [back1997handbook, simon2013evolutionary]. Evolutionary algorithms date back to the beginning of computer science [neumann1966theory] and have been used to tackle non-smooth and non-convex problems across many domains. Methods in this class typically rely on function evaluations only. Their application to hyperparameter optimization includes methods like CMA-ES [hansen1996adapting], population-based training (PBT) [jaderberg2017population] and particle swarm optimization [lorenzo2017particle], which we review below.
6.1 Evolutionary algorithms
Traditionally, evolutionary algorithm (EA) distinguish between evolutionary strategies (ES) and genetic algorithms (GA). GA refers to a class of methods that deal with binary problems and discrete objective functions while ES is concerned with real-valued, continuous objectives.
The terminology of EA techniques is borrowed from biology: the set of points is called population and candidates are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically at the end of each generation the fitness of all individuals is observed, making EA inherently parallel algorithms. Traditional implementations typically also require synchronization. After observing the fitness of the population a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include:
  • Selection, global transformations that prime a new generation, typically by discarding individuals with low fitness and selecting a subset of fitter individuals to become “parents” for the next generation;
  • Recombination or crossover, global transformations that consist in generating new individuals combining slices of two or more parents;
  • Mutation, local modifications designed to escape local optima that introduce random changes in the individual, for instance adding independent Gaussian noise.
Other strategies like elitism (keep the best individuals unchanged) and niching (maintain a diverse population) are often used in conjunction.
Various implementations of ES/GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for more details), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next subsections, we review two particular implementations that have received notable attention.
6.1.1 Covariance matrix adaptation evolution strategy (CMA-ES)
CMA-ES [hansen1996adapting, hansen2016cma, loshchilov2016cma] maintains a multivariate Gaussian distribution from which it samples individuals for the new generations. That is, the generation at iteration $k+1$ is drawn according to
$\lambda_{k+1}^{j}\sim\mu_{k}+\sigma_{k}\mathcal{N}(0,\Sigma_{k})\quad\text{for}j\in[N]$
where $N\in\mathbb{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size.
The parameters of the distribution are evolved at each iteration following a selection and recombination principle. Suppose $\{\lambda_{k}^{j}\}_{j=1}^{N}$ are sorted in increasing objective value, i.e. $f(\lambda_{k}^{1})\leq f(\lambda_{k}^{2})\leq\dots\leq f(\lambda_{k}^{N})$. Then the mean is updated as a convex combination of the best $1\leq M\leq N$ points:
$\mu_{k+1}=\sum_{i=1}^{M}w_{i}\lambda_{k}^{i},$
where the weights $w_{i}$ are positive and sum to 1. The threshold $M$ and the recombination weights are parameter of the method. hansen2016cma suggests $M=N/2$ and $w_{i}\propto M-i+1$ as reasonable defaults.
The covariance matrix is adapted in a similar fashion by reestimating it from the $M$ best points of the previous generation. That is:
$\Sigma_{k}=\sum_{i=1}^{K}w_{i}(\lambda_{k}^{i}-\mu_{k})(\lambda_{k}^{i}-\mu_{k})^{\intercal}$ (6.1)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1) [hansen2019pycma].
6.1.2 Population-based training (PBT)
jaderberg2017population propose an asynchronous method that evolves a population of models with different hyperparameter configurations as they are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by standard gradient descent) with applications of evolutionary transformations to both the parameter of the models and the hyperparameter of the learning algorithm. PBT is an instance of online HPO which construct an hyperparameter schedule rather than finding a single configuration. We will discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Let $A=A_{T}\circ\dots\circ A_{1}$ be a decomposition of the base learning algorithm [^1] Such decomposition is exact when each $A_{t}$ is applied with the same, fixed, hyperparameter configuration.. Each $A_{t}$ could be, for instance, an epoch of gradient-based training of a neural network. Let $N$ be a population size and $\{(\theta_{1}^{j},\lambda_{1}^{j})\}_{k\in[N]}$ be a set of pairs of model parameters and algorithm hyperparameters — the individuals in PBT — initialized from suitable distributions. At each generation, individuals are updated (asynchronously) as
$\theta_{k}^{t+1}=A(\theta_{k}^{t},\lambda^{t}_{k})\quad\text{for }k\in[N].$
After this, some evolutionary transformations are applied to the pairs of parameters and hyperparameters using as fitness the evaluation of the HPO objective $\mathscr{M}$ on the hypotheses provided by the $\theta_{k}^{t+1}$’s. jaderberg2017population propose two type of operations:
  • “exploit,” a selection strategy that involves copying over parameters and hyperparameters from other better performing individuals, for instance, by replacing the worst-performing 20% with randomly drawn individuals from the best-performing 20% (truncation selection);
  • “explore,” a mutation tranformation that either perturbs the hyperparameter configuration or resamples it from the initial distribution. In both cases, the parameters are left unvaried.
PBT further uses elitism. tao2020learning combine PBT with gradient-based optimization of the hyperparameters, which we will discuss in the next chapter.
6.2 Particle swarm optimization
Particle swarm optimization (PSO) methods draws inspiration from the social behaviour of animals, like ant colonies or bird flocks [kennedy1995particle], as well as from the evolution of particle trajectories in a physical system. The general idea is that each candidate solution evolves according to a simple equation that weights a momentum term, or velocity, and one or more attractor terms to steer the trajectory toward better performing points.
Assume for simplicity that $\Lambda=\mathbb{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameters and their velocities, both drawn randomly from suitable distributions (e.g. uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization with the following equation:
$\lambda_{t+1}^{j}$ $=\lambda_{t}^{j}+v^{t}_{j}$
$v_{t+1}^{j}$ $=\alpha_{1}v_{t}+\alpha_{2}r_{l}(\bar{\lambda}^{j}-\lambda_{t})+\alpha_{2}r_{g}(\lambda^{*}-\lambda_{t})$
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathbb{R}^{m}$ are two randomly drawn vectors and $\bar{\lambda}^{j}$ and $\lambda^{*}$ are the particle’s best and population best points found so far, respectively. The second term in the velocity update equation can be interpreted as a local memory attractor, while the third term is a global attractor. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search.
6.3 Discussion
Despite their simplicity and their lack of theoretical guarantees, population-based methods enjoy broad applicability and strong empirical performances. In addition, this class of methods is designed to readily scale to parallel and distributed compute systems. Although traditional EA and PSO implementation assume synchronicity, one may easily derive asynchronous versions of many algorithms. For instance in PSO one may replace the global best $\lambda^{*}$ with the best known solution to individual $j$. However, population-based algorithms require a considerable amount of customization due to the presence of several method parameters, task-specific heuristics and design choices, e.g. to define meaningful mutation or perturbation strategies, undermining their accessibility and applicability to novel algorithms and problems. Besides, they typically require a considerable computational power, benefiting usually from large population sizes.
Finally, we note that one may consider successive halving and related methods, which we covered in Chapter 5, as also falling in the category of population-based methods. However, the foundations of those methods are linked to the bandit problem and to hedging strategies [li2017hyperband], rather than taking inspiration from evolutionary principles.
Chapter 7 Gradient-based Optimization
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Of course assuming that $f(\lambda)$ be differentiable excludes discrete hyperparameters such as discrete choices (for example, which nonlinear activation function should be used in a layer of a neural network) or integer-valued hyperparameters (such as the dimension of a layer). There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. For example, [maclaurin2015gradient] suggested to consider the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning (see also Section 8.4). For example for example, the idea of tuning initial weights from several learning tasks was shortly after successfully exploited in [finn_model-agnostic_2017] in their model-agnostic meta learner.
7.1 Preliminaries
At a high level, gradient-based HPO methods start from an initial hyperparameter configuration $\lambda_{0}\in\Lambda$ and iteratively and sequentially update such configuration as follows:
$\lambda_{r}=\mathrm{Proj}_{\Lambda}\left[\lambda_{r-1}-\nu\hat{\nabla}f(\lambda_{r-1})\right]\qquad r=1,2,\dots$ (7.1)
where ^$\hat{\nabla}f$ informally denotes some approximation of the gradient of the HPO objective $f$, $\nu>0$ is a step-size and $\mathrm{Proj}_{\Lambda}$ is the projection onto $\Lambda$. Gradient-based optimization of the loss function inside the learning algorithm $A$ is extremely common, especially in deep learning. Hence, to avoid confusion, we will call ^$\hat{\nabla}f$ the hypergradient.
Not every HPO problem is suitable for gradient-based optimization. Recalling the general formulation of the HPO problem of Section 2.2, one underlying assumption shared by all gradient-based HPO methods is that the performance metric $\mathscr{M}$ be differentiable w.r.t. the model output of $A$. In this chapter, to simplify the notation we take the output of $A$ to be a parameter vector in some weight space $\mathcal{W}\subseteq\mathbb{R}^{d}$ and let the performance metric take values in the weight space directly, that is $\mathscr{M}:\mathcal{W}\times\mathscr{D}\ \to\mathbb{R}$. Hence
$f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})=\mathscr{M}(A(\mathcal{D},\lambda),\mathcal{V}).$ (7.2)
Hereafter, we will drop training and validation sets.
In order to compute hypergradients, it is also necessary to specify at least some details of the learning algorithm $A(\mathcal{D},\lambda)$ that appears in (2.4) and that we have mostly treated as an opaque box until now. Gradient-based HPO methods are distinguished mainly by the assumptions they make about $A$, which then translate into different computational schemes for ^$\hat{\nabla}f$. Next, we review the main approaches.
7.2 Hypergradients via implicit differentiation
Many learning algorithms may be conceptually expressed as solutions to optimization problems
$A(\mathcal{D},\lambda)\in\operatorname*{arg\,min}_{u\in\mathcal{W}}L(u,\lambda),$ (7.3)
where $L:\mathcal{W}\times\Lambda\to\mathbb{R}$, with $\mathcal{W}\subseteq\mathbb{R}^{d}$ is an objective function (the training error) that depends on the data.  [^1] We leave such dependence implicit for improving readability. If $L$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
$\nabla_{u}L(u,\lambda)=0.$ (7.4)
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $L$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}L\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [krantz2012implicit] to (7.4) which assures us that (locally) there exist a function $w:U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}L(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. In such case, we can locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
$[\nabla f(\lambda)]^{\intercal}=\partial_{\lambda}f(\lambda)=\partial_{w}\mathscr{M}(w(\lambda),\lambda)\partial_{\lambda}w(\lambda)\qquad\forall\lambda\in U,$ (7.5)
where, again due to the implicit function theorem,
$\partial_{\lambda}w(\lambda)=-\left[\partial_{u}^{2}L(u,\lambda)_{|u=w(\lambda)}\right]^{-1}\partial^{2}_{u,\lambda}L(u,\lambda)_{|u=w(\lambda)}.$ (7.6)
Equations (7.5) and (7.6) give us the analytic expression of the exact gradient of the HPO objective $f(\lambda)$. Such expression is, however, only implicit. In fact, we almost never have an explicit formula for the map $w(\lambda)$.
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging (7.6) into (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}L\right]^{-1}\in\mathds{R}^{d}$ and approximately solve the linear system
$\partial^{2}_{u}L(u,\lambda)_{|u=w_{T}}q=-\nabla_{u}\mathscr{M}(u)_{|u=w_{T}}.$ (7.7)
Call $q_{T,K}$ an approximate solution of (7.7). Then an approximate gradient of the outer objective is given by
$\nabla_{T,K}f(\lambda)=\left[\partial^{2}_{u,\lambda}L(u,\lambda)_{|u=w_{T}}\right]^{\intercal}q_{T,K}.$ (7.8)
The two possible sources of approximations stem from ($i$) the solution of the inner problem and ($ii$) of the linear system (7.7). Particular implementation of this computation scheme depend on how (7.7) is solved. One classic procedure is the conjugate gradient (CG) method [hestenes1952methods], which features a linear rate of convergence that depends on the conditioning number of the Hessian of the inner objective.
This computational approach has been studied in [larsen_design_1996, larsen1998adaptive, bengio_gradient-based_2000, chapelle2002choosing, seeger2007cross]. More recently, pedregosa_hyperparameter_2016 proposed an approximation scheme whereby solutions to the inner problems and to the linear equation (7.7) become more precise as the number of HPO iterations $r$ grows, proving convergence for algorithms which feature strongly convex (inner) objectives.
7.3 Hypergradients via iterative differentiation
Methods based on iterative differentiation further assume that solutions to the learning algorithm problem (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
$s_{t}=\Phi_{t}(s_{t-1},\lambda,\mathcal{D})\ \ t=1,\dots,T.$ (7.9)
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathds{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$ and a (possibly empty) set of auxiliary variables the $v_{t}$, such as velocities and moments for algorithms such as Adam and RMSProp. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition map $\Phi:(\mathds{R}^{d}\times\mathds{R}^{m}\times\mathscr{D})\rightarrow\mathds{R}^{d}$ is assumed to be differentiable and describes the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g. for stochastic gradient descent) and it internally requires the computation of the inner objective gradients
$g_{t}\doteq\nabla_{w}\bigl{.}L(w,\lambda,\mathcal{D}_{t})\bigr{|}_{w=w_{t-1}}.$
Note that $\Phi$ depends both explicitly and via $s_{t-1}$, and $\lambda$ might contain parameters related to the optimizer itself such as the learning rate. As an example, the dynamics (7.9) for stochastic gradient descent with momentum are
$\begin{array}[]{lcl}v_{t}&=&\mu v_{t-1}+g_{t}\\ w_{t}&=&w_{t-1}-\eta v_{t-1}\end{array}$ (7.10)
where $\mu$ and $\eta$ are the momentum term and the learning rate, and may be included in $\lambda$ if desired.
In a nutshell, iterative differentiation schemes for gradient-based HPO rely on the application of automatic differentiation (AD) to the composite map $\mathscr{M}\circ\Phi_{T}\circ\dots\circ\Phi_{1}$. We briefly survey AD next, and refer the reader to [Griewank08:_evaluat_deriv] for an in-depth account.
7.3.1 Automatic differentiation
Automatic (or algorithmic) differentiation (AD) [Griewank08:_evaluat_deriv, baydin2017automatic] is the standard tool for computing gradients in deep learning. Suppose we have a generic differentiable function $G:\mathds{R}^{N}\mapsto\mathds{R}^{M}$ with $(\eta_{1},\dots,\eta_{M})=G(\xi_{1},\dots,\xi_{M})$ and we are interested in computing the Jacobian matrix, with entries $\frac{\partial\eta_{i}}{\partial\xi_{j}}$. In AD, we begin by describing $G$ as a computational graph $(V,E)$, which is a directed acyclic graph whose nodes $V$ include inputs $\xi_{j}$, outputs $\eta_{i}$, and intermediate variables that are computed by applying elementary functions to inputs or other intermediate variables, that can be scalars, vectors or tensors. Elementary functions are typically atomic operations such as (tensor) sum, (Hadamard) product, and (elementwise) trascendental functions, whose partial derivatives are readily known. Edges $E$ in the graph describe functional dependencies, i.e., they link intermediate variables to the elementary functions they are fed into. Forward propagation, following a topological sort of the graph, allows us to compute $G(\xi_{1},\dots,\xi_{M})$. Two AD strategies are available, forward and reverse mode, which are based on the definition of local derivatives and a message-passing procedure based on the standard chain rule of calculus.
In forward mode AD, the local derivative of $v_{k}$ is defined as
$\dot{v}_{k}\doteq\frac{\partial v_{k}}{\partial\xi_{j}}$
and can be updated by the recursion
$\dot{v}_{k}=\sum_{\ell\in\pi_{k}}\dot{v}_{\ell}\frac{\partial v_{k}}{\partial v_{\ell}}$ (7.11)
initialized as $\dot{\xi_{j}}=1$, where $\pi_{k}$ are $v_{k}$’s parents, and iterated following a topological sort of $(V,E)$. In the above equation, if $v_{k}$ and $v_{\ell}$ are vector-valued, $\frac{\partial v_{k}}{\partial v_{\ell}}$ denotes a Jacobian matrix. The desired derivatives can be finally read as $\dot{\eta}_{i}$. The same propagation allows us to compute the $j$-th column of the Jacobian and to obtain the whole Jacobian, propagation needs to be repeated for each input $\xi_{j}$. Assuming elementary functions can be computed in $O(1)$, the total computational cost is $O(NE)$. Forward mode is thus advantageous for functions with few inputs and many outputs. If on the other hand $N=O(E)$, as it happens for example for the objective function of neural networks, forward mode is as inefficient as computing derivatives numerically. It is therefore not surprising that forward mode is never used to train neural network models (although there is a remarkable exception when training recurrent networks on data streams [williams_learning_1989], a setting where reverse mode cannot be applied).
In reverse mode (the strategy used by the well known backpropagation algorithm [rumelhart1986learning]) the local derivative of $v_{k}$, usually called adjoint, is defined as
$\bar{v}_{k}\doteq\frac{\partial\eta_{i}}{\partial v_{k}}$
and can be updated by the backward recursion
$\bar{v}_{k}=\sum_{\ell\in\gamma_{k}}\bar{v}_{\ell}\frac{\partial v_{\ell}}{\partial v_{k}}$ (7.12)
initialized as $\bar{\eta_{i}}=1$, where $\gamma_{k}$ are $v_{k}$’s children, and iterated following a reversed topological sort of $(V,E)$. The desired derivatives can be finally read as $\bar{\xi}_{j}$. The same propagation allows us to compute the $i$-th row of the Jacobian and to obtain the whole Jacobian, propagation needs to be repeated for each output $\eta_{i}$, for a total computational cost of $O(ME)$. Reverse mode is thus advantageous for functions with many inputs and few outputs. However, since the computation proceeds in a backward way, the memory used to store the values of intermediate nodes $v_{k}$ cannot be freed until the end of the propagation, yielding a memory cost of $O(V)$.
7.3.2 Computing hypergradients
We are now ready to apply the principles of AD to the differentiation of the validation error with respect to $\lambda$. A simplified view of the computational graph mapping $\lambda$ into the response $f(\lambda)$ is shown in Figure 7.1, where minibatches $\mathcal{D}_{t}$ and initial weights $w_{0}$ are supplied as given constants [^2] Note that $w_{0}$ could be equally well included in $\lambda$ but here we keep it distinct for clarity.. Here the number of inputs $M$ is the dimension of the hyperparameter vector and there is a scalar ($N=1$) output. Forward and reverse calculations are presented in [franceschi2017forward] and repeated below.
Figure 7.1: Computational graph of the response function for forward and reverse mode hypergradient calculation.
In the case of forward mode, the local variable is $\dot{s}_{t}$ and is udpated by the recursion
$\dot{s}_{t}\doteq\frac{\partial s_{t}}{\partial\lambda}=\left.\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}\dot{s}_{t-1}\right|_{a=s_{t-1}}+\left.\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}\right|_{b=\lambda}$ (7.13)
where $\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}$ and $\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}$ are $d\times d$ and $d\times m$ Jacobian matrices, respectively. The base step is $\dot{s}_{0}=\frac{\partial s_{0}}{\partial\lambda}$ which is zero unless the hyperparameters directly affect the initial weights (e.g., their range, if randomly initialized, or even their values as in maclaurin2015gradient). The desired gradient is finally obtained at the last node:
$\nabla f(\lambda)=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}\dot{s}_{T}$
where $w_{T}$ is the proper slice of $s_{T}$.
In the case of reverse mode, the adjoint $\bar{s}_{t}$ is initialized as
$\bar{s}_{T}=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}$
and updated as
$\bar{s}_{t}\doteq\frac{\partial f}{\partial s_{t}}=\left.\bar{s}_{t+1}\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}\right|_{a=s_{t-1}}.$
Since the variable $\lambda$ is a parent of all $s_{t}$, Eq. (7.12) requires us to sum all the contributions:
$\nabla f(\lambda)=\sum_{t=1}^{T}\bar{s}_{t}\left.\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}\right|_{b=\lambda}.$ (7.14)
It must be remarked the striking similarity between hypergradient calculations and gradient calculations in recurrent neural networks (RNN). Indeed, the computational graphs are identical, where $s_{t}$ matches the “state” of an RNN, $\lambda$ the weights, and $\Phi$ the state update function. Equations (7.13) and (7.14) closely resemble real-time recurrent learning [williams_learning_1989] and backpropagation through time [werbos1990backpropagation], respectively.
7.4 Fixed-point methods
Another possible approach that sits in between the iterative and the implicit differentiation schema is given by the so-called fixed-point method. The method is based on the observation that one can rewrite the condition (7.4) as a fixed point equation of an appropriate contractive dynamics
$w(\lambda)=\Phi(w(\lambda),\lambda),$ (7.15)
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla L$.  [^3] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to (7.15) yields to
$\frac{dw(\lambda)}{d\lambda}=\left(I-\frac{\partial\Phi(w(\lambda),\lambda)}{\partial w}\right)^{-1}\frac{\partial\Phi(w(\lambda),\lambda)}{\partial w\partial\lambda}$
and, repeating the procedure described above, Equation (7.7) becomes
$\left(I-\partial_{w}\Phi(w_{T},\lambda)\right)^{\intercal}q=\nabla\mathscr{M}(w_{T}).$ (7.16)
Starting from any point $q_{T,0}\in\mathds{R}^{d}$, one can iterate
$q_{T,k}=\partial_{w}\Phi(w_{T},\lambda)^{\intercal}q_{T,k-1}+\nabla\mathscr{M}(w_{T})\qquad k\in[K]$ (7.17)
to find an approximate fixed-point of (7.16) and then plug $q_{T,K}$ into Equation (7.8) to obtain an hypergradient.
This computational scheme emerges in more general cases where $\Phi$ is not necessarily an optimization dynamics: $\Phi$ may represent a stable recurrent neural network [miller2019stable], a graph neural network [scarselli2009graph] or an equilibrium model [bai2019deep]. In these contexts, the fixed-point algorithm is usually known as recurrent backpropagation [almeida1987learning, liao2018reviving].
Warm-starting $q_{T,0}$ with the vector obtained at the previous optimization iteration (of $\lambda$) may speed-up the convergence of (7.17).
7.5 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g. for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems.
[maybe move and shorten] Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to avoid completely the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to $(j+1)C-1$ to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered in [shaban2019truncated], together with a more radical truncation strategy (similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90]), which however does not recover exact hypergradients.
Chapter 8 Further Topics
In the last six chapters we covered the main classes of algorithms for hyperparamter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization, online HPO and hyperparameter schedules and neural architecture search. We conclude the chapter with a section on meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting.
8.1 Optimizing for multiple objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example when working with hardware-constraint systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front – the set of all non-dominated solutions – to help users take an informed trade-off decision.
More formally, consider a function $f:\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be a $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda\_1$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination.
Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally: $\mathcal{P}_{f}=\{\lambda\in\Lambda\,|\,\nexists\,\lambda^{\prime}\in\Lambda:\lambda^{\prime}\succ\lambda\}$.
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $A\subset\Lambda$ of objective vectors near the Pareto front. The quality of $A$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $\mathbf{r}$, $\mathcal{H}(A)$ equals the hypervolume dominated by $A$:
$\mathcal{H}(A)=\text{Vol}\left(\{\lambda\in\mathcal{Y}=\mathbb{R}^{n}\,|\,\exists\,\lambda^{\prime}\in A:\lambda^{\prime}\succeq\lambda\wedge\lambda\succeq\mathbf{r}\}\right).$
Computing $\mathcal{H}(A)$ involves partitioning the space into hypercubes, an operation that scales exponentially with the number of objectives $n$. This exponential scaling poses a key bottleneck for hypervolume-based multi-objective optimizers.
Examples
A popular example of MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to acceptance-based and error-based, can be in conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some (albeit) imperfect extent and expose empirical trade-offs between various definitions can still be important from societal perspectives.
Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a joint or conditionally independent random function model. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.
8.1.1 Multi-objective optimization
Early efforts on HPO focused on designing evolutionary optimization techniques for neural networks and SVMs [Yao1999:Evolving, Nolfi2002:Evolution, Igel2005:Evolutionary, friedrichs2005evolutionary]. Evolutionary techniques have also been applied to multi-objective HPO problems [Jin2006:Multi]. To mitigate the large computational cost of evaluating a single hyperparameter configuration (e.g., training a DNN), sample-efficient Bayesian optimization (BO) techniques have been introduced [hutter_sequential_2011, snoek_practical_2012, bergstra2011algorithms].
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective (SO) BO techniques. The scalarization is defined by using a weight vector $\mathbf{w}\in\mathbb{R}^{n}$, solving the MO minimization problem consists of solving the following optimization problem:
$\min_{\lambda\in\Lambda}V(f(\lambda),\mathbf{w})$ (8.1)
where $V:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a scalarization function. While these methods are comparatively easy to implement and scale gracefully with the number of objectives, they do not utilize information about the overall geometry of the current Pareto front approximation. Some common choices for $V$ are:
  • inner product between objective and weight vectors (also known as random weights): $V_{\text{RW}}(f(\lambda),\mathbf{w})=f(\lambda)^{\top}\mathbf{w}$;
  • the scalarization used in [Knowles2006:Parego], called ParEGO: $V_{\text{ParEGO}}(f(\lambda),\mathbf{w})=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}\mathbf{w})$; the scalarization used in [Golovin2020]: $V_{\text{Golovin}}(f(\lambda),\mathbf{w})=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}$.
Note that by optimizing different scalarizations, solutions along the Pareto front can be obtained. The choice of $V$ influences the subset of the Pareto front that is found, and the right selection might change depending on the task at hand.
A second class of MBO techniques builds on a performance measure of Pareto front approximations, namely the dominated hypervolume [Zitzler1998:Multiobjective]. Both the expected hypervolume improvement (EHI) [Emmerich2011:Hypervolume] and probability hyper-improvement (PHI) [Keane2006:Statistical] operate by extending their single-objective BO counterparts—expected improvement [Mockus1978:Application, Jones1998:Efficient] and probability of improvement [Kushner1964:New], respectively. Other methods include step-wise uncertainty reduction (SUR) [Picheny2015:Multiobjective], smsEGO [Ponweiser2008:Multiobjective, Wagner2010:Expected] and expected maximin improvement (EMMI) [Svenson2010:Multiobjective].
Finally, information-theoretic MBO approaches aim to select points that reduce uncertainty about the location of the Pareto front. Methods in this class tend to be more sample efficient and scale better with the number of objectives. PAL [Zuluaga2013:Active] iteratively reduces the size of a discrete uncertainty set. PESMO [Hernandez2016:Predictive] adapts the predictive entropy search (PES) [Hernandez2014:Predictive] criterion. Pareto-frontier entropy search (PFES) [Suzuki2019:Multi] is suitable when dealing with decoupled objectives. MESMO [Belakaria2019:Max] builds on the max-value entropy-search criterion [Wang2017:Max] and enjoys an asymptotic regret bound. Building on MESMO, two very recent works have proposed MF-OSEMO [Belakaria2020:Multi] and iMOCA [Belakaria2020:Information], two multi-fidelity based information-theoretic MBO techniques which internally use multi-fidelity Gaussian processes.
8.1.2 Constrained hyperparameter optimization
Our goal in a constrained optimization scenario is to minimize the target black-box $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda):\Lambda\rightarrow\mathbb{R}$ and $c(x):\Lambda\rightarrow\mathbb{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$f_{\star}=\min_{\lambda\in\Lambda}\left\{f(\lambda)\;\|\;c(\lambda)\leq\delta\right\},$ (8.2)
where $\delta\in\mathbb{R}$ is a threshold on the constraint. For instance, in the case of BO, the latent functions $f(\lambda)$ and $c(\lambda)$ can be conditionally independent in our surrogate model, with different GP priors placed on them.
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assumes that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent function can be represented as GPs with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a DNN fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash, neither do we usually know the exact amount of memory available on the compute instance. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.2) as follows:
$y_{\star}=\min_{x{}\in\mathcal{X}}\left\{y(x{})\;\|\;P(z_{c}=+1|x{})=\sigma(c(x{}))\leq\sigma(\delta)\right\}.$
This formulation is similar to the one proposed in [Gelbart14]. The parameter $\sigma(\delta)\in(0,1)$ controls the size of the (random) feasible region for defining $y_{\star}$. Finally, note that in the example of out of memory errors training failures, the criterion observation $z_{y}$ is obtained only if $z_{c}=-1$: if a training run crashes, a validation error is not obtained for the queried configuration. Apart from an experiment in [Gelbart14], the case of contrained BO work covering the binary feedback case is covered in [perrone_cmes].
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. If the constraint is denoted by $c(\lambda)\leq\delta$, a separate regression model is used to learn the constraint function $c(\lambda)$ (typically a GP), and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=P\{c(\lambda)\leq\delta\}EI(\lambda)$, where $P\{c(\lambda)\leq\delta\}$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.
Several issues with cEI are detailed in [Lobato15a]. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter equivalent to ours features in [Gelbart14], but is only used to recommend the final hyperparameter configuration. Another approach was proposed in [Lobato15a], where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case are developed in [Picheny2016] and [Ariafar2019]. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. A better option is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
8.2 Hyperparameter schedules and online HPO
The vast majority of learning algorithms do not produce an hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapter we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.1.2). When this is the case, it can make sense to consider hyperparameter schedules rather than fixed values. One of the most prominent example in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A:\mathscr{D}\times\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}:\mathcal{H}\times\Lambda\times\mathscr{D}\rightarrow\mathcal{H}$. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\mathcal{D},\lambda)=\Phi_{T}(\Phi_{T-1}(\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1},\mathcal{D}),\lambda_{T-1},\mathcal{D}),\lambda_{T},\mathcal{D}).$ (8.3)
for some $T>0$, where we generally view $\lambda\in\Lambda^{T}$ as an hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other predefined sequence of operations..
Hyperparameter iterates may be left “free to float”, in which case one can formulate a related HPO problem following the general setup of Section 2.2, but this time searching for configurations in $\Lambda^{T}$ rather than in $\Lambda$. This is seldom done in practice due to the complications arising from an expanded search space, few exceptions being [maclaurin2015gradient, wu2018learning]. More commonly, the iterates are expressed as a function of the step $t$ and a set of other fixed hyperparameters $\lambda^{\prime}\in\Lambda^{\prime}$ Learning rate schedules are a primary example. For instance, a triangular schedule (i.e. warmup followed by linear decay) is given by:
$\eta_{t}=\left\{\begin{array}[]{ll}\eta_{\mathrm{max}}(t/T_{\mathrm{max}})&\text{ for }\,t\geq T_{\mathrm{max}}\\ \eta_{\mathrm{max}}(1-(t-T_{\mathrm{max}})/(T-T_{\mathrm{max}}))&\text{ for }\,T_{\mathrm{max}}<t\leq T\end{array}\right.$
where $\lambda^{\prime}=(\eta_{\mathrm{max}},T_{\mathrm{max}})$. Once again, when defining a related HPO problem, one may fall back to the setting of Section 2.2, optimizing over $\Lambda^{\prime}$ rather than $\Lambda$. Along these lines, one may even let the hyperparameter iterates be a function of a state (e.g. the weights of the trainee model at step $t$), an approach that has been explored e.g. by hansen2016using using reinforcement learning to fit a learning rate adjustment policy or lorraine2018stochastic using hyper-networks and gradient-based HPO.
However, perhaps one main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparamter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors. Online HPO remains an area of active research.
8.3 Neural architecture search
Deep learning has revolutionized the traditional approach of hand-crafting features by enabling the automatic learning of features during training. However, the process of creating neural network architectures still depends significantly on manual design, which involves a substantial amount of trial and error. We can extend the idea of hyperparameter optimization to automatically design neural network architectures, a process known as neural architecture search (NAS).
NAS shares similarities with hyperparameter optimization (HPO), but focuses solely on the hyperparameters that define the architecture. However, NAS faces challenges distinct from most HPO problems, as its search spaces are often significantly higher dimensional, exclusively discrete, and of highly conditional structure (e.g., the hyperparameter controlling the number of units in a given layer is only applicable if that layer exists, as controlled by another hyperparameter).
8.3.1 Search Space
A key question in NAS is how to define a search space to parameterize a neural network architecture. An expressive search space might lead to entirely new and innovate architecture but might be also prohibitively expensive to explore. On the other hand, a too restrictive space will lower the chance to improve upon manually design architectures. We discuss two popular approaches to define NAS search spaces [white-arxiv23]: Marco Search Spaces and Cell Search Spaces.
Macro search spaces.
We assume a pre-defined architecture topology, such as for example residual neural network or a transformer architecture. Macro search spaces consider design decisions to configure the architecture on a macro level, for example whether to use pre- or post-layer normalization or how many heads to use for all multi-head attention layers.
For example, MNASNet [tan-cvpr19] separates an architecture into blocks and for each block optimizes different hyperparameters, such as the convolution type or kernel size. NASBOT [kandasamy-neurips18] defines the entire network as a DAG where nodes perform a certain operation and edges represent direct dependencies between nodes, or in other words, the output of one node is the input of an adjacent node. For each node, the NAS algorithm has to select the operation type and the number of convolutional filters or fully connected nodes.
The advantage of macro search spaces is that they allow parameterizing each block or layer of the architecture individually. However, while this leads to a very expressive space, it will be quite expensive to search.
Cell search space.
Compared to the Macro Search Space, the Cell Search Space considers a block or cell which is a smaller computational unit that is stacked multiple times on top of each other to form the final architecture. zoph:2018:learning_transferable defined each cell as a directed acyclic graph where for each node they select a certain operation to be performed, for example 1x1 convolution or 3x3 convolution and how to place the edges in the graph. To reduce the overall complexity, zoph2016neural distinguished between two different types of cell, which share the same structure but different strides to reduce the spatial dimension of the input.
Instead of assigning operations to each node, liu2018darts assigns operations to every edge of the graph. Nodes that accumulate the feature vectors from all incoming edges. Their proposed method DARTS then uses gradient descent to select an operation for each edge.
ying-icml19 proposed NASBench-101 which consists of a cell search space with a maximum of 7 nodes and 3 operation choices per node. Furthermore, they limited the total number of edges to 9, which led to a total of 400K different architectures. In an expensive offline step, all architectures were evaluated on CIFAR-10 to form a table. This allows running NAS methods by simply looking up the performance of an architecture instead of training it, accelerating the search process by multiple orders of magnitude.
In follow-up work,  dong-iclr20 proposed NASBench-201 which reduced the number of nodes but allowed for more operations per node. While the total number of architectures in this search space is substantially smaller than in NASBench-101, there are no constraints on the total number of edges, which makes it more accessible for different search algorithms. Furthermore, NASBench-201 evaluated each architecture on CIFAR-10, CIFAR-100 and Imagenet-16.
8.3.2 Search strategies
How do we explore the search space? This is particularly challenging in the case of NAS because of the large number of possible architectures in the search space and the expensive nature of training neural network models.
In one of the first works on NAS, zoph2016neural used a controller based on a recurrent neural network to suggest new architectures. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. real2019regularized proposed an evolutionary approach that maintains a population of architectures. In each iteration, a random architecture is drawn from the population and mutated along a single dimension. While these approaches led to architectures outperforming state-of-the-art models at that point, the original methods required substantial computational resources since each network was trained from scratch.
To accelerate the search process, we can utilize recent advancements from HPO, such as multi-fidelity optimization. For example, there is a body of work [white-neurips21] that tries to predict learning curves of neural networks to stop the training process early.
A more radically different approach to accelerate NAS is so-called weight-sharing NAS. The idea is to define a single super-network that contains all possible architectures in the finite search space. Each path through the super-network represents a single architecture. In the original work by pham-icml18, this method accelerates the search by more than a factor of 1000.
In a similar vein, liu2018darts proposed DARTS, which adds a continuous architecture parameter for every operation at each edge in the super-network. These architecture parameters are jointly optimized with the weights of the neural network in a bi-level fashion using variations of algorithms we described in Chapter 7. The inner loop fits the weights of the super-network to minimize the training loss, and the outer loop optimizes the architecture parameters to minimize the validation loss. After training the super-network, the final architecture is obtained by selecting the $k$-most important operations per edge. This architecture is then usually re-trained from scratch.
Another approach is so-called two-stage NAS [cai-iclr20, yu-eccv20], which first trains the super-network and then runs gradient-free optimization methods, such as random search or evolutionary algorithms, to select the optimal sub-network without any further training. For example, cai-iclr20 proposed the once-for-all model, which provides a single super-network that allows selecting the optimal architecture with weights for a given hardware, minimizing both latency on the target device and validation error.
8.4 Meta-learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to hyperparameter optimization. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta learning, see e.g., [hospedales:2021:metalearning_neural_networks] and references therein.
8.4.1 Definition
In meta-learning (also called learning-to-learn), we have a sequence of tasks $\mathcal{T}^{1},\mathcal{T}^{2},\dots,\mathcal{T}^{N}$, where each task $\mathcal{T}^{j}=(p^{j},\mathcal{L}^{j})$ consists of a distribution $p^{j}(x,y)$, $x\in\mathcal{X}$, $y\in\mathcal{Y}^{j}$ over data points and a loss function $\mathcal{L}^{j}$. Note that all tasks share the same input space but their label spaces may differ. Tasks are sampled from a meta distribution $\mathcal{P}$ on tasks $(\mathcal{T})$. Unlike a traditional supervised algorithm (which, given a dataset, will produce a prediction function $h$), a meta-learning algorithm outputs a learning algorithm that is hoped to learn a good prediction function when presented with a new (unseen) task $\mathcal{T}=(p,\mathcal{L})$ sampled from the same meta distribution $\mathcal{P}$. The meta-training phase consists in computing the learning algorithm $A=ML(\mathcal{D}^{1},\dots,\mathcal{D}^{N})$ where $\mathcal{D}^{j}\sim p^{j}$ and $ML$ is the meta-learning algorithm. The meta-test phase, in turn, consists in computing the prediction function using the new task: $h=A(\mathcal{D})$, with $\mathcal{D}\sim p$.
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multi-task learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several source tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\leavevmode\nobreak\ \sim p$ will be available for the new task (few-shot learning, e.g. $|\mathcal{D}|=5$), or even a single example (one-shot learning). The main insight is that when the tasks are related, as capture by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit].
8.4.2 Meta-learning as bilevel programming
As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}:\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}:\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}:\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathbb{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multi-task network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define
$f(\lambda)=\sum_{j=1}^{N}\mathcal{L}^{j}\left(w^{j},\lambda,\mathcal{D}^{j}_{\text{val}}\right).$
Figure 8.1: Tasks and datasets for meta-learning on Omniglot [lake:2019:omniglot_challenge_3year]. At the end of the meta-training phase, a small training set for the the new task $\mathcal{T}$ is provided and performance is evaluated on test data $\mathcal{D}_{\text{test}}$. A similar figure was shown in [ravi_optimization_2016] for tasks derived from Imagenet.
We are now ready to formulate the following bilevel program
$\min_{\lambda}$ $f(\lambda)$ (8.4)
s.t. $w^{j}\in\text{arg}\min_{u}\mathcal{L}^{j}\left(w^{j},\lambda,\mathcal{D}^{j}_{\text{tr}}\right)$ (8.5)
where the lower level (or inner) objectives $\mathcal{L}^{j}$ are empirical risk minimization on each training tasks, and the upper (or outer) objective is the validation error on the same tasks. If $\Phi$ and $\gamma$ are realized by deep learning modules, the inner problem would be solved by some iterative gradient-based algorithm, implying that the solution would only be found approximately. Moreover, in general, the above problem is not guaranteed to have a solution. In [franceschi2018bilevel] it is shown that a solution is guaranteed to exist if $\gamma$ is linear (i.e., the classification head consists of a single dense layer) and an algorithm based on hypergradients (using the method described in  Section 7.3) is presented to solve the meta-learning problem. A closely related algorithm is proposed in [finn_model-agnostic_2017] although it is not explicitly formulated as a bilevel program and the emphasis is on fine-tuning to the target task an architecture pretrained on the sequence of previous tasks. A solution based on implicit gradients (see Section 7.2) is offered in [rajeswaran2019meta].
Lastly, while we presented meta-learning as shared feature learning, we note that other tasks’ sharing knowledge mechanisms are possible. Notable examples are biased regularization [denevi2018learning, denevi2019learning, balcan2019provable], in which the meta-parameter $\lambda$ describes a common mean among the tasks, hyper-network approaches, in which $\lambda$ is associated to a function that outputs the parameters of a task-specific models [zhao2020meta, zakerinia2024more], and conditional meta-learning [denevi2020advantage, denevi2022conditional, rusu2018meta, wang2020structured], in which $\lambda$ is a conditioning function of side information specific to the task at hand.
Chapter 9 Hyperparameter Optimization Systems
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO is a fast-moving field, and the current landscape is changing frequently, so the provided list here is by no means exhaustive and only represents a snapshot.
9.1 Open Source Libraries and Frameworks
One of the first libraries that provides Bayesian optimization for hyperparameter optimization was the Python package Spearmint [snoek_practical_2012], which implements Bayesian optimization using Gaussian processes (see Section 4.1). At the same time, bergstra2011algorithms provided HyperOpt, offering Bayesian hyperparameter optimization by using the tree-parzen estimator instead of Gaussian processes. Hutter2011:Sequential introduced SMAC, which implemented Bayesian optimization strategies using Random Forests. Initially targeting for more generic algorithm configuration problems, such as configuring SAT solvers, it also offered functionality for HPO of machine learning methods. While the initial version of SMAC was written in Java, there is also a newer version fully implemented in Python [smac3], that also provides more advanced functionalities, such as multi-fidelity optimization (see Chapter 5).
Some of the first Python libraries that implemented multi-fidelity HPO, such as Fabolas [klein2017fast] or Hyperband [li2016hyperband], was the RoBO library by klein-bayesopt17 and Dragonfly by kandasamy-jmlr19.
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on the large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distributed the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
Recent pytorch implementations of gradient-based hyperparameter optimziation algorithms include the libraries HyperTorch [grazzi2020iteration], higher [grefenstette2019generalized] and Betty [choe2023betty]. JaxOpt [blondel2022efficient] is a Jax implementation that focuses on implicit differentiation.
Most HPO libraries or frameworks expect the objective function either as function handle, for example in Python, or as a single training script. The search space is usually defined outside of the actual training script and hyperparameters are passed as input arguments. Optuna [akiba-sigkdd19] follows a different approach by a so-called define-by-run API that allows users to dynamically define the search space instead of providing it in advance.
9.2 Commercial Systems
Since hyperparameter optimization plays such an integral role in the machine learning workflow, most machine learning platforms, for example Weights and Biases, provide some functionality to select hyperparameters. There are also several commercial systems specific for automated hyperparameter optimization. For example, Google Vizier [google_vizier] is an internal service that contributed to several Google products and research efforts. oss_vizier also introduces a open-source version of Google Vizier as a stand-alone Python package. AWS SageMaker provides a fully automated service called Automated Model Tuner [perrone2021sagemaker] for gradient-free optimization problems, which allows for large-scale HPO leveraging built-in SageMaker functionalities. Another example is Sigopt, which provides optimization-as-a-service for hyperparameter optimization utilizing Bayesian optimization techniques.
9.3 Benchmarking
A key challenge for an empirical driven research field such as HPO to make progress is the thorough evaluation of new approaches and comparison to baseline methods. However, due to the expensive nature of the underlying optimization problem (see Chapter 2), it is often prohibitively expensive to obtain statistically reliable results including multiple repetitions, which are necessary to characterize HPO algorithms that often come with intrinsic randomness.
To reduce the computational requirements and enable large-scale experimentation, recent work tries to emulate expensive HPO problems by using cheap-to-evaluate approximations of the underlying objective function. eggensperger-aaai15 introduced the idea of surrogate benchmarks that replace the original benchmark with the predictions of a machine learning model. This model is first trained in an offline step on data collected from the original benchmark. Unlike the probabilistic model used in Bayesian optimization, where data is generated iteratively, all data in this approach is generated simultaneously, and the model is trained only once. While using a machine learning model to approximate the objective function introduces potentially some bias, single function evaluations are now orders of magnitude cheaper, since they only require a single prediction.
Instead of using a machine learning model to approximate the true objective, klein-arxiv19a first discretzied the search space of a fully-connected neural network leading to a finite number of hyperparameter configurations and then performed an exhaustive search to collect the performance and runtime of each configuration. This allows to replace the actual training and validation of a hyperparameter configuration by simple table look-ups during the HPO process. To mimic the observation noise that comes with the HPO of neural networks, each HPO configuration was evaluated multiple times. During the HPO process, random entries for each hyperparameter configuration are returned to simulate the observation noise. The idea of tabularized benchmarks has since been extended to neural architecture search [ying-icml19, dong-iclr20] (see Section 8.3) leading to several new benchmarks. While tabularized benchmarks return for each query truly observed values of the objective function instead of prediction of the model, pfisterer-automl22 argued that the discretization of the search space eliminates the local structure of the objective function, leading actually to a higher bias than using surrogate models. The idea of surrogate benchmarks has been extended to define a generative model by [klein-nips19], which allows sampling an arbitrary set of new benchmarks.
To standardize the interface to different benchmarks, eggensperger_towards_2013 introduced HPOLib, which consolidates different benchmarks from the literature. In follow up work [eggensperger-neurips21], HPOBench containerized these benchmarks to increase reproducibility of benchmarking results. HPO-B [pineda-neurips21] provides tablularized benchmarks based on OpenML datasets [vanschoren-sigkdd13a]. In a similar vein, zimmer-ieee21 introduced LCBench, which provides full learning curves of neural network models trained on OpenML datasets. salinas-automl22 also provided a unified interface to tabular and surrogate benchmarks in Syne-Tune, using a fast data format to accelerate the table look-ups. Another suite for surrogate-based benchmarking is YAHPO [pfisterer-automl22], consisting of different scenarios, including combined hyperparameter optimization and algorithm selection.
Chapter 10 Conclusions
Hyperparameter optimization is a mature field with well-established algorithmic solutions and industry-level software frameworks available. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (model-based, multi-fidelity, population-based, and gradient-based) and briefly highlighted connections to closely related areas such neural architecture search and meta-learning. There remain open issues and interesting directions for improvement which we have not covered and briefly sketch below.
Response functions for unsupervised learning.
HPO requires a well defined response function to be minimized (see Section 2.2), which in turn requires a well defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning. A notable example is anomaly detection, an area rich of methods, results, and applications [pang:2021:deep_learning_anomaly]. The anomaly or novelty detection problem essentially consists of identifying the support of the distribution behind the data, so that low-probability points can be detected as outliers. For this problem, [ma:2023:need_unsupervised_outlier] survey a number of internal (i.e., not using any ground truth labels) approaches for assessing outlier detection models and conclude that “none of the existing and adapted strategies would be practically useful.”
A second important example is generative models. In this case the goal is to construct a model of the data distribution such that sampling from that model produces fake data points that are indistinguishable from real data. Again, the literature is rich in terms of models, ranging from early generative adversarial networks [DBLP:journals/corr/GoodfellowPMXWOCB14] to normalizing flows [pmlr-v37-rezende15] and more recently probabilistic diffusion models [ho:2020:denoising_diffusion_probabilistic]. In the particular case of images, popular metrics used to compare different generative models are the Inception Score [NIPS2016_8a3363ab] and the Fréchet inception distance (FID) [DBLP:conf/nips/HeuselRUNH17]. These metrics only capture one aspect of the generative process, namely fidelity. However, the problem of assessing the quality of a model is inherently more complicated and multidimensional: a perfectly faithful model that only produces samples that are too close to training data points is useless (implying an additional dimension related to diversity or creativity). There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports (of real and fake data) are estimated and compared (fake points off the support of real points are interpreted as false positives, and real points out of the support of fake points are interpreted as false negatives). The idea has sparkled a number of practically computable metrics that [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to optimize hyperparameters. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions would require reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionalty issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced in [yang:2020:evaluation_generative_models] for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data.
A notable exception in the realm of unsupervised learning is the case of obtaining disentangled representations, which is the problem of identifying a small number of explanatory factors behind observed data [bengio:2013:representation_learning_reviewa, locatello:2019:challenging_common_assumptions], such as shape, color and position for image classification problems. For this particular case, evaluation metrics have been developed based on informativeness, separability and inference, and interpretability [do:2019:theory_evaluation_metrics]. For this particular problem, a hyperparameter tuning algorithm was developed in [duan:2019:unsupervised_model_selection] using ranking scores. Still, the development of general purpose HPO algorithms for unsupervised learning remains an open problem.
Learning to (hyperparameter) optimize
In section 8.4 we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning an hyperparamter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and a response function of a (or several) standard learning algorithm. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g. validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on experience optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves, which we discussed in Chapter 5 when covering multi-fidelity methods; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modelling and learning to optimize, a sub-field of meta-learning, [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
The usage of rich meta-data, past experience, and leveraging similarities across standard learning algorithms may hold the promise of discovering HPO algorithm that are more efficient than classic approaches. However, significant challenges remain in developing truly general and robust meta-learning approaches for HPO, including handling the diversity of problem types, scalability to high-dimensional spaces, and transferability across vastly different domains and algorithms.
HPO and foundation models.
The emergence of foundation models [bommasani2021opportunities] as a new frontier for machine learning applications and research opens up a series of novel scenarios and challenges for hyperparameter optimization.
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become standard paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices, which are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparamter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (cf. Chapter 7) that optimizes for discrete “trigger” tokens, Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; cf. Section 5.2.2); Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches (with discrete gradient estimators) may prove useful in these directions,
In addition, the emergence of agents and agent networks where various LLMs interact also by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs (APIs). These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the the communication protocols and the interactions graph between agents. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.
[…]