### v2

# Hyperparameter Optimization in Machine Learning

Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determines the effectiveness of systems based on these technologies. Manual hyperparameter search is often unsatisfactory and becomes infeasible when the number of hyperparameters is large. Automating the search is an important step towards advancing, streamlining, and systematizing machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples, insights into the state-of-the-art, and numerous links to further reading. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model-, population-, and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
[…]
Chapter 1 Introduction
Over the last decade, machine learning has established itself as a key driver of innovation in science and industry. The digitalization of human activities and interactions has led to the generation of data at an unprecedented scale. These data could be stored ever more cheaply thanks to advances in micro-electronics, as well as the advent of cloud computing. Progress in hardware and low-power chip design also led to an exponential increase of computing capabilities of respectively cloud infrastructure, notably graphical processing units (GPUs), and devices on the edge, such as mobile phones. These two trends, cheaper and better hardware for storage and compute, combined with strong recent advances in algorithms and architectures for deep learning, lay the foundation for the successful application of statistical machine learning in practice and its adoption in artificial intelligence. It enabled learning algorithms to sift through vast amounts of data to discover and extract patterns that can then be used to make data-driven decisions without or with limited human intervention.
One notable illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed by machine learning experts at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of configuration parameters, called hyperparameters, were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization, a technique that sequentially predicts and then evaluates the performance of algorithms such as AlphaGo when it uses specific hyperparameters. Until a certain cost or time budget is exhausted, Bayesian optimization repeatedly trades off exploration and exploitation to determine what is the next most interesting set of hyperparameters to evaluate based on the system performances observed so far.
In practice, the generalization capabilities of any machine learning algorithm also depend on hyperparameters. Whereas conventional (shallow) algorithms such as support vector machines [shawe-taylor:2004:kernel_methods_pattern_analysis] need hand-engineered features to transform the raw data in a suitable format, deep neural networks learn in an “end-to-end” fashion by ingesting the raw data directly [lecun2015nature]. However, in both cases, the composition of the model (e.g., the choice of feature extractors or the neural network architecture) and the empirical loss minimizer depend on configuration parameters that are not adjusted by the algorithm itself. Ultimately, their choice contributes to determine how effectively the algorithm can learn a specific natural language understanding, computer vision, or speech recognition task.
1.1 A simple illustrative example
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as is commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression trained with stochastic gradient descent to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Table 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $\ell_{1}$ or $\ell_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Table 1.1(b). Interestingly, they showed that tuned logistic regression using a bag of words representation of the text performed on par with a convolutional neural network and only slightly worse than a sequential convolutional neural network. Hence, by carefully optimizing hyperparameters, the authors could achieve a result close to the state of the art with a much simpler model. We will come back to this example in more details in the next chapter.
(a) Hyperparameters.
(b) Accuracy.
Figure 1.1: Results on sentiment analysis reported in [yogotama2015-emnlp15]. Hyperparameters that were optimized and associated domains are listed on the left table. The feature extraction depends on the range of $n$-grams considered through $n_{min}$ and $n_{max}$. Stop words could be removed or not and words weighted according different weighting schemes. Two choices of regularization were considered and the amount of regularization was optimized. The convergence tolerance was also optimized. The binary classifiers compared are either SVM (support vector machine), LR (logistic regression), NN (multi-layer perceptron) or CNN (convolutional neural network). The authors found that when optimizing all these hyperparameters, simple logistic regression trained by stochastic gradient descent was performing as well as convolutional neural networks, which were the state of the art at the time of publication.
Hyperparameters play a critical role in machine learning as they do not only determine the generalization capabilities of trained models, but they may actually determine what constitutes the state-of-the-art. Indeed, results reported in an empirical study might look very differently for another choice of hyperparameters and so would be the conclusions drawn. Unfortunately, it is not uncommon for published results to omit specific hyperparameters used in the experiments, which may hinder reproducibility. Moreover, when experiments support the claim that a proposed method surpasses previous ones, the instability of results with respect to hyperparameters may cast doubts on these claims and raise questions about the current state-of-the-art [haibe-kains:2020:transparency].
The subject of this monograph is hyperparameter optimization (HPO). The availability of algorithmic tools to automate this task and their adoption will help machine learning researchers dealing with hyperparameters in a principled way. Importantly, these tools will also significantly help towards ensuring that new published advances are reproducible and thus unequivocally contribute to the field. Along with the experimental set-up, we argue that researchers would need to publish the HPO algorithm they used, the amount of resources it was allowed to consume [dodge:2019:show_your_work], and the hyperparameter search space that was considered [gundersen:2018:state_art]. Our objective is to provide a comprehensive overview of established techniques and recent advances in HPO. By doing so we hope we will help establish the practice of systematically using automatic HPO algorithms in research and industry, besides offering an introductory guide to understanding the fundamentals of this field.
1.2 What is hyperparameter optimization?
Let $A(\lambda)$ denote a learning algorithm configured by the hyperparameter vector $\lambda$ and consider a function $\mathscr{M}$ that evaluates its performance on a learning task. We call the composition $f(\lambda)=\mathscr{M}\circ A(\lambda)$ the response function. HPO consists in optimizing $f$ with respect to $\lambda$.
What distinguishes the HPO problem from conventional optimization problems, such as those arising in convex or combinatorial optimization, is its nested nature: to evaluate the response function $f$ at a specific point $\lambda$, it is required to run the learning algorithm $A$, which in turn has to solve another optimization problem to fit a model to the available data. In practice, running $A$ can be time-consuming. For example, training ResNet-50 from scratch on CIFAR-10 on an Amazon Elastic Cloud Compute (EC2) g4dn.xlarge instance takes more than 2 hours. [^1] A ResNet is a deep residual neural network commonly used in computer vision. ResNet-50 has 50 convolutional layers with skip connections (one for every stack of three layers). We refer the interested reader to the original paper by he2016deep.
1.3 Why is hyperparameter optimization difficult?
In general, computing $f$ might not only be time-consuming, but also expensive in terms of amount of resources used or cost. Its analytical form is typically either unknown or impractical to handle, it might have multiple local optima, it might exhibit wide flat regions, or it might be even non-smooth or non-continuous as a function of $\lambda$. As we will see in the following, for some classes of algorithms, we only require that $f$ is observable. Returning to the example of AlphaGo, this means that it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) and the probability of winning a game is known; what matters is that, for our specific choice of $\lambda$, we can observe if AlphaGo won or lost the game. Therefore, in general, one cannot apply conventional gradient-based optimization algorithms to solve HPO problems (but we will see that with additional assumption on $f$ this is, in fact, possible).
In addition, evaluations of $f$ can be stochastic, and the search space for $\lambda$ can itself be quite complex. Some hyperparameters could be continuous (e.g., learning rate), but others could be integral (e.g., number of layers in a deep neural network) or categorical to encode choices. Even more structure can emerge in the search space: conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units of the $j$-th layer of a neural network is only defined if the network has at least $j$ layers. If the domain to search over would be a small discrete set, one could carry out an exhaustive search by evaluating all the possible hyperparameter configurations in this set. Unfortunately, this is rarely the case in reality as the number of hyperparameter configurations to consider can be large or even uncountable.
All these challenges mean that, in practice, the focus of HPO is more often about finding “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than finding (global) optima. Hence, HPO often consists in determining a small set of good hyperparameters on which to evaluate $f$.
1.4 Historical remarks
Modern machine learning often considers a hierarchy of model families, indexed by model parameters at the lower level and hyperparameters at the higher level. Traditional model selection, such as Akaike’s Information Criterion (AIC) and the Bayesian Information Criterion (BIC), focuses on choosing the “model size” (i.e., a single integer-valued hyperparameter that determines the number of model parameters) that results in a good generalization [akaike:1974:new_look_statistical, schwarz1978estimating]. Interestingly, the origins of some recent HPO methods can be traced back to at least two decades before the introduction of AIC and BIC: the HPO problem in machine learning can be seen as a special case of optimal experimental design, an area that has a long tradition in statistics since the introduction of the response surface methodology [box:1951:experimental_attainment_optimum]; see myers16:_respon for a recent account.
Many ideas useful for tuning hyperparameters originated from a different class of problems: the configuration of heuristics and meta-heuristics in problem solving in Computer Science. The goal in these cases is not generalization but finding a solution with a small amount of computational resources. For example the LEX system [mitchell:1983:learning_experimentation_acquiringa] already embodied some fundamental ideas behind model-based approaches to HPO. LEX ran the to-be-configured algorithm to produce examples from which a generalizer learned to produce better heuristics. Model-based approaches have been subsequently applied to global optimization and propositional satisfiability [hutter:2009:experimental_investigation_modelbased], and to supervised learning [thornton2013auto]. These model-based approaches are based on or closely related to Bayesian optimization [mockus:1975:bayesian_methods_seeking, sacks:1989:design_analysis_computer, jones1998efficient]; see also the review by shahriari:2016:taking_human_out. Bayesian optimization is a technique at the core of current model-based HPO approaches in machine learning (see Section 4.2). Yet another class of approaches is based on the idea of racing, where candidate configurations are run in parallel and, as soon as sufficient statistical evidence is collected against a candidate, the run is stopped and the candidate replaced by a new one. This idea was initially formulated for model selection in machine learning [maron:1994:hoeffding_races_accelerating], and then applied to configuration of meta-heuristics [birattari:2002:racing_algorithm_configuring]. Finally, population-based algorithms also originated in the area of problem solving with seminal works of rechenberg65 and fogel:1965:intelligent_decisionmaking_simulation, boosted by the introduction of genetic inspired operators by holland75. While their application to optimization in machine learning has a long tradition, the use of population-based algorithms in HPO is more recent and preceded by the idea of gradient-based approaches [bengio_gradient-based_2000].
1.5 Outline of the monograph
In this chapter, we gave a broad overview of the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function $f(\lambda)$ based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require a much smaller number of hyperparameters (compared to grid or random search) to be evaluated to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, stochastic algorithms for global optimization that draw inspirations from natural phenomena. When applied to HPO, some techniques in this family such as population-based training maintain a set of predictors whose hyperparameters are updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how to adapt HPO to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency). Furthermore, we also discuss connections with with neural architecture search and meta-learning, and touch upon transfer learning of hyperparameters for foundation models. The two closing chapters are dedicated to an overview of HPO software frameworks (Chapter 9) and emerging topics and open questions in the field (Chapter 10).
1.6 Related surveys
The topic of this monograph has been surveyed in some other papers, briefly in [hutter_beyond_2015] and more extensively in  [bischl:2023:hyperparameter_optimization_foundations, yang:2020:hyperparameter_optimization_machine]. Arlot and Celisse arlot:2010:survey_crossvalidation_procedures surveyed the model selection problem from a complementary statistical perspective. Specific aspects have been reviewed in other review papers, in particular we would like to mention Bayesian optimization [Frazier:08, shahriari:2016:taking_human_out, wang:2023:recent_advances_bayesian], neural architecture search [ren:2021:comprehensive_survey_neurala, heuillet:2024:efficient_automation_neural], evalutionary algorithms [li:2023:survey_evolutionary_deep], and meta-learning [hospedales2020meta] HPO is a core topic in automated machine learning (AutoML). The subject is covered extensively in the book by Hutter et al. hutter2019automated and more recently in [he:2021:automl]
1.7 Notation
As usual, we denote by $\mathbb{R}$ and $\mathbb{N}$ the set of real and natural numbers, respectively. For $K\in\mathbb{N}^{+}$, we indicate by $[K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We reserve the Greek letter $\lambda$ to indicate hyperparameters and $f$ for the response function. We use $\mathbb{P}$ to indicate probability functions. Whenever confusion may arise, we further add a subscript like $\mathbb{P}_{x\sim\nu}$ to specify that the random variable $x$ evaluated by $\mathbb{P}$ follow the distribution $\nu$. We use $\mathbb{E}$ with the same convention to indicate the expected value. We mostly reserve the letter $p$ to indicate density or probability mass functions. We denote by $\mathcal{N}(\mu,\Sigma)$ the Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
Chapter 2 Hyperparameter Optimization
Here, we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We discuss the nested structure that characterizes an HPO problem and the archetypal settings encountered in machine learning, contrasting this to more general algorithm configuration problems. To fix ideas, we revisit in greater detail the example of the introduction and showcase different application scenarios from the literature. We conclude the chapter with an overview of practical desiderata for HPO algorithms that will guide our discussions in the remainder of the monograph.
2.1 Learning algorithms with hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm relies on in order to generalize outside of the training data. Formally, let:
  • $\mathscr{D}$ be a the collection of all finite datasets that the algorithm can process;
  • $\Lambda$ be an hyperparameter space, often though of as the Cartesian product of individual hyperparameters spaces;
  • $\mathcal{H}$ be a model space, typically indexed by a parameter vector $w$, that represents the class of models that the algorithm may output.
We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h_{\lambda}$, capable of accomplishing a given task:
$A:\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda).$ (2.1)
We invariably call a $\lambda\in\Lambda$ a hyperparameter vector or a hyperparameter configuration. By writing $h_{\lambda}=A(\mathcal{D},\lambda)$ in Eq. (2.1) we emphasize the dependence of the model returned by the algorithm on the hyperparameter vector $\lambda$, but omit the dependence on the training data $\mathcal{D}$. This choice is consistent with the fact that often one considers the dataset fixed and “given” when setting up an HPO problem.
2.1.1 Logistic regression for sentiment analysis: an HPO perspective
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Let us denote the data by $\mathscr{D}:=\bigcup_{N\in\mathbb{N}^{+}}(\mathcal{X}\times\mathcal{Y})^{N}$, where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ is the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda_{1}\times\dots\times\Lambda_{7}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}:\mathcal{X}\to\mathbb{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
=$\mathcal{H}=\left\{h_{\lambda}:\mathcal{X}\to\mathcal{Y}\,:\,h_{\lambda}(x)=\textbf{1}_{\mathbb{R}^{+}}(w^{\intercal}\Phi_{\lambda}(x)),\;w\in\mathbb{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
where $\mathbf{1}_{S}$ is the indicator function: $\mathbf{1}_{S}(y)=1$ if $y\in S$ and 00 otherwise. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ vary in their respective spaces $\Lambda_{1}$, $\Lambda_{2}$, $\Lambda_{3}$ and $\Lambda_{4}$.
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the $n$-grams considered, the more input features and, consequently, parameters the model will have. This range is controlled by the hyperparameters $\lambda_{1}$ and $\lambda_{2}$. Unigrams alone ($\lambda_{1}=\lambda_{2}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer $n$-grams typically requires increasingly larger corpora to avoid overfitting, since the number of possible features (sequences of words of length $n$) grows exponentially with $n$.
The remaining hyperparameters, $\lambda_{5},\lambda_{6}$ and $\lambda_{7}$, do not directly affect the model space, but rather determine the way in which the optimization is carried out. Specifically, the algorithm $A$ outputs the model whose parameter vector $w$ minimizes the training objective $\mathcal{J}_{\lambda}:\mathbb{R}^{n_{\lambda}}\times\mathscr{D}\to\mathbb{R}$ defined as follows:
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log\mathbb{P}\left(y|\Phi_{\lambda}(x);w\right)+\lambda_{6}\Omega_{\lambda_{5}}(w),$ (2.3)
where $\Omega_{\lambda_{5}}(w)$ is either the $L^{1}$ or the $L^{2}$ norm depending on the value of $\lambda_{5}$, and the probability that the model assigns a document to express a positive sentiment, encoded by $y=1$, is given by
$\mathbb{P}\left(1|\Phi_{\lambda}(x);w\right)=\frac{1}{1+e^{-w^{\intercal}\Phi_{\lambda}(x)}}.$
Hyperparameter $\lambda_{5}$ exposes a binary choices between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values. Figure 2.1 offers a geometric interpretation of these regularization schemes. The hyperparameter $\lambda_{6}$ controls the strength of such regularization; if, for instance, $\lambda_{5}=L^{1}$, one can expect the number of 00’s in the solution vector to be proportional to the value of $\lambda_{6}$.
Figure 2.1: Geometrical representations of the $L^{2}$ (left) and $L^{1}$ (right) regularizers. $w^{*}_{0}$ represents the minimizer of the training loss $\mathcal{J}_{\lambda}$ when the regularization coefficient $\lambda_{6}$ is 00 (i.e. no regularization). $w^{*}_{\rho_{1}}$ and $w^{*}_{\rho_{2}}$ are, instead, the minimizers of the regularized losses for two different values of $\lambda_{6}=\rho_{i}$ for $i\in\{1,2\}$, with $0<\rho_{1}<\rho_{2}$. These corresponds to the projections of $w^{*}_{0}$ onto $L^{2}$ and $L^{1}$ balls with varying radii that are inversely proportional to $\lambda_{6}$.
Finally, $\lambda_{7}$ controls the convergence tolerance of the numerical solver used to minimize the objective (2.3).
Although logistic regression is a relatively simple algorithm, its practical application already exposes several hyperparameters. Other algorithms may require many more hyperparameters. This happens for example in deep learning, where feature extraction is internalized in the underlying models, and the optimization problems become more challenging.
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda_{6}$ in our example could in principle take any non-negative real value: it does not need to be limited to the interval $[10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in (2.3) instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
Hence, by exposing a (relatively) larger number of hyperparameters of the learning algorithm, such as parameters associated with the feature extractors or parameters associated with the optimizer or the regularizer, one can substantially expand the algorithm’s model space $\mathcal{H}$. These additional degrees of freedom may in turn raise the odds of finding a better model for the task at hand, at the price of having to search over a possibly much larger hyperparameter space.
2.2 The hyperparameter optimization problem
Hyperparameter configurations need to be evaluated according to a given performance metric, that we denote by $\mathscr{M}$. This metric takes as input a trained model and a validation set, and returns a scalar value, such as the validation error.  [^1] In Section 8.1 we discuss the case where $\mathscr{M}$ returns multiple scalars. The objective of an HPO problem, the response function, is then constructed from the performance metric. Specifically, given a learning algorithm $A:\mathscr{D}\times\Lambda\to\mathcal{H}$, a metric :×$\mathscr{M}:\mathcal{H}\times\mathscr{D}\to\mathds{R}$, a training set $\mathcal{D}$ and a validation set $\mathcal{V}$, the hyperparameter optimization problem is defined as follows:
$\min_{\lambda\in\Lambda}f(\lambda)$ $=$ $\mathscr{M}\left(h_{\lambda},\mathcal{V}\right)$ (2.4)
such $\mbox{such that}\quad h_{\lambda}$ $=$ $A(\mathcal{D},\lambda).$ (2.5)
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until convergence. Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of out-of-the-box continuous algorithms for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 where we discuss gradient-based HPO techniques that explicitly leverage this structure. A key observation is that we cannot collapse the optimization problem into one level, where we jointly minimize both model parameters and hyperparameters. For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda_{6}=10^{-5}$ and $\lambda_{7}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.
2.2.1 Estimating generalization
Most of the material and techniques for HPO we covered in this monograph may be used regardless of the specific nature and meaning of $\mathscr{M}$ and, hence, of $f$. However, since many hyperparameters control the inductive bias of their respective learning algorithms, the archetypal setting consists in a response function that estimates the generalization of the learning algorithm’s output models.
Figure 2.2: Three examples of response functions. From left to right the underlying algorithms are: ridge regression, two-layers neural network regression, Monte Carlo tree search [chen_learning_2017].
Informally, a hypothesis generalizes if it “behaves sensibly” on data outside $\mathcal{D}$.
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g., in generative modelling), but it is easy in the case of supervised learning, where we assume to have access to a task loss :×$\mathcal{L}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\mathds{R}$. We can define the generalization error of $h_{\lambda}$ as the expected loss:
$\mathcal{E}(h_{\lambda})\doteq\mathbb{E}_{(x,y)\sim p}\left[\mathcal{L}(h_{\lambda},(x,y))\right],$ (2.6)
where $p$ is the unknown task distribution from which the training dataset $\mathcal{D}$ was drawn independently. Widely used metrics such as classification error or mean-squared-error for regression can be interpreted as empirical estimates of (2.6):
$\mathscr{M}(h_{\lambda},\mathcal{V})=\frac{1}{|\mathcal{V}|}\sum_{(x,y)\in\mathcal{V}}\mathcal{L}(h_{\lambda},(x,y)),$ (2.7)
where $\mathcal{V}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. For example, the sentiment analysis problem addressed in [yogotama2015-emnlp15] is an instance of supervised binary classification and the task loss is the 00-$1$ error, namely $\mathcal{L}(h_{\lambda},(x,y))=\mathbf{1}\{h_{\lambda}(x)\neq y\}$. Note that in many cases, the internal structure of the learning algorithm, $A(\mathcal{D},\lambda)$, also requires the minimization of a loss function $\mathcal{L}^{\prime}$ with respect to the model’s parameters, which may or may not coincide [^2] For example, in the case of classification, the 0-1 loss is intractable to minimize even for linear classifiers and we usually resort to a convex surrogate [nguyen:2009:surrogate_loss_functions] such as the cross-entropy. with $\mathcal{L}$.
Figure 2.2 shows some examples of other response functions. The leftmost plot concerns a simple ridge regression algorithm to regress median housing prices using the Boston dataset [harrison1978hedonic]. The scalar hyperparameter controls the $L^{2}$ regularization and $\mathcal{L}(h_{\lambda},(x,y))=\|h_{\lambda}(x)-y\|^{2}$. The plot shows the mean loss computed on the training and validation data, respectively. The center plot represent the response function of a two-layer neural network on the same learning scenario. Here the two hyperparameters are an $L^{2}$ regularization coefficient (real-valued, vertical axis) and the number of units in the hidden layer (integer, horizontal axis). The rightmost plot, reproduced from [chen:2018:bayesian_optimization_alphago] represents instead the winning probability of a Go agent introduced in Chapter 1, and somewhat deviates from the standard setting we just discussed.
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of (2.6); e.g., underestimation when the model is under-regularized. This is illustrated in Figure 2.2, left. In order to obtain an unbiased estimate of the expected loss (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$. Otherwise, one carries over unwanted dependencies between training and validation.
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$. 333 Note that a third test set would be needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of (2.6) at the expense of increased computational cost.
2.2.2 Other types of response functions
One may also be interested in optimizing hyperparameters with respect to performance measures not necessarily linked to generalization. For instance, $\mathscr{M}$ could measure the runtime of $A$ when, e.g., tuning optimization hyperparameters [xu2008satzilla, hutter2009paramils], or the inference latency or the memory requirements of $h_{\lambda}$ when looking to deploy models for devices on the edge [benmeziane2021comprehensive]. chen:2018:bayesian_optimization_alphago take $\mathscr{M}$ to be an estimate of the winning rate of the AlphaGo agent that plays with hyperparameters $\lambda$ versus another AlphaGo agent that plays with a fixed baseline hyperparameter configuration $\lambda_{0}$. Figure 2.2 (right) plots the response function (i.e. the winning rate) along two hyperparameter axes.
Furthermore, one may be interested in optimizing the hyperparameters of a learning algorithm simultaneously with respect to multiple measures (e.g., accuracy, inference latency and memory footprint) or in ensuring that $h_{\lambda}$ satisfies some constraints regarding one or multiple measures. These scenarios give rise to multi-objective and constrained hyperparameter optimization, advanced topics which we will discuss in Chapter 8.
2.2.3 General algorithm configuration
Several algorithms, not only those employed in machine learning, require configuration constants that may affect both the quality of the solutions and the computational resources used to find solutions on specific instances [schede:2022:survey_methods_automated]. While in HPO the emphasis is on the former aspect, in general algorithm configuration (AC) the focus is more typically on compute. A typical problem suitable for AC is propositional satisfiability (SAT), where proper choices of parameters can improve solving time significantly [hutter:2009:paramils_automatic_algorithm, hutter_sequential_2011]. Additional interesting examples can be found in the area of discrete optimization, where several parameters may be used to configure metaheuristics [Birattari2009TuningM]. Entering into details would be out of scope for this survey. Still, we would like to point out a striking similarity between AC and HPO. In both cases, optimal (hyper) parameters depend on some unknown probability distribution, such as the data distribution for HPO, or the distribution of future instances of a certain problem in AC. This requires resorting to data (such as a validation set or a set of problem instances) to formulate a proxy empirical optimization problem.
2.3 Hyperparameter optimization in the literature
Here below we report a few notable milestones and examples of “success stories” of HPO in the literature.
  • bergstra_random_2012, later extended in [bergstra_making_2013], brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, where widespread approaches to HPO. We cover grid, random and more advanced quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They improved the manual found hyperparameter by a human expert which marked state-of-the-art at this point by reducing the test error by $3\%$. This papers sparked interest in Bayesian optimization for HPO. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
  • As we mentioned in the introduction, chen:2018:bayesian_optimization_alphago used Gaussian-process-based Bayesian optimization to optimize the hyperparameter of the Monte-Carlo Tree-Search of Alpha GO. This improved the win-rate of Alpha GO from $50\%$ to $66.5\%$ in self-play games.
  • zoph2016neural parameterized the architecture of a convolutional and recurrent neural networks to discover new architectures. To sample new architectures they trained a recurrent neural network controller based on reinforcement learning. This work marked the beginning of a sub-field of research dubbed neural architecture search (NAS), which we discuss in Section 8.3.
  • jamieson2016non and later li2017hyperband introduced bandit-based approaches to hyperparameter optimization. These methods draw inspiration from the multi-armed bandit literature, particularly the successive halving algorithm [karnin-icml13]. They adaptively allocate resources to promising configurations while terminating unpromising ones early, emphasizing the concepts of budget and efficient compute allocation. These works have been particularly influential in the development of multi-fidelity optimization methods, which we discuss in Chapter 5.
2.4 Practical desiderata for HPO algorithms
The nested optimization formulation (2.4) abstracts away some aspects which are practically relevant. Together, these characteristics set aside hyperparameter from other optimization scenarios. For this reason one should weigh an HPO algorithm alongside multiple directions. We summarize these desiderata in Table 2.1.
Characteristic Brief Description
Sample efficiency How many evaluations of $f$ does the algorithm need to find satisfactory configurations?
Runtime efficiency How long does the algorithm take to find satisfactory configurations?
Parallelizability: How efficiently does the algorithm exploit parallel computational resources if:
   | Synchronous •  all parallel evaluations need to start at the same time?
   | Asynchronous •  evaluations can be started any time a resource becomes available?
Multi-Fidelity Support Can the algorithm take advantage of cheaper approximate evaluations of $f$?
Applicability Does the algorithm have restriction on the type of search space $\Lambda$?
Accessibility Which type of access does the method need to the underlying learning algorithm and/or training data?
Ease of use How difficult is to implement and tune the algorithm? What expertise is required to use it effectively?
Table 2.1: List of desirable properties of an HPO algorithm.
Efficiency and budget.
The response function $f(\lambda)$ is typically stochastic, depending on random effects in the computation of $A(\mathcal{D},\lambda)$ (e.g., weight initialization, data ordering, e.g., when performing stochastic gradient descent, etc.). It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ – expressed in time, number of function evaluations, monetary costs, etc. – rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a sub-optimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
Multi-fidelity.
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access $A_{b}(\mathcal{D},\lambda)$, that is, the model after $b$ epochs of training have elapsed, for free for all $b\leq 64$. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where, in this case, $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5. One aspect that is worth noting when choosing an HPO algorithm is if it can leverage low-fidelity evaluation, and how the quality of such evaluation may impact the final result.
Synchronous and asynchronous parallelization.
We can trade off time efficiency against computation cost by running several evaluations of $f$ in parallel on (say) $W$ workers. For a synchronous system, all $W$ workers start new jobs at the same time. While this reduces wall-clock time, it often leads to idle time when workers need to wait for the slowest job to finish. In an asynchronous system, each worker can be assigned a new job independently from the status of the others. However, HPO decision-making for parallel execution systems comes with additional challenges. In the synchronous case, a batch of $W$ new configurations has to be suggested in one go. Configurations in a batch should be “diverse”, in that they reveal non-redundant information about the optimum [Ginsbourger:10, Desautels:14, Gonzales:16, Wilson2018]. While only a single new configuration needs to be proposed in the asynchronous case, up to $W-1$ responses of running evaluations are not yet available [snoek_practical_2012]. Not taking such pending evaluations into account runs the risk of increased redundancy in the proposed configurations. HPO can also be distributed across several computational nodes that may reside in networked computers. In this case, algorithms should also be parsimonious in terms of communication to reduce overhead as much as possible.
Applicability and accessibility.
The applicability and accessibility of an HPO algorithm refer to the requirements it may have on the type of problem or the level of access needed to the underlying learning algorithm. Some algorithms may be restricted to certain types of search spaces, e.g., continuous vs integer or categorical domains, or a mix of these. Furthermore, some algorithm can exploit conditionality in the search space, while others cannot. Some HPO methods like gradient-based techniques (Chapter 7) presuppose a white-box access to the underlying learning algorithm in terms of computational structure and/or data processing, while others assume the learning algorithm to be a black box. In between, there are a class of algorithms that require a gray-box access, like the ability to start and stop the execution of the learning algorithm and be able to observe intermediate statistics. Although HPO has been traditionally described and tackled as a black-box optimization problem, more restrictive assumptions may be met in specific cases and HPO algorithm that exploit these assumptions typically show efficiency gains.
Ease of use
The ease of use desideratum refers to the complexity involved in implementing, configuring, and effectively using an HPO algorithm. Some methods may be conceptually simple but require careful tuning of their hyperparameters or auxiliary components. Others may be more complex to implement from scratch or require deep expertise in the underlying techniques (e.g., Bayesian and multi-fidelity optimization) to use them effectively.
2.5 Tackling the search manually
In the next chapters we will introduce the main algorithms for HPO. But in this section we take a few sentences to state the obvious. Searching for the best hyperparameters by hand can be alluring: getting started requires no new frameworks or assumptions, and one may be under the impression that the search could proceed organically, driven by hunches.
Yet manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Manual search also suffers from poor reproducibility, as it is hard to track the steps an user followed to reach some final configuration.
Chapter 3 Elementary Algorithms
In this chapter, we present simple non-adaptive, model-free and “memoryless” approaches to HPO, in that subsequent runs are not informed by any previous search history. These methods correspond to basic search schemes in global optimization that are conceptually simple, widely applicable, (mostly) easy to implement and straightforward to parallelize. Furthermore, these methods may work as initialization schemes for more advanced techniques like Bayesian optimization, as we shall see in later chapters.
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favour of a fixed, simple distribution over the search space, drawing configurations from such distribution until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offers benefits over random search, although to date we are not aware of any theoretical result supporting this.
3.1 Grid Search
Perhaps the first formal description of grid search (GS) dates back to the 1930’s, appearing in a foundational work by fisher1936design on experimental design, under the name of factorial design. GS consists in evaluating $f(\lambda)$ over a predefined grid of hyperparameter configurations $\upsilon\subset\Lambda$. For each entry of $\lambda\in\Lambda=\Lambda_{1}\times\dots\Lambda_{m}$, the user defines a list of values:
$\bar{\lambda}^{i}=\{\bar{\lambda}^{i}_{j}\}_{j=1}^{n_{i}},\quad\bar{\lambda}^{i}_{j}\in\Lambda_{i},\quad\text{for }\;i\in[m].$
The search grid is then given by all the combinations on the Cartesian product generated by these $m$ sets, namely $|\upsilon|=\prod_{i=1}^{m}n_{i}$. The learning algorithm can be executed sequentially or in parallel on the resulting set of points. The hyperparameter configuration that minimizes the objective function $f$ over $\upsilon$ is returned as $\hat{\lambda}$; that is
$\hat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\upsilon}f(\lambda).$ (3.1)
GS is a simple memoryless method that can be used to explore low-dimensional hyperparameter spaces. It requires very little implementation effort and parallelization is straightforward. The number of total trial generated by GS, however, grows linearly in the size of the trial sets $\bar{\lambda}^{i}$ and is exponential in $m$, making the method impractical for even moderate values of $m$. In addition, while Boolean and categorical hyperparameters naturally lend themselves to an evaluation on a predefined grid, unbounded integer and real-valued hyperparameters require manually setting bounds and discretization.
Extensive experimental practice has shown that some continuous hyperparameters, such as the learning rate, benefit from discretization on a logarithmic scale [bengio2012practical]. While requiring less expert knowledge than manual search, GS is very sensitive to the definition of the grid. To reduce the computational overhead, GS can be paired with manual search: the user alternates GS steps with adjustments and refinement of the grid values, e.g., using a finer grid “centered” around the best found configuration, or moving the search bounds of continuous hyperparameters, until a satisfactory result is found.
The top row of Figure 3.1 depicts equally distanced grid “samples” with 4, 9 and 16 points, on the unit square.
3.2 Random Search
Perhaps conceptually the simplest procedure for HPO, random search (RS) appears already considered in the 50’ as a method for global optimization [brooks1958discussion]. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$ and it is readily parallelized. The incumbent solution after $K$ trials is given by
$\hat{\lambda}_{K}=\operatorname*{arg\,min}_{(\lambda_{1},\dots\lambda_{k})\sim\upsilon^{K}}f(\lambda_{i}),$ (3.2)
where the trial points $\lambda_{i}$ are drawn i.i.d. from $\upsilon.$ While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or, much more frequently in HPO, after a budget is exhausted (see Section 2.4). For instance, one may run RS for 12 hours.
Even if in principle $\upsilon$ could be any distribution, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions (uniform or log-uniform distributions) that factorize independently along the configuration components, i.e. $\upsilon=\prod_{j=1}^{m}\upsilon_{i}$. For instance, if $\lambda$ contains, in order, a categorical choice between $K$ different neural architectures, the learning rate for stochastic gradient descent and a global dropout factor for regularization then $\upsilon$ could be set as
$\upsilon=\mathcal{U}\{[k]\}\times\mathrm{Log}\,\mathcal{U}(10^{-5},10^{-1})\times\mathcal{U}(0,0.5).$
where $\mathcal{U}\{\cdot\}$ and $\mathcal{U}(\cdot)$ are the discrete and continuous uniform distributions and $\mathrm{Log}\,\mathcal{U}(\cdot)$ is the log-uniform (or reciprocal) distribution on the indicated intervals. Using a distribution rather than a grid allows for a more natural treatment of a larger class of hyperparameters: RS does not require discretization of continuous spaces and also in principle allows also for distributions with non-compact supports.
The second row of Figure 3.1 shows uniform random samples with 4, 9 and 16 points on the unit square.
3.2.1 On the efficiency of random search (compared to grid search).
In an influential article, bergstra_random_2012 advocate for using RS as the “default” baseline method for HPO instead of manual or grid search. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter $\lambda=(\lambda_{1},\lambda_{2})$ it could be that $f(\lambda)\approx g(\lambda_{1})$, for some $g:\Lambda_{1}\to\mathbb{R}$.
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to a positive outcome of the experiment, i.e., where $f$ is sufficiently small. Then, we can assess the odds that a run of RS “lands” at least one configuration in the good region by comparing the volume of such region with that of the full search space under random sampling [brooks1958discussion]. We can model the probability of success of random search after $K$ draws as follows:
$\mathbb{P}_{(\lambda_{1},\dots\lambda_{K})\sim\upsilon^{K}}\left(\hat{\lambda}_{K}\in\Lambda^{*}\right)=1-\left(1-\mathbb{P}_{\lambda\sim\upsilon}(\lambda\in\Lambda_{*})\right)^{K},$ (3.3)
where $\mathbb{P}_{\upsilon}$ denotes the probability of sampling from $\upsilon$, typically uniform or log-uniform. Eq. (3.3) says that the minimum number of trials needed to assure a positive outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda_{*}$ under $\upsilon$. This can be independent of the search space dimension if the response function has a low effective dimensionality. This argument may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures.
For a general search space, we should expect $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m-m^{\prime}}$ is the “good region”. The $m^{\prime}$-dimensional 1/2-hypercube represents the important hyperparameters, while the $m-m^{\prime}$-dimensional unit hypercube represents the unimportant ones, as these coordinates do not matter in hitting $\Lambda^{*}$. If $\upsilon=\mathcal{U}(\Lambda)$ is the uniform distribution over $\Lambda$, then $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})=2^{-m^{\prime}}$. Consequently, the number of trials needed to obtain a positive outcome with at least $95\%$ probability is
$K\geq\log(0.05)/\log(1-2^{-m^{\prime}}).$
For $m^{\prime}=3$ one already needs $K\geq 23$ draws to achieve a $95\%$ probability of success. This illustrates that the number of trials needed grows exponentially in $m^{\prime}$. At the same time, this also shows that if $m^{\prime}\ll m$ RS is much more efficient compared to GS.
3.2.2 Unlucky runs
One potential disadvantage of RS is that its randomized nature means that “unlucky runs” may end up with entire regions of $\Lambda$ being under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed.
Let us illustrate this phenomenon. Continuing with our illustrative examples, say we divide the unit hypercube in $2^{m}$ 1/2 hypercubes $I_{j}$, for $j\in[2^{m}]$. We wish to have at least one configuration per hypercube (i.e., anywhere inside the 1/2 hypercube). While GS would need exactly $2^{m}$ points to achieve this, the probability under uniform sampling that RS hits all the hypercubes in $2^{m}$ draws is
$\mathbb{P}_{(\lambda_{1},\dots,\lambda_{2^{m}})\sim\upsilon^{2^{m}}}(\forall\,j\in[2^{m}],\;\;\exists\,k\in[2^{m}]\,\text{ s.t. }\,\lambda_{k}\in I_{j})=\prod_{j=1}^{2^{m}-1}(1-j/2^{m}).$
For example, for $m=3$ such probability is already $3/32$,. This means that, in expectation, RS with $8$ draws covers all the hypercubes only once every ten runs.
3.3 Quasi-random search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{k\to\infty}\min_{i\in[k]}f(\lambda_{i})=f(\lambda^{*})$. RS with a uniform distribution is consistent, while GS is not. However, unlike GS, random search for any fixed budget may generate a set of configurations that misses large portions of the search space. Intuitively, good sequences should cover evenly the search space. Also their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality – a property that GS misses. Quasi-random search methods seek to combine the advantages of grid and random search, with configuration-drawing strategies that combine randomness with deterministic steps.
Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks; see also [tuningplaybookgithub, Chapter 7] The resulting methods are conceptually very similar to both GS and RS, being highly parallelizable, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{k}$ are generated according to more complex procedures. We review two possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. Note that this is typically not at all an hard limitation: starting with samples on the unit interval, one can easily derive samples for a variety of distributions with appropriate reparameterizations [see, e.g. devroye2006nonuniform].
3.3.1 Latin hypercube sampling (LHS)
LHS [mckay1979comparison] directly combines grid and random search, in that it divides the search space into a grid, but samples configurations at random within regions of the grid.
LHS works as follows: choose a sequence length $n>1$, which determines the number of hyperparameters generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}:[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
$\lambda_{k}\sim\mathcal{U}(H_{k}),\;\text{ where }\;H_{k}=I_{\sigma_{1}(k)}\times I_{\sigma_{2}(k)}\times\cdots\times I_{\sigma_{m}(k)},$ (3.4)
where $\mathcal{U}(H_{k})$ is the uniform distribution on $H_{k}$. This generates $n$ trial points; one may then generate further sequences of $n$ configurations by repeating the process. Importantly, Latin hypercube sampling ensures that the projection on each axis of the $n$-long sequences places exactly one sample per interval.
The second last row of Figure 3.1 shows Latin hypercube samples on the unit square with 4, 9 and 16 points; namely, for $n\in\{2,3,4\}$.
3.3.2 Sobol’ sequences
Sobol’ sequences [sobol1979systematic] are a collection of separate one-dimensional sequences, one for each axis: when constructing a $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$ – representing the iterate – let $k=\sum_{l}a_{l}(k)2^{j}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 00’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
$[\lambda_{k}]_{j}=\sum_{l}\frac{[C_{j}\cdot a(k-1)]_{l}}{2^{l+1}}$
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., the additions are modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to [Lemieux:1338331, Chapter 5] for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences by adding random shifts, i.e. $\lambda^{\prime}_{k}=\lambda_{k}+u$ (modulo the unit hypercube) with $u\sim\mathcal{U}([0,1]$ and by “scrambling” the binary representation vector $a(k)$, that is, randomly permuting its entries [owen1995randomly].
The bottom row of Figure 3.1 plots 4, 9 and 16 points Sobol sequences with random shift and scrambling on the unit square.
3.3.3 Other schemes
bousquet2017critical experimented also with Halton and Hammersley sequences, finding the latter (with scrambling and random shifting) to be overall robust and well performing on the tested deep learning scenarios. Other quasi-random low-discrepancy sequences are presented in [Lemieux:1338331]. LHS, Sobol’, Halton and Hammerley sampling are implemented in the stats.qmc module of SciPy.
Figure 3.1: Non-adaptive sampling methods on the unit square for different sample sizes: 4 (left), 9 (center) and 16 (right) points. Each row represents a different schema. Grid and Latin hypercube sampling (LHS) may require re-sampling every time we increase the sample size. Instead, for random and Sobol we can keep on adding points, “reusing” previous samples. For each plot, we also show projections of the samples onto the two axes. We can see how grid samples cover much fewer points on the projections w.r.t. the other sampling methods. We can also observe the “unlucky run” phenomenon of (uniform) random sampling, where after nine draws the upper-right quadrant is still unvisited. LHS and Sobol (with random shifts and scrambling) provide more even coverage of the space and its unidimensional projections.
Chapter 4 Model-based Methods
The techniques described in the previous chapter do not make any attempt to model the response function $f(\lambda)$: in other words, evaluations of $f$ on previously selected configurations do not influence the search. In contrast, model-based (MB) techniques construct an explicit surrogate model for $f(\lambda)$ – or for closely related quantities – based on the information collected from previous evaluations. The surrogate model is a first central component of model-based HPO methods. We will denote such model by
$\hat{f}:\Lambda\times\mathcal{X}\to\mathcal{Y},$ (4.1)
where $\mathcal{X}$ may represent additional inputs (e.g., the performance of partial executions of $A$, a baseline or the value of the incumbent solution) and $\mathcal{Y}$ is an appropriate output space. For instance, $\mathcal{Y}$ could be (a distribution over) $\mathbb{R}^{+}$ if $\hat{f}$ models directly $f$, or the interval $[0,1]$ if $\hat{f}$ predicts the probability that a configuration improves over a given baseline.
The landscape of model-based methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we cover in Section 4.2. In BO, the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation; implementations include Gaussian processes (Section 4.2.2) and tree Parzen estimators (Section 4.2.5). Regardless of the particular implementation, one critical feature of surrogate models is that they should be much cheaper to evaluate than $f$ itself. BO methods do not locally minimize the surrogate function to select the next configuration to evaluate, but rather mediate it through a so-called acquisition function (AF), a second central component of a BO method and defined as a mapping:
$\alpha:\Lambda\times\hat{\mathcal{F}}\times\mathcal{X}\to\mathbb{R}$ (4.2)
where $\hat{\mathcal{F}}$ is an appropriate space of surrogate models, and, again, $\mathcal{X}$ collects auxiliary inputs, that is used to select the next configuration to evaluate $f$ on. Typically, $\alpha$ encodes a trade-off between exploration and exploitation:
  • exploration seeks to evaluate $f$ on points with high uncertainty, with the primary aim to improve the surrogate model;
  • exploitation seeks to leverage knowledge gathered so far to improve on the incumbent solution.
We discuss acquisition functions in Section 4.2.3.
We briefly touch upon other approaches to model-based optimization that do not directly fall in the BO category in Section 4.3. We conclude the chapter with a practical example that contrasts BO to random search 4.4 and an overall discussion in Section 4.5. In particular, we discuss parallelizability of model-based methods, a challenging practical bottleneck of MB-HPO methods (Section 4.5.1).
4.1 Model-based HPO: a general schema
Algorithm 1 Model-based-HPO
1:Response function $f$
2:Warm-up dataset $\mathcal{C}_{0}=\{(\lambda_{j},f(\lambda_{j}))\}$
3:Class of surrogate functions $\hat{\mathcal{F}}$ and fitting routine
4:Acquisition function $\alpha$
5:$\beta\leftarrow\min_{j}f(\lambda_{j})$triangleright Initialize incumbent
6:$\mathcal{C}_{1}\leftarrow\mathcal{C}_{0}$
7:for $k\leftarrow 1$ to $T$ do
8:     Fit $\hat{f}_{k}$ to $\mathcal{C}_{k}$triangleright Update step ($\pi_{2}$)
9:     $\lambda_{k}\leftarrow\operatorname*{arg\,max}_{\lambda^{\prime}}\alpha(\lambda^{\prime}|\hat{f}_{k-1},\beta)$triangleright Decision step ($\pi_{1}$)
10:     Evaluate $f(\lambda_{k})$triangleright Expensive step
11:     if $f(\lambda_{k})\leq\beta)$ then
12:         $\beta\leftarrow f(\lambda_{k})$
13:     end if
14:     $\mathcal{C}_{k+1}\leftarrow\mathcal{C}_{k}\cup(\lambda,f(\lambda))$
15:end for
The vast majority of model-based methods is, in principle, sequential: the surrogate model suggests the next point to evaluate while being simultaneously adapted online as new data from the execution of the learning algorithm $A$ is collected and processed.
It can be useful to think about a MB-HPO method as an iterative process over both the hyperparameter space $\Lambda$ and the surrogate function space $\hat{\mathcal{F}}$. To represent this process, we introduce two mappings $\pi_{1}:\mathcal{F}\times\mathcal{X}\to\Lambda$ and $\pi_{2}:(\Lambda\times\mathbb{R})^{*}\times\hat{\mathcal{F}}\times\mathcal{X}\to\hat{\mathcal{F}}$, one corresponding to the iteration in the hyperparameter space, the other to the update of the surrogate model. Then, we can describe the iterates of a MB-HPO method as follows:
$\left(\begin{array}[]{c}\lambda_{k+1}\\ \hat{f}_{k+1}\end{array}\right)=\left(\begin{array}[]{c}\pi_{1}(\hat{f}_{k},\upsilon_{k})\\ \pi_{2}\left(\mathcal{C}_{k},\hat{f}_{k},\upsilon_{k}\right)\end{array}\right),$ (4.3)
where the variables $\nu_{k}$ represent auxiliary inputs and $\mathcal{C}_{k}$ a set of paris of configurations and response function evaluations. The surrogate update mapping $\pi_{2}$ may often entail an online update of $\hat{f}$, therefore depending only on the last collected point $(\lambda_{k},f(\lambda_{k}))$ and the current surrogate model $\hat{f}_{k}$.
Typically $\hat{f}_{0}$ is initialized by fitting it to several evaluations obtained in a “warm-up stage”, e.g., where configuration are generated by random search or Latin hypercube sampling. Then, the points $\mathcal{C}_{k}=\mathcal{C}_{0}\cup\{(\lambda_{j},f(\lambda_{j}))\}_{j=1}^{k}$ that comprise past configurations and evaluations of the response function (including warm-up evaluations) can be though of as a dataset to fit the surrogate model. We summarize the typical steps of a model-based method in Algorithm 1, where the decision map $\pi_{i}$ is implemented by the optimization of an AF $\alpha$. The algorithm generates a sequence of $T$ configurations. As in the previous chapter, we denote by $\hat{\lambda}_{k}$ the best configuration found after $k$ steps.
4.2 Bayesian optimization
BO has long been used as a method to globally optimize black-box functions, that is, functions for which we may only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda)$. It dates back at least to the seventies with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. While limited to small-to-medium dimensional settings, BO is an effective approach to optimize (highly) multi-modal, non-convex and possibly expensive-to-compute functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give a brief overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for an intuitive introduction of the mathematical aspects of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning.
In Bayesian optimization, $\hat{y}$ is a stochastic process rather than a function and $\hat{y}(\lambda)$ is a random variable. Here randomness arises because of the epistemic uncertainty of the surrogate model (due to the small size of $\mathcal{D}$) about the response at $\lambda$. This surrogate model is fitted to available observations by Bayesian conditioning, so that afterwards it represents the Bayesian posterior distribution, which is used to decide the next value to probe. Such decision is most often expressed as the maximization of an acquisition function, as we introduced before. The AF balances exploration — collecting information to improve the model itself, by proposing points away from previous observations — and exploitation — picking points close to the current minimizer in order to refine the best solution. Thus, referring to Eq. (4.3), a typical BO method implements $\pi_{1}$ with the optimization of an AF and $\pi_{2}$ with a Bayesian update of the stochastic surrogate model.
First, we discuss standard choices to design and fit surrogate models in BO, including Gaussian processes and tree Parzen estimators. Then, we focus on AFs, especially on the expected improvement [mockus1978application]. One may say that a most classic instantiation of BO consists in modelling the surrogate with a Gaussian process and using as AF the expected improvement. This approach is well-suited for the global optimization of smooth multivariate real functions, and it provides an accessible setting to approach BO for hyperparameter optimization. Nevertheless, it may not be the most suitable for some HPO problems; thus we discuss limitations along the way.
4.2.1 Bayesian linear model
We start by introducing a simple linear model $h_{w}$ in a given feature space $\mathcal{V}$, to which we add independent Gaussian noise with variance $\sigma^{2}$:
$h_{w}(\lambda)=\varphi(\lambda)^{\intercal}w+\varepsilon;\qquad\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}),$ (4.4)
where $\varphi:\Lambda\to\mathcal{V}$ is a feature map. While using linear maps directly in the input space would result in an overly simplistic model – we are trying to model an objective, the validation error, which can be (highly) non-convex and multimodal – working in a different feature space allows us to capture much richer behaviors but still retain most of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in Eq. (4.4) models potential errors in the observations (see also 4.2.4).
Suppose now we are at the $k$-th iteration of the algorithm, after having already collected $k$ observations $\mathcal{C}=\{(\lambda_{i},f(\lambda_{i}))\}_{i\in[k]}$. Let $Y_{k}=\left(f(\lambda_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ be the vector of observed values, and denote by $X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)$ the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
$y_{*}\sim\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k}),$
where $x_{*}=\varphi(\lambda_{*})$ is a query point in the feature space. This is going to be useful later in Section 4.2.3, where an AF employs $\hat{f}_{k}$ to select the next configuration $\lambda_{k+1}$.
Modelling uncertainty away from observations.
We could follow an empirical risk minimization approach, obtaining the weights $w_{k}\in\mathbb{R}^{\mathrm{dim}(\mathcal{V})}$ of the model Eq. (4.4) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k+1}=h_{w_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w_{k})$, capable of capturing only our belief of the amount of noise present in the observations. Rather, we would like the stochasticity of $\hat{f}$ to (primarily) represent our uncertainty about the behavior of $f$ in unexplored regions. To better grasp the difference between these two aspect, consider the case where $f$ is a (non-linear) smooth deterministic function.  [^1] The following argument is mutated from jones1998efficient. If we were to follow an ERM approach, we could either omit the noise term $\varepsilon$ in Eq. (4.4), or leave it for capturing modelling errors. However, the first scenario, would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously observed. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k+1}$ would trivially be inaccurate at $\lambda_{i}$ for $i\leq k$ (as $f$ is deterministic). Thus, both ERM modelling options would leave us with a surrogate model that is almost surely inaccurate.
A way to escape this apparent stalemate is to replace the independent Gaussian noise with one that depends on the evaluation point, i.e. $\varepsilon=\varepsilon(\lambda)$. Then, when fitting the model to the $k$ observations, one may have $\varepsilon_{k}(\lambda_{i})=0$ for $i\leq k$ while letting $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, by continuity ($\varepsilon_{k}$ is smooth and continuous as difference of $C^{1}$ functions), $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behaviour.
Thus, we can reason about all the linear models (in the feature space) that agree with the observations, weighted by their “realization probability”.
More formally, we marginalize over the weights $w$ and set
$\hat{f}_{k+1}(\lambda_{*})=\int p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)p_{w}(w|X_{k},Y_{k})\,\mathrm{d}w.$ (4.5)
Given our modeling assumption (4.4) and the hypothesis that the noise is independent, the first distribution simply reads as
$p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)=p_{y_{*}}(\cdot|x_{*},w)=\mathcal{N}(h_{w}(\lambda_{*}),\sigma_{\varepsilon}^{2})=\mathcal{N}(x_{*}^{\intercal}w,\sigma_{\varepsilon}^{2}).$ (4.6)
By the Bayes’ rule, $p_{w}(\cdot|X_{k},Y_{k})$ is instead given by
$p_{w}(\cdot|X_{k},Y_{k})=\frac{p_{y}(\cdot|X_{k},w)p_{w}(\cdot|X_{k})}{p_{y}(\cdot|X_{k})}.$ (4.7)
Regarding the numerator of Eq. (4.7), we have that, for the same argument of Eq. (4.6),
$p_{y}(\cdot|X_{k},w)=\mathcal{N}(X_{k}^{\intercal}w,\sigma_{\varepsilon}^{2}I).$
This is the likelihood of the observations given the model. We must instead make a decision regarding the second term of the numerator, that is the distribution of the model parameters. We make the very natural assumption of independence from the design matrix and set
$p_{w}(\cdot|X_{k})=p_{w}=\mathcal{N}(0,\Sigma),$ (4.8)
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with Eq. (4.8), we can now also compute the denominator of (4.7) by marginalizing $p_{y}(\cdot|X_{k})$ over $w$. We obtaining
$p_{y}(\cdot|X_{k})=\int p_{y}(\cdot|X_{k},w)p_{w}(w)\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I)$ (4.9)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the kernel corresponding to the feature map $\phi$ as follows:
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\varphi(\lambda)^{\intercal}\Sigma\,\varphi(\lambda^{\prime})=\langle\varphi(\lambda),\varphi(\lambda^{\prime})\rangle_{\Sigma}$ (4.10)
and rewrite Eq. (4.9) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathbb{R}^{k\times k}$ as
$p_{y}(\cdot|X_{k})=\mathcal{N}(0,K+\sigma_{\varepsilon}^{2}I).$
We finally have all the terms to compute (4.5); which is, again, Gaussian:
$\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k})=\\ \mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)k_{*})$ (4.11)
where $k_{*}=\{\kappa_{\Sigma}(\lambda_{*},\lambda_{i})\}_{i=1}^{k}\in\mathbb{R}^{k}$ is the vector of kernel evaluations between the query and the observations points. Equation (4.11), despite its apparent complexity, describes a rich posterior distribution over the values of $f$ that is analytically computable and whose design depends essentially only on the definition of the kernel map (4.10).
Examples of kernel functions include the $\mathcal{C}^{\infty}$ squared exponential
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\sigma_{0}e^{-r_{\Sigma}^{2}(\lambda,\lambda^{\prime})/2};\qquad r_{\Sigma}^{2}(\lambda,\lambda^{\prime})=\lambda^{\intercal}\Sigma\lambda^{\prime},$
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$ and the Matérn family, which contains kernels (defined as functions of $r_{\Sigma}$, above) with various degree of smoothness [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow to sidestep the definition of (possibly obscurer) feature maps and covariance matrices. The series of “free parameters”222 That is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel. that appears in (4.11) – usually called hyperparameters in the Bayesian literature (not to be confused with $\lambda$) – do not have to be fixed constants, but may be estimated from the observations [williams2006gaussian, shahriari:2016:taking_human_out] (automatic relevance determination). Thus, the design choice left to the experimenter mainly boils down the choice of the kernel function.
4.2.2 Gaussian processes
Instead of worrying about feature maps and weights, we can directly work with a kernel function, a notion which gives rise to Gaussian processes (GP) [williams2006gaussian, murphy2012machine]. In fact, the Bayesian linear model converges to a GP in the limit of infinitely many features, if the prior distribution on the weights is scaled appropriately [neal1996bayesian].
A Gaussian process is a Gaussian distribution over functions $g(\lambda)$, fully specified by a mean function $\mu(\lambda)$ and a covariance function $\kappa(\lambda,\lambda^{\prime})$, or kernel function. It has the property that for any finite set $\lambda_{1},\dots,\lambda_{n}\in\Lambda$, the joint distribution of the vector $[g(\lambda_{1}),\dots,g(\lambda_{n})]$ is multivariate Gaussian with mean $[\mu(\lambda_{i})]_{i=1,\dots,n}$ and covariance matrix $[\kappa(\lambda_{i},\lambda_{j})]_{i,j=1,\dots,n}$. Our linear model derivation above gives rise to a zero-mean Gaussian process (since the prior over weights (4.8) has mean zero) with covariance function $\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta$, where $\delta(\lambda,\lambda^{\prime})=1$ if $\lambda=\lambda^{\prime}$ and 00 otherwise. In short, we can write
$\hat{f}\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta).$
The surrogate model update equation $\pi_{2}$ (see (4.3)) may then be written as
$\hat{f}_{k+1}=\pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\xi_{2}\right)\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta\,|\,X_{k},Y_{k}),$ (4.12)
which expresses the update as the conditioning of the Gaussian process on the observations $X_{k}$, $Y_{k}$. While this posterior is in principle an infinite-dimensional object, we only ever need its marginal distributions over single points $\lambda_{*}$ (or over small sets, such as pairs) in order to compute AFs and make sampling decisions. These marginals are Gaussian with mean and (co)variance given by simple linear algebra expressions (4.11).
Limitations of Gaussian processes.
Gaussian processes also come with a number of drawbacks that may limit their applicability to real-world scenarios. Most importantly, computing the posterior for $k$ datapoints costs $\mathcal{O}(k^{3})$, as can be seen in (4.11). If we are able to sample $f$ a few thousand times or more, GP-based BO starts to become intractable. It is also difficult to design kernel functions for high-dimensional $\Lambda$, or if the search space contains many integer, categorical, or conditional hyperparameters. In such situations, BO with other surrogate models can perform better, as shown in a comparative study by eggensperger_towards_2013.
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions, and take on the scalability issues in snoek_scalable_2015 by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms and bergstra_making_2013 focus both on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators (see also Section 4.2.5). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next section) to improve performance in the first stage of HPO, when the surrogate model may still be rather unreliable.
4.2.3 Acquisition functions
Suppose we fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $\mathcal{D}=\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function $\alpha(\lambda|\hat{f}_{k})$. In other words, our suggestion mapping $\pi_{1}$ becomes
$\lambda_{k+1}=\pi_{1}(\hat{f}_{k},\nu_{2})=\arg\max_{\lambda\in\Lambda}\,\alpha(\lambda|{\hat{f}_{k}}).$ (4.13)
In this section, we restrict our attention to Gaussian process surrogate models, even though most AFs discussed here can be applied to other models just as well. As we will see, most common AFs depend on the posterior distribution of $\hat{f}_{k}$ only through its marginal mean and variance
$\mu_{k}(\sigma)=\mathbb{E}\left[\hat{f}_{k}(\lambda)\right],\quad\sigma_{k}(\lambda)^{2}=\mathbb{V}\left[\hat{f}_{k}(\lambda)\right],$
given by (4.11). Here, $\mathbb{E}$ and $\mathbb{V}$ denotes expectation and variance over the posterior distribution of $\hat{f}_{k}(\lambda)$ respectively.
A useful AF has to guide the choice of the next candidate so as to gain as much new information about high-performing hyperparameters as possible, given a limited sample $\mathcal{D}$ obtained so far. In particular, it needs to balance the trade-off between exploration (i.e., looking away from points in $\mathcal{D}$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $\mathcal{D}$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
$\alpha_{\text{exploit}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda).$
Maximizing $\alpha_{\text{exploit}}$ constitutes an extreme form of exploitation, which typically gets stuck in suboptimal local optima. On the other extreme, we could sample where we know least about $f$:
$\alpha_{\text{explore}}(\lambda|\hat{f}_{k})=\sigma_{k}(\lambda).$
Maximizing $\alpha_{\text{explore}}$ is entirely focused on exploration and may waste many evaluations in regions of poorly performing hyperparameters. However, a linear combination of the two,
$\alpha_{\text{LCB}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda)+\eta\sigma_{k}(\lambda),\quad\eta>0,$
is a competitive and frequently used AF, called lower-confidence bound (LCB) [Srinivas2009].
Expected improvement.
A classic AF, introduced in the seminal work by mockus1978application and frequently used in HPO libraries [snoek_practical_2012] and services [AMT2020], is the expected improvement (EI), defined as
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})\doteq\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
Here, $\beta\in\mathbb{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far, $\beta=f(\hat{\lambda}_{k})$. In contrast to LCB, EI does not come with a free parameter to be set. In addition, it can be shown to have a simple closed-form expression in terms of $\mu_{k}(\lambda)$ and $\sigma_{k}(\lambda)$:
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})=\sigma_{k}(\lambda)\nu(\lambda)\Phi\left(\nu(\lambda)\right)+\sigma_{k}(\lambda)\phi\left(\nu(\lambda)\right)$ (4.14)
where
$\nu(\lambda)\doteq\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}$
and where $\Phi$ and $\phi$ denote the cumulative distribution function and the probability density function of the standard Gaussian distribution, respectively.
Equations (4.13) where $\alpha$ is the expected improvement and (4.12) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. As an illustrative example, two steps of the Bayesian optimization algorithm using EI as an AF are shown in Figure 4.1.
Figure 4.1: Left: Suppose we have measured the response function (plotted in red) on a small set of points. Fitting a Gaussian process on these points yields a mean function (plotted in black) with and a variance function ($\pm$ one standard deviation is filled in gray). If $\lambda$ falls between 7 and 11, the uncertainty is very high, which yields the right bump in the expected improvement (plotted in green), allowing us to explore a region where potentially a smallest response might be found. The left bump of the expected improvement falls in an exploitation region, which in this example looses against exploration. Hence, the next candidate is $\lambda=8.5$. Right: once the response is measured at the new candidate, the Gaussian process is trained again. Now uncertainty shrinks significantly and the new expected improvement has just one bump around 2.7, which is already very close to the minimum of the true response. This example takes inspiration from Figure 11 in [Jones1998:Efficient].
Other acquisition functions.
An earlier and simpler variant of EI is the probability of improvement [Kushner1964], given by
$\alpha_{\text{PI}}(\lambda|\beta,\hat{f}_{k})\doteq P\left[\hat{f}_{k}(\lambda)<\beta\right]=\Phi(\nu(\lambda)).$ (4.15)
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points333 Which costs $\mathcal{O}(b^{3})$ time and $\mathcal{O}(b^{2})$ memory, if $b$ is the size of the set. [unify notation please] or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. While these have a sound theoretical foundation and have been shown to be more sample efficient, they typically come with higher computational cost and are more complex to implement than standard options such as EI or LCB. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out].
Optimizing acquisition function in practice
Techniques for maximizing AFs include grid search, where the AF is evaluated at a predefined set of points, and random search, which evaluates the AF at randomly selected points. More sophisticated methods involve the use of gradient-based optimization, which leverages the gradient of the AF to find local maxima more efficiently.
4.2.4 Aleatoric and epistemic uncertainty
In (4.4) we have introduced a noise term $\epsilon$ in order to account for the stochastic nature of the learning algorithm itself. This derives from the internal randomness of the training algorithm, (for example due to initial weights when training a deep network). In general, the amount of this aleatoric uncertainty may depend on the current hyperparameters and should not be confused with the epistemic uncertainty which arises because we may have not (yet) evaluated a large enough set of hyperparameter vectors. The latter uncertainty is crucial to allow explorations in regions not yet assessed, while the former should not mislead us to acquire points where uncertainty cannot be reduced.
A heteroscedastic approach can help disentangle these types of uncertainties, as proposed by [griffiths:2021]. In their approach, distinct Gaussian processes are trained aiming to predict the unknown response and the noise separately.
4.2.5 Tree Parzen estimator
The Tree Parzen Estimator (TPE) is a probabilistic model to efficiently explore complex search spaces. It models the objective function as a pair of conditional probability distributions: one for configurations that yield better results than a threshold and another for those that perform worse. By iteratively selecting hyperparameter values that maximize the expected improvement, TPE focuses the search on promising regions. This approach is particularly well-suited for high-dimensional, non-convex optimization problems and balances exploration and exploitation to accelerate convergence toward optimal configurations. While Gaussian process based BO methods model the distribution $p(y|\lambda)$ of observations $y$ given configurations $\lambda$, the Tree Parzen Estimator (TPE) [bergstra2011algorithms, bergstra_making_2013] fits two conditional densities over $\Lambda$:
$l_{\beta}(\lambda)=p(\lambda|y<\beta),\quad g_{\beta}(\lambda)=p(\lambda|y\geq\beta).$ (4.16)
Here, the baseline $\beta$ is set to a lower $\gamma$-quantile of the observed targets $\{y_{i}\}$ (e.g., $\gamma=0.15$), giving rise to a $(\gamma,1-\gamma)$ split of the data, on which $l_{\beta}(\lambda)$, $g_{\beta}(\lambda)$ are fit respectively. While TPE uses a tree of Parzen estimators for conditional search spaces, later work made use of kernel density estimators [falkner2018bohb]. The density ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$ is a strictly increasing function of the expected improvement444 However, note that we cannot set $\beta$ to the minimum value observed so far, as in the GP-BO case, since this implies $\gamma=0$ and no data to fit $l_{\beta}(\lambda)$ with. AF with baseline $\beta$, and TPE maximizes this AF approximately by sampling a set of configurations from $l_{\beta}(\lambda)$ and suggesting the one with the largest ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$. TPE is conceptually simple, easy to run with parallel evaluations, and can naturally deal with conditional search spaces. Subsequent work has shown that density ratios can be estimated by way of probabilistic binary classification, which is simpler and better supported by ML methodology than density estimation [Tiao:21].
4.3 Other approaches to model-based HPO
Not all model-based methods follow a Bayesian optimization framework. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of $f$, while hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer to perform HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, linked to gradient-based HPO may also be interpreted as an attempt to learn a surrogate model of the entire learning algorithm (applied to a specific dataset), whereby the output space of $\hat{f}$ coincides with the hypothesis space of $A$ itself, that is $\hat{\mathcal{Y}}=\mathcal{H}$ in (4.1). We cover gradient-based HPO in Chapter 7.
Furthermore, not all methods follow the iterative schema of Eq. (4.3). Another possibility is to first learn $\hat{f}$ and then use it to suggest one or several hyperparameter configurations in a second stage, following a parallel scheme. Examples of this latter approach include [feurer2015initializing, fusi2018probabilistic]. These methods, however, assume the presence of multiple tasks and are closer in nature to meta or transfer learning settings.
More sophisticated strategies to exploit parallel compute resources, either with random search or BO, are described in the next chapter.
4.4 Example: model-based vs random search
We illustrate the difference between model-based and random search in a concrete example of tuning regularization parameters of an ML model to prevent overfitting. We tuned the L1 and L2 regularization terms, respectively alpha and lambda, of the XGBoost algorithm on the UCI direct marketing binary classification dataset from Kaggle. The goal was to minimize the AUC. Figure 4.2 presents a comparison between random search and BO (model-based), which models the objective function through a GP and uses the expected improvement as the acquisition function.
Each experiment was conducted using 50 different random seeds, and the results were reported as the average and standard deviation. The left and middle plots reveal that BO consistently suggests better-performing hyperparameters than random search. The right plot further demonstrates that BO consistently outperforms random search across all evaluated hyperparameter configurations.
While BO is particularly effective in a sequential setting, random search remains a valuable approach, especially in distributed environments where many configurations can be evaluated in parallel, significantly reducing wall-clock time.
Figure 4.2: Top: Hyperparameters suggested by random search (left) and BO (right). The $x$-axis and $y$-axis are the hyperparameter values for alpha and lambda and the color of the points represent the quality of the hyperparameters (yellow: better AUC; violet: worse AUC). Bottom: Best model score obtained so far ($y$-axis, where lower is better) as more hyperparameter evaluations are performed ($x$-axis).
4.5 Discussion
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an AF depending on the surrogate model, as we saw in Section 4.2.3. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of MB techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
While the foremost representatives of these type of methods – tied to BO – are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in the next chapter. Before doing so, we spend some words on what is possibly the biggest practical limitation of (naive) model-base methods: parallelizability.
4.5.1 Decision making under parallelism
In its original formulation BO is sequential, which is unsuitable for many real-world scenarios where parallel computation resources are readily available in order to increase wall-clock time efficiency. A number of extensions of BO to parallel evaluation settings have been proposed. Although more time efficient, these techniques tend to be less sample efficient than sequential BO. The extensions we discuss aim thus to find a diverse set of configurations to preserve sample efficiency as much as possible.
For simplicity, assume a fixed set of $W$ workers available for parallel evaluations. Recall from Section 2.4 that systems may either schedule evaluation jobs synchronously or asynchronously. In the synchronous case, we need to suggest batches of $W$ configurations to be evaluated next. A simple way to do this would be to pick the $W$ top-ranked configurations when optimizing our AF. However, this strategy risks suggesting very similar configurations, and their parallel evaluation may reveal little more information than any single one of them. A heuristic approach to suggest a more diverse batch is to run Thompson sampling $W$ times with independently drawn sample paths [Kandasamy:18], but this can be quite expensive. Another idea is to select batch members greedily one at a time. For each configuration placed in the batch, the AF is modified by a local penalty function, ensuring that subsequently selected points are at least some distance away [Gonzales:16].
Independent of how configurations are suggested, synchronous job scheduling tends to be substantially less efficient than asynchronous. Namely, the time to process a batch is determined by the slowest one, while up to $W-1$ workers are idle waiting. Such differences in training times are particularly pronounced when searching over parameters determining the model size. In an asynchronous system, whenever a worker becomes available, it can be assigned a new configuration. While BO needs to make single suggestions only (as opposed to batches), this has to be done in the presence of up to $W-1$ pending evaluations, i.e. jobs which have been started for previously suggested configurations, but did not return a metric value yet. Ignoring pending evaluations runs the same risk of selecting highly redundant configurations in subsequent rounds. Fortunately, dealing with pending evaluations is quite related to ensuring diversity in batch selection, and local penalty methods apply just as well [Gonzales:16, Alvi:19]. Simpler approaches start from the observation that in sequential BO, diversity is enforced by conditioning on observed points. The resulting very low posterior variance shrinks the AF to small values near such points. In the kriging believer heuristic [Ginsbourger:10, Desautels:14], we impute the response for each pending evaluation by its current posterior mean and condition on this pseudo-datapoint. A somewhat more principled approach is fantasizing [snoek_practical_2012, Wilson2018], where pending target values are integrated out by Monte Carlo: they are sampled several times from the joint posterior over pending configurations, and the AF is averaged over these samples.
Chapter 5 Multi-fidelity Methods
One of the main challenges of HPO is that training and validating a machine learning algorithm each time the objective function is evaluated can be expensive. This is particularly true for modern machine learning algorithms applied to large datasets, such as deep neural networks, which can often take hours or days to train. In this chapter, we consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are much faster to evaluate than computing the full objective function and often provide a strong enough signal to guide the optimization process.
We begin with an overview of different types of fidelities in Section 5.1. Section 5.2 explores methods for multi-fidelity optimization based on randomly sampling configurations. Finally, in Section 5.3, we examine multi-fidelity approaches that leverage probabilistic models to guide the search.
5.1 Multi-fidelity optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is fast turnaround time. To achieve this, practitioners typically work with a subset of the training data during their initial explorations. They also do not wait for training to fully converge but instead check validation performance after a few iterations or epochs. While these early metrics are not entirely reliable, they are useful for identifying promising regions of the search space.
In this section, we discuss how these manual practices can be transformed into algorithmic tools that enable automated early stopping of underperforming runs.
Multi-fidelity HPO [li2017hyperband, klein2017fast, swersky2014freeze] extends the HPO formulation from Section 2.2 by taking the budget that we spend to train a hyperparameter configuration into explicit account. We assume access to inexpensive approximations of the true objective function $f(\lambda)$, denoted as $f_{b}(\lambda)$, where the fidelity level is determined by the budget $b\in[b_{min},b_{max}]$.
If we set the budget to its maximum value, $b=b_{max}$, we recover the original objective function, $f_{b_{max}}(\lambda)=f(\lambda)$. From this perspective, standard HPO can be seen as a special case of multi-fidelity HPO. Additionally, we assume that the approximation of $f(\lambda)$ improves as $b$ increases. However, the cost $c(\lambda,b)$ associated with evaluating $f_{b}(\lambda)$—such as training time—also increases with $b$.
While we can freely choose $b$ during the optimization process, our primary interest lies in the performance at the full budget $b_{max}$. More formally, our goal is (still) to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$, which coincides with the formalization of Chapter 2.
In the next section, we explore commonly used fidelity measures that can be leveraged to accelerate the optimization process.
5.1.1 Iterations
For iterative machine learning algorithms, such as neural networks, we can interpret the number of iterations as the budget and use the performance at each iteration as an approximation of the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. In the case of neural networks, we typically consider epochs as iterations rather than individual mini-batch updates to reduce the overhead of validation error computations.
Figure 5.1 (left) illustrates the validation loss after each epoch of training for different hyperparameter configurations of a multi-layer perceptron. Even after just a few epochs, we can visually distinguish between poorly performing and well-performing configurations. However, to reliably identify the best configuration, it is often necessary to evaluate the top-performing candidates for the full number of epochs.
5.1.2 Training dataset subsets
Figure 5.1: Left: Validation loss of different neural networks evaluated after each epoch of training. Right: Validation error of a support vector machine trained on different subsets of the training data. Each color represents a different hyperparameter configuration.
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 (right) shows the validation error of a support vector machine trained on increasing subsets of the MNIST dataset, with each color representing a different hyperparameter configuration. As expected, the validation error decreases as the training dataset size increases. More interestingly, however, the ranking of hyperparameter configurations remains relatively consistent across dataset subsets [bornschein-icml20].
In the context of deep neural networks, this observation is related to recent research on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which empirically demonstrate that the training loss of autoregressive transformer models [vaswani_attention_2017] often follows a power-law function with respect to the size of the training dataset.
In multi-fidelity optimization, we can exploit this phenomenon by interpreting the budget as the training dataset size [klein2017fast, swersky_multi-task_2013]. To evaluate $f_{b}(\lambda)$, we vary only the training dataset size while keeping the validation set fixed to avoid introducing additional noise. Most methods resample the training dataset uniformly at random for each evaluation of $f_{b}(\lambda)$, though more sophisticated strategies, such as importance sampling, can also be used to select dataset subsets [ariafar-aistats21]. Importantly, global statistics such as the class distribution must be preserved to avoid biasing performance estimation.
5.1.3 Cross-validation
As discussed in Section 2.2.1, cross-validation (CV) is a widely used technique for estimating the generalization error of a machine learning method and reducing the risk of overfitting. Rather than relying on a single training/validation split, we divide the dataset into $K$ subsets or folds, using each fold for validation in turn and averaging our target criterion across them. The main drawback of CV is its computational cost, which is $K$ times higher than using a single fold. However, we can approximate it using a multi-fidelity approach by treating the number of folds, $b=1,\dots,K$, as the budget.
Unlike the previous two fidelities, where performance typically improves with increasing budget, CV does not necessarily exhibit this behavior. Instead, the key assumption is that the variance of the generalization performance estimate decreases as the number of folds increases.
5.2 Methods based on random search
In this section, we review algorithms that leverage multiple fidelities and propose new configurations through random sampling. Random search based multi-fidelity methods were first introduced in the multi-armed bandit (MAB) literature. We discuss how a simple MAB approach can dynamically halt the evaluation of poorly performing configurations. Additionally, we discuss how these methods can be adapted for efficient execution in large-scale distributed settings. The methods covered here assume that the evaluation costs, $c(\lambda,b)$, for any fixed $\lambda$ are linear (or affine linear) in $b$.
5.2.1 Successive halving
Successive halving (SH) [karnin-icml13, jamieson2016non] is an MAB strategy that aims to identify the best hyperparameter configuration from a finite set of configurations. Given the maximum and minimum budgets, $b_{max}$ and $b_{min}$, for evaluating a single hyperparameter configuration, we first define a set of rung levels:
=$\mathcal{R}=\{b_{min}\eta^{k}\mid k\in[K]\}$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{K}$, where $K\in\mathbb{N}$. Here, $\eta\geq 2$ defines the halving constant and determines how aggressively SH prunes configurations.
Algorithm 2 Successive-Halving-HPO
1:Initial budget $b_{min}$, maximum budget $b_{max}$
2:Set of $n$ candidate configurations $\mathcal{C}_{0}=\{\lambda_{1},\lambda_{2},\dots\lambda_{n}\}$
3:Halving constant $\eta\geq 2$
4:for $i\in[|\mathcal{R}|]$ do
5:     $\{f_{r_{i}}(\lambda):\lambda\in\mathcal{C}_{i-1}\}$triangleright Evaluate configurations on the rung
6:     $k\leftarrow\lfloor|\mathcal{C}_{i-1}|/\eta)\rfloor$triangleright Number of configurations to keep
7:     $\mathcal{C}_{i}\leftarrow\underset{\lambda\in\mathcal{C}_{i-1}}{\arg\mathrm{top}\text{-}k}f_{r_{i}}(\lambda)$triangleright Keep $k$ best performing configurations
8:end for
Algorithm 2 shows pseudo code for SH. Given a set $\mathcal{C}_{0}$ of $n$ hyperparameter configurations, usually drawn uniformly at random, we evaluate them at the first rung level, i.e., each configuration is evaluated with a budget of $b_{min}$ (Line 2). For convenience, the number of initial configurations is often set to $n=\eta^{K}$, but this can also be arbitrary. Afterwards, SH promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance, to the next higher rung level $b_{min}\eta$ (Line 4). These steps are iterated until we reach the highest rung level, $b_{max}$. See Figure 5.2 for a visualization.
While conceptually simple, SH has shown strong performance in the literature and even outperformed more sophisticated Bayesian optimization techniques that only operate on the full budget [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration at the first rung level should be representative of its performance at the highest rung level. Otherwise, configurations that perform best at $b_{max}$ may be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $\frac{1}{\eta}$ of the configurations to the next rung.
5.2.2 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while if $b_{min}$ is too large, our advantage over standard random search may diminish. Hyperband [li2017hyperband] addresses this issue by running different instances, $l=0,\dots,K$, of SH in sequence. The $l$-th instance (or bracket) has rung levels
$\mathcal{R}_{l}=\{b_{min}\eta^{k}\mid k\in\{l,\dots,K\}\},$
so $l=0$ corresponds to SH with $K+1$ rung levels, while $l=K$ corresponds to random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at a modest extra expense.
5.2.3 Synchronous successive halving
How is parallel computation exploited by SH? Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,M)$ jobs in parallel, where $M$ is the number of free slots remaining in the currently processed rung. Here, $W-M$ workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has only a single slot, so $W-1$ workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.
Hyperband inherits the idle worker issue noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which (at the expense of some implementation complexity) eliminates idle time for asynchronous scheduling.
5.2.4 Asynchronous successive halving (ASHA)
As discussed above, neither SH nor Hyperband can fully exploit parallel computation because they require synchronization among workers evaluating hyperparameter configurations for a specific run.
This also means that if a configuration is clearly superior to all others on a rung $r_{i}\in\mathcal{R}$, it may take a long time before it is promoted to the next rung level $r_{i+1}$. In fact, all configurations on the current rung $r_{i}$ must be evaluated before any promotion decisions can be made.
For example, consider Figure 5.3. In a distributed setting, we refer to the evaluation of a hyperparameter configuration as a trial, and in this scenario, we have two workers evaluating four different trials on Rung 0. If Trial-2 takes significantly longer than the others, worker 1 remains idle until all trials on Rung 0 are completed, delaying progress to Rung 1.
Figure 5.3: Scheduling behavior of Synchronous SH with two workers. A darker shade of blue indicates better-performing trials. Some trials may take significantly longer than others, resulting in idle workers until all trials in the current rung are completed.
These shortcomings are addressed by Asynchronous Successive Halving (ASHA) [Li:19]. In this algorithm, rungs do not have a predefined size but grow dynamically over time. Additionally, promotion decisions are made more aggressively.
For example, assume that $\eta=2$; a configuration is immediately promoted to the next rung if it ranks in the top $50\%$ of the current rung. In other words, configurations are promoted as early as possible: once a rung has two entries, one of them advances to the next level. Figure 5.4 illustrates this process: as soon as Trial-1 is completed, evaluation of Trial-0 continues on Rung 1—assuming Trial-0 performs better than Trial-1 — rather than evaluating Trial-3 on Rung 0.
After evaluating Trial-3, no configurations can be promoted because, at this point, only two configurations have been evaluated on Rung 0, and one on Rung 1. As a result, the evaluation proceeds with a new trial at the lowest level.
Figure 5.4: Scheduling behavior of ASHA. As in Figure 5.3, a darker shade of blue indicates better-performing trials. ASHA eliminates idle time by immediately promoting a trial to the next rung as soon as it meets the promotion criteria, rather than waiting for all trials in the current rung to be completed.
In ASHA, all available computational resources are utilized immediately, ensuring that workers are never idle. Additionally, superior configurations are quickly promoted toward $b_{max}$ without the need to wait for synchronization points. In practice, ASHA often significantly outperforms synchronous SH or Hyperband [Klein:20].
5.3 Model-based methods
So far, the multi-fidelity methods we have examined sample hyperparameter configurations purely at random. As we saw in Chapter 4, more sophisticated methods that use a probabilistic model to sample configurations typically require fewer function evaluations to approach the global optimum compared to random search-based approaches. In this section, we will discuss multi-fidelity strategies that leverage one or multiple probabilistic models to guide the search.
5.3.1 Fabolas
Fabolas [klein2017fast] models the performance $f_{b}(\lambda)$ of a hyperparameter configuration $\lambda$ across different subsets $b\in[b_{min},b_{max}]$ of the training data.
Compared to classical Bayesian optimization (see Chapter 4), we cannot simply optimize standard acquisition functions as described in Section 4.2.3. For example, in the case of expected improvement, the highest improvements are expected at $b=b_{max}$. As a result, this approach would lead to over-exploitation, primarily evaluating configurations at the highest budget.
Instead, Fabolas is based on the entropy search method by Hennig2012, which models the probability
$p_{\min}(\lambda^{*}\mid\mathcal{C})=\mathbb{P}(\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda))$
of $\lambda$ being the global optimum, where $\mathcal{C}$ are the collected observations (see Section 4.2). The key idea of entropy search is to select, in each iteration, the configuration that provides the most information about $p_{\min}$, which we quantify by computing the KL divergence $KL\big{[}p_{\min}(\lambda\mid\mathcal{C})||\mathcal{U}(\lambda)\big{]}$ between $p_{\min}$ and a uniform distribution $\mathcal{U}(\lambda)$, which serves as a prior assuming that each point $\lambda\in\Lambda$ is equally likely to be the optimum. In other words, entropy search selects, in each iteration, the candidate point $\lambda$ that shifts $p_{\min}$ from a uniform distribution to a highly peaked distribution around the global optimum.
For Fabolas, we model $p_{\min}$ only at the highest budget $b=b_{max}$, since we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$. To balance information gain with the cost of evaluating $\lambda$, Fabolas maintains a second Gaussian process $p(c\mid\lambda,b)$ to model the wall-clock time $c(\lambda,b)$ required to evaluate $f_{b}(\lambda)$. In each iteration, Fabolas optimizes the following acquisition function:
$\alpha(\lambda,b)=\frac{KL\big{[}p_{\min}(\lambda\mid\mathcal{C})||\mathcal{U}(\lambda)\big{]}}{\mu_{c}(\lambda,b)}$
to select both the configuration $\lambda$ and budget $b$ that provide the most information about $p_{\min}$ while remaining cost-effective to evaluate, where $\mu_{c}(\lambda,b)=\mathbb{E}_{p(c\mid\lambda,b)}[c(\lambda,b)]$ is the expected cost of evaluating $\lambda$ using budget $b$.
5.3.2 BO-HB: Bayesian Optimization Hyperband
To combine the sample efficiency of Bayesian optimization with the strong anytime performance of Hyperband, falkner2018bohb fit an independent TPE model (see Section 4.2.5) on each rung level $b_{i}\in\mathcal{R}$ to sample new configurations.
More specifically, once the proposed method BO-HB observes $N_{\min}$ data points at the rung level $b_{i}$, it trains two kernel density estimators—$l(\lambda)$ and $g(\lambda)$—to model the probability of a configuration $\lambda$ achieving a validation error lower or higher than a certain threshold, as defined in Eq. 4.16. However, the density estimators are only trained on all data points at the given rung level.
New configurations are then sampled according to the density ratio $l(\lambda)/g(\lambda)$ instead of being chosen uniformly at random. This process continues until $N_{\min}$ observations have been collected at the highest rung level $b_{max}$, after which new configurations are sampled exclusively from the model trained at this level.
While the original implementation of TPE by bergstra2011algorithms used a univariate kernel density estimator, falkner2018bohb proposed using a multivariate kernel density estimator to better capture interactions between hyperparameters. To fit such a model, at least $N_{\min}=m+1$ data points are required, where $m$ is the dimensionality of the search space.
Figure 5.5 presents a comparison by falkner2018bohb of BO-HB with other approaches for optimizing the hyperparameters of a neural network. Hyperband initially outperforms other methods since it returns configurations much earlier than non-multi-fidelity methods such as random search or TPE. However, compared to model-based approaches like TPE, Hyperband eventually converges toward random search, as it samples configurations uniformly at random and, in later stages, requires evaluating them on larger budgets to improve performance. BO-HB, on the other hand, retains the early speed-up advantage of Hyperband while achieving a performance level similar to TPE in the later stages of the optimization process.
Figure 5.5: Comparison of random search (RS), TPE, Hyperband (HB) and BO-HB overtime from falkner2018bohb. BO-HB keeps the strong anytime performance of Hyperband but because of the model samples better configurations over time.
Chapter 6 Population-based Methods
In this chapter we discuss a class of randomized methods for global optimization that evolve a population of candidate solution, drawing inspiration from the process of natural evolution [back1997handbook, simon2013evolutionary]. Evolutionary algorithms date back to the beginning of computer science [neumann1966theory] and have been used to tackle non-smooth and non-convex problems across many domains. Methods in this class typically rely on function evaluations only. Their application to hyperparameter optimization includes methods like CMA-ES [hansen1996adapting], population-based training (PBT) [jaderberg2017population] and particle swarm optimization [lorenzo2017particle], which we review below.
6.1 Evolutionary algorithms
Traditionally, evolutionary algorithm (EA) distinguish between evolutionary strategies (ES) and genetic algorithms (GA). GA refers to a class of methods that deal with binary problems and discrete objective functions while ES is concerned with real-valued, continuous objectives.
The terminology of EA techniques is borrowed from biology: the set of points is called population and configurations are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically at the end of each generation the fitness of all individuals is observed, making EA inherently parallel algorithms – traditional implementations usually requiring synchronization. After observing the fitness of the population, a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include:
  • Selection, global transformations that prime a new generation, typically by discarding individuals with low fitness and selecting a subset of fitter individuals to become “parents” for the next generation;
  • Recombination or crossover, global transformations that consist in generating new individuals combining slices (e.g. subsets of hyperparameter coordinates) of two or more parents;
  • Mutation, local modifications designed to escape local optima that introduce random changes in the individual, for instance adding independent Gaussian noise.
Other strategies like elitism (keep the best individuals unchanged) and niching (maintain a diverse population) are often used in conjunction to the explore/exploit strategies outlined above.
We described this class of algorithm by the procedure outlined in Population-Based-HPO in Algorithm 3. W can think about individuals and generations as draws from $N$ distributions $p_{k}^{j}$, where $N$ is the population size. These distributions implement the strategies described above, and evolve at each generation using global information from the current generation. For instance, elitism can be represented by setting $p_{k}^{j}=\delta_{\lambda_{k}^{j}}$ for the best performing configurations, where $\delta$ is a Dirac delta distribution. In contrast to random or quasi-random search methods covered in Chapter 3, population-based methods keep in general $N$ distinct distributions (rather than a global one) that are updated iteratively (rather than fixed).
Algorithm 3 Population-Based-HPO
1:Response function $f$
2:Population size $N$ and number of generations $T$
3:Initial distributions $\{p^{j}_{0}\}_{j\in[N]}$
4:for $k\leftarrow 1$ to $T$ do
5:     for $j\leftarrow 1$ to $N$ do $\triangleright$ Typically executed in parallel
6:         $\lambda_{k}^{j}\sim p^{j}_{k-1}$triangleright Draw an individual
7:         Evaluate $f(\lambda_{k}^{j})$triangleright Evaluate its fitness (if needed)
8:     end for
9:     Update $p^{j}_{k}$ from $p^{j}_{k-1}$ using $\{(\lambda^{i}_{k},f(\lambda^{i}_{k}))\}_{i\in[N]}$triangleright “Evolve” distributions
10:end for
Various implementations of ES/GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for a discussion on neural architecture search), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next sections, we review three particular implementations that have received notable attention.
6.2 Covariance matrix adaptation evolution strategy (CMA-ES)
CMA-ES [hansen1996adapting, hansen2016cma, loshchilov2016cma] maintains a global multivariate Gaussian distribution from which it samples individuals for the new generations. That is, the generation at iteration $k+1$ is drawn according to
$\lambda_{k+1}^{j}\sim p_{k}^{j}=\mu_{k}+\sigma_{k}\mathcal{N}(0,\Sigma_{k})\quad\text{for }j\in[N]$
where $N\in\mathbb{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size. Referring to Algorithm 3, this means in CMA-ES $p^{j}_{k}=\mathcal{N}(\mu_{k},\sigma_{k}^{2}\Sigma_{k})$ for all $j\in[N]$.
The parameters of the distribution are evolved at each iteration following a selection and recombination principle. Suppose $\{\lambda_{k}^{j}\}_{j=1}^{N}$ are sorted in increasing objective value, i.e. $f(\lambda_{k}^{1})\leq f(\lambda_{k}^{2})\leq\dots\leq f(\lambda_{k}^{N})$.  [^1] We recall we seek to minimize $f$, hence the first points are the best points. Then the mean is updated as a convex combination of the best $1\leq M\leq N$ points:
$\mu_{k+1}=\sum_{i=1}^{M}w_{i}\lambda_{k}^{i},$
where the weights $w_{i}$ are positive and sum to 1. The threshold $M$ and the recombination weights are parameter of the method. hansen2016cma suggests $M=N/2$ and $w_{i}\propto M-i+1$ as reasonable defaults.
The covariance matrix is adapted in a similar fashion by reestimating it from the $M$ best points of the previous generation. That is:
$\Sigma_{k}=\sum_{i=1}^{K}w_{i}(\lambda_{k}^{i}-\mu_{k})(\lambda_{k}^{i}-\mu_{k})^{\intercal}$ (6.1)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1); see [hansen2019pycma].
6.3 Population-based training (PBT)
Figure 6.1: Evolution of models trained with PBT for Atari game-playing policies. Red dots represent initial points in the parameter and hyperparameter spaces. Paths represents progressions via parameter training. Branching points represent applications of the exploit and explore operations, namely, a copy of the parameters and hyperparamters of better performing individuals onto worse performing one and a subsequent perturbation of their hyperparameter configurations. The performance score is color-coded, with light blue color representing better performing models. Reproduced from [jaderberg2017population].
jaderberg2017population propose a parallel method that maintains a population of model and hyperparameter pairs that are evolved as the models are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by gradient descent) with applications of evolutionary transformations to both the parameter of the models and the hyperparameter of the learning algorithm. PBT is an online hyperparamter optimization algorithm, as hyperparameters are updated alongside model parameters. In other words, in contrast to many techniques we discussed so far, PBT yields an hyperparameter sequence (or schedule) rather than a single hyperparamter configuration. 222 We discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Suppose one can “decompose” the learning algorithm of interest into a series of stages: $A=A_{T}\circ\dots\circ A_{1}$, where each map $A_{t}$ represents a stage of the learning process that can be warm-started given the outcome of the previous stage. This is particularly straightforward for neural network training: given a space of weights $\mathcal{W}$, each $A_{t}:\mathcal{W}\times\Lambda\to\mathcal{W}$ could be, for instance, an epoch of gradient-based optimization of the network weights (given configuration hyperparameters). Let $N\in\mathbb{N}$ be a population size and $\{(\theta_{0}^{j},\lambda_{0}^{j})\}_{j\in[N]}$ be a set of initial model weights and hyperparameters, where $\theta_{0}^{j}\in\mathcal{W}$ and $\lambda_{0}^{j}\in\Lambda$. These pairs represent the “individuals” in PBT and they can be initialized drawing from suitable distributions.
In PBT, generations correspond to stages of the learning algorithm. At each generation $t\in[T]$ the weights are updated as follows
$\tilde{\theta}_{t}^{j}=A_{t}(\theta_{t-1}^{j},\lambda_{t-1}^{j})\quad\text{for}j\in[N].$
Then, evolutionary transformations are applied to both parameters and hyperparameters before proceeding to the subsequent learning stage, on the basis of a fitness function $\mathscr{M}$ (e.g. a validation error) evaluated on the current model weights. Suppose we sort individuals in ascending order so that $\mathscr{M}(\tilde{\theta}_{t}^{1})\leq\mathscr{M}(\tilde{\theta}_{t}^{2})\leq\dots\leq\mathscr{M}(\tilde{\theta}_{t}^{N})$. jaderberg2017population propose two type of operations:
  • exploit, a selection strategy that involves copying the parameters and hyperparameters of better performing individuals onto worst performing ones; for instance, one may set $(\theta^{N}_{t},\lambda^{N}_{t})=(\tilde{\theta}^{1}_{t},\lambda^{1}_{t-1})$. jaderberg2017population suggest to replaces the worst-performing 20% with individuals randomly drawn from the best-performing 20%;
  • explore, a mutation transformation that brings over the parameter form the last learning stage, but either resamples the hyperparamter from the starting distribution, or perturbs them, e.g. setting $(\theta_{t}^{j},\lambda_{t}^{j})=(\tilde{\theta}_{t}^{j},\lambda_{t-1}^{j}+\varepsilon)$ for a suitable perturbation $\varepsilon$.
PBT further uses elitism, leaving the top performing individuals unchanged; that is, setting for some $M\in[N]$, $(\theta_{t}^{i},\lambda_{t}^{i})=(\tilde{\theta}_{t}^{i},\lambda_{t-1}^{i})$ for $i\in[M]$. Figure 6.1 present a visualization of a phylogenetic tree of game-playing policy functions for Atari games. The performance metric is the expected episodic reward, and the models deep neural networks trained using the Feudal Networks (FuN) framework [vezhnevets2017feudal].
tao2020learning combine PBT with gradient-based optimization of the hyperparameters, which we will discuss in the next chapter.
6.4 Particle swarm optimization
Particle swarm optimization (PSO) methods draws inspiration from the social behaviour of animals, like ant colonies or bird flocks [kennedy1995particle], as well as from the evolution of particle trajectories in a physical system. The general idea is that each candidate solution evolves according to a simple equation that weights a momentum term, or velocity, and one or more attractor terms to steer the trajectory toward better performing points.
Assume for simplicity that $\Lambda=\mathbb{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameter and velocity pairs, both drawn randomly from suitable distributions (e.g. uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization that uses the following equation:
$\lambda_{k+1}^{j}$ $=\lambda_{k}^{j}+v_{k}^{j}$ (6.2)
$v_{k+1}^{j}$ $=\alpha_{1}v_{k}+\alpha_{2}r_{l}\circ(\bar{\lambda}_{k}^{j}-\lambda_{k})+\alpha_{2}r_{g}\circ(\lambda^{*}-\lambda_{k})$ (6.3)
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathbb{R}^{m}$ are two randomly drawn vectors and $\circ$ is the Hadamard product. The configuration $\bar{\lambda}_{k}^{j}$ and $\lambda_{k}^{*}$ are, respectively, the particle’s best and population best points found so far, that is,
$\bar{\lambda}_{k}^{j}=\operatorname*{arg\,min}_{\lambda\in\bar{\Lambda}^{j}_{k}}f(\lambda)\quad\text{for }j\in[N],$
where $\bar{\Lambda}^{j}_{k}=\left\{\lambda^{j}_{t}\,:\,t\in[k-1]\right\}$, and
$\lambda^{*}_{k}=\operatorname*{arg\,min}_{\lambda\in\Lambda^{*}_{k}}f(\lambda),$
where $\Lambda^{*}_{k}=\left\{\lambda_{t}^{j}\,:\,t\in[k-1],\,j\in[N]\right\}$. The second term in the velocity update of Eq. (6.3) can be interpreted as a local memory attractor that pushes the current configuration toward the individual best configuration found so far, while the third term is a global attractor that pushes toward the global best configuration found so far. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search. Referring to Algorithm 3, each individual distribution is implicitly modelled by the Equations (6.2)-(6.3).
6.5 Discussion
Despite their simplicity and their lack of theoretical guarantees, population-based methods enjoy broad applicability and strong empirical performances. In addition, this class of methods is designed to readily scale to parallel and distributed compute systems. Although traditional EA and PSO implementation assume synchronicity, one may easily derive asynchronous versions of many algorithms. For instance in PSO one may replace the global best $\lambda^{*}$ with the best known solution to individual $j$. However, population-based algorithms require a considerable amount of customization due to the presence of several method parameters, task-specific heuristics and design choices, e.g. to define meaningful mutation or perturbation strategies, undermining their accessibility and applicability to novel algorithms and problems. Besides, they typically require a considerable computational power, benefiting usually from large population sizes.
Finally, we note that one may consider successive halving and related methods, which we covered in Chapter 5, as also falling in the category of population-based methods. However, the foundations of those methods are linked to the bandit problem and to hedging strategies [li2017hyperband], rather than taking inspiration from evolutionary principles.
Chapter 7 Gradient-based Optimization
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Assuming that $f(\lambda)$ be differentiable excludes the direct optimization of discrete hyperparameters such as discrete choices (for example, which nonlinear activation function should be used in a layer of a neural network) or integer-valued hyperparameters (such as the dimension of a layer). There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. For example, maclaurin2015gradient suggested to consider the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning, which we expand on in Section 8.4). For example, the idea of tuning initial weights from several learning tasks was shortly after successfully exploited in [finn_model-agnostic_2017] in their model-agnostic meta learner.
The crux of gradient-based HPO lies in computing approximate gradients of the response function w.r.t. hyperparameters. In this chapter, we cover three main technical approaches: implicit differentiation (Section 7.2), iterative differentiation (Section 7.3) and fixed-point iterations (Section 7.4). Next, we explore some use cases in Section 7.5. We touch on extensions that allow for the gradient-based optimization of discrete hyperparameters in Section 7.6, and conclude with a discussion centred on computational considerations in Section 7.7.
7.1 Preliminaries
At a high level, gradient-based HPO methods start from an initial hyperparameter configuration $\lambda_{0}\in\Lambda$ and iteratively and sequentially update such configuration as follows:
$\lambda_{r}=\mathrm{Proj}_{\Lambda}\left[\lambda_{r-1}-\nu\hat{\nabla}f(\lambda_{r-1})\right]\qquad r=1,2,\dots$ (7.1)
where ^$\hat{\nabla}f$ informally denotes some approximation of the gradient of the HPO objective $f$, $\nu>0$ is a step-size and $\mathrm{Proj}_{\Lambda}$ is the projection onto $\Lambda$. Gradient-based optimization of the loss function inside the learning algorithm $A$ is extremely common, especially in deep learning. Hence, to avoid confusion, we will call ^$\hat{\nabla}f$ the hypergradient.
Not every HPO problem is suitable for gradient-based optimization. Recalling the general formulation of the HPO problem of Section 2.2, one underlying assumption shared by all gradient-based HPO methods is that the performance metric $\mathscr{M}$ be differentiable w.r.t. the model output of $A$. In this chapter, to simplify the notation we take the output of $A$ to be a parameter vector in some weight space $\mathcal{W}\subseteq\mathbb{R}^{d}$ and let the performance metric take values in the weight space directly, that is $\mathscr{M}:\mathcal{W}\times\mathscr{D}\ \to\mathbb{R}$. Hence
$f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})=\mathscr{M}(A(\mathcal{D},\lambda),\mathcal{V}).$ (7.2)
Hereafter, we will drop training and validation sets.
In order to compute hypergradients, it is also necessary to specify at least some details of the learning algorithm $A(\mathcal{D},\lambda)$ that appears in (2.4) and that we have mostly treated as an opaque box until now. Gradient-based HPO methods are distinguished mainly by the assumptions they make about $A$, which then translate into different computational schemes for ^$\hat{\nabla}f$. Next, we review the main approaches.
7.2 Hypergradients via implicit differentiation
Many learning algorithms may be conceptually expressed as solutions to optimization problems
$A(\mathcal{D},\lambda)\in\operatorname*{arg\,min}_{u\in\mathcal{W}}\mathcal{J}(u,\lambda),$ (7.3)
where $\mathcal{J}:\mathcal{W}\times\Lambda\to\mathbb{R}$, with $\mathcal{W}\subseteq\mathbb{R}^{d}$ is an objective function (the training error) that depends on the data – we have seen a concrete example in (2.3).  [^1] We leave dependence on data implicit for improving readability. If $\mathcal{J}$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
$\nabla_{u}\mathcal{J}(u,\lambda)=0.$ (7.4)
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $\mathcal{J}$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}\mathcal{J}\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [krantz2012implicit] to (7.4) which assures us that (locally) there exist a function $w:U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. In such case, we can locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
$[\nabla f(\lambda)]^{\intercal}=\partial_{\lambda}f(\lambda)=\partial_{w}\mathscr{M}(w(\lambda),\lambda)\partial_{\lambda}w(\lambda)\qquad\forall\lambda\in U,$ (7.5)
where, again due to the implicit function theorem,
$\partial_{\lambda}w(\lambda)=-\left[\partial_{u}^{2}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}\right]^{-1}\partial^{2}_{u,\lambda}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}.$ (7.6)
Equations (7.5) and (7.6) give us the analytic expression of the exact gradient of the HPO objective $f(\lambda)$. Such expression is, however, only implicit. In fact, we almost never have an explicit formula for the map $w(\lambda)$.
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging (7.6) into (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}\mathcal{J}\right]^{-1}\in\mathds{R}^{d}$ and approximately solve the linear system
$\partial^{2}_{u}\mathcal{J}(u,\lambda)_{|u=w_{T}}q=-\nabla_{u}\mathscr{M}(u)_{|u=w_{T}}.$ (7.7)
Call $q_{T,K}$ an approximate solution of (7.7). Then an approximate gradient of the outer objective is given by
$\nabla_{T,K}f(\lambda)=\left[\partial^{2}_{u,\lambda}\mathcal{J}(u,\lambda)_{|u=w_{T}}\right]^{\intercal}q_{T,K}.$ (7.8)
The two possible sources of approximations stem from ($i$) the solution of the inner problem and ($ii$) of the linear system (7.7). Particular implementation of this computation scheme depend on how (7.7) is solved. One classic procedure is the conjugate gradient (CG) method [hestenes1952methods], which features a linear rate of convergence that depends on the conditioning number of the Hessian of the inner objective.
This computational approach has been studied in [larsen_design_1996, larsen1998adaptive, bengio_gradient-based_2000, chapelle2002choosing, seeger2007cross]. More recently, pedregosa_hyperparameter_2016 proposed an approximation scheme whereby solutions to the inner problems and to the linear equation (7.7) become more precise as the number of HPO iterations $r$ grows, proving convergence for algorithms which feature strongly convex (inner) objectives.
7.3 Hypergradients via iterative differentiation
Another approach to computing hypergradients involves explicitly accounting for an iterative training dynamics. Methods based on iterative differentiation further assume that solutions to the learning algorithm problem (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
$s_{t}=\Phi_{t}(s_{t-1},\lambda,\mathcal{D})\ \ t=1,\dots,T.$ (7.9)
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathds{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$ and a (possibly empty) set of auxiliary variables the $v_{t}$ – e.g. velocities and moments for algorithms such as Adam and RMSProp. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition maps $\Phi_{t}:(\mathds{R}^{d}\times\mathds{R}^{m}\times\mathscr{D})\rightarrow\mathds{R}^{d}$ are assumed to be differentiable and describe the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g. for stochastic gradient descent) and it internally requires the computation of the inner objective gradients
$g_{t}\doteq\nabla_{w}\bigl{.}\mathcal{J}(w,\lambda,\mathcal{D}_{t})\bigr{|}_{w=w_{t-1}}.$
Note that $\Phi$ depends both explicitly and via $s_{t-1}$, and $\lambda$ might contain parameters related to the optimizer itself such as the learning rate. As an example, the dynamics (7.9) for stochastic gradient descent with momentum are
$\begin{array}[]{lcl}v_{t}&=&\mu v_{t-1}+g_{t}\\ w_{t}&=&w_{t-1}-\eta v_{t-1}\end{array}$ (7.10)
where $\mu$ and $\eta$ are the momentum term and the learning rate, and may be included in $\lambda$ if desired.
In a nutshell, iterative differentiation schemes for gradient-based HPO rely on the application of automatic differentiation (AD) to the composite map $\mathscr{M}\circ\Phi_{T}\circ\dots\circ\Phi_{1}$. We briefly survey AD next, and refer the reader to [Griewank08:_evaluat_deriv] for an in-depth account.
7.3.1 Automatic differentiation
Automatic (or algorithmic) differentiation (AD) [Griewank08:_evaluat_deriv, baydin2017automatic] is the standard tool for computing gradients in deep learning. Suppose we have a generic differentiable function $G:\mathds{R}^{N}\mapsto\mathds{R}^{M}$ with $(\eta_{1},\dots,\eta_{M})=G(\xi_{1},\dots,\xi_{M})$ and we are interested in computing the Jacobian matrix, with entries $\frac{\partial\eta_{i}}{\partial\xi_{j}}$. In AD, we begin by describing $G$ as a computational graph $(V,E)$, which is a directed acyclic graph whose nodes $V$ include inputs $\xi_{j}$, outputs $\eta_{i}$, and intermediate variables that are computed by applying elementary functions to inputs or other intermediate variables, that can be scalars, vectors or tensors. Elementary functions are typically atomic operations such as (tensor) sum, (Hadamard) product, and (elementwise) trascendental functions, whose partial derivatives are readily known. Edges $E$ in the graph describe functional dependencies, i.e., they link intermediate variables to the elementary functions they are fed into. Forward propagation, following a topological sort of the graph, allows us to compute $G(\xi_{1},\dots,\xi_{M})$. Two AD strategies are available, forward and reverse mode, which are based on the definition of local derivatives and a message-passing procedure based on the standard chain rule of calculus.
In forward mode AD, the local derivative of $v_{k}$ is defined as
$\dot{v}_{k}\doteq\frac{\partial v_{k}}{\partial\xi_{j}}$
and can be updated by the recursion
$\dot{v}_{k}=\sum_{\ell\in\pi_{k}}\dot{v}_{\ell}\frac{\partial v_{k}}{\partial v_{\ell}}$ (7.11)
initialized as $\dot{\xi_{j}}=1$, where $\pi_{k}$ are $v_{k}$’s parents, and iterated following a topological sort of $(V,E)$. In the above equation, if $v_{k}$ and $v_{\ell}$ are vector-valued, $\frac{\partial v_{k}}{\partial v_{\ell}}$ denotes a Jacobian matrix. The desired derivatives can be finally read as $\dot{\eta}_{i}$. The same propagation allows us to compute the $j$-th column of the Jacobian and to obtain the whole Jacobian, propagation needs to be repeated for each input $\xi_{j}$. Assuming elementary functions can be computed in $O(1)$, the total computational cost is $O(NE)$. Forward mode is thus advantageous for functions with few inputs and many outputs. If on the other hand $N=O(E)$, as it happens for example for the objective function of neural networks, forward mode is as inefficient as computing derivatives numerically. It is therefore not surprising that forward mode is never used to train neural network models (although there is a remarkable exception when training recurrent networks on data streams [williams_learning_1989], a setting where reverse mode cannot be applied).
In reverse mode (the strategy used by the well known backpropagation algorithm [rumelhart1986learning]) the local derivative of $v_{k}$, usually called adjoint, is defined as
$\bar{v}_{k}\doteq\frac{\partial\eta_{i}}{\partial v_{k}}$
and can be updated by the backward recursion
$\bar{v}_{k}=\sum_{\ell\in\gamma_{k}}\bar{v}_{\ell}\frac{\partial v_{\ell}}{\partial v_{k}}$ (7.12)
initialized as $\bar{\eta_{i}}=1$, where $\gamma_{k}$ are $v_{k}$’s children, and iterated following a reversed topological sort of $(V,E)$. The desired derivatives can be finally read as $\bar{\xi}_{j}$. The same propagation allows us to compute the $i$-th row of the Jacobian and to obtain the whole Jacobian, propagation needs to be repeated for each output $\eta_{i}$, for a total computational cost of $O(ME)$. Reverse mode is thus advantageous for functions with many inputs and few outputs. However, since the computation proceeds in a backward way, the memory used to store the values of intermediate nodes $v_{k}$ cannot be freed until the end of the propagation, yielding a memory cost of $O(V)$.
7.3.2 Computing hypergradients
We are now ready to apply the principles of AD to the differentiation of the validation error with respect to $\lambda$. A simplified view of the computational graph mapping $\lambda$ into the response $f(\lambda)$ is shown in Figure 7.1, where minibatches $\mathcal{D}_{t}$ and initial weights $w_{0}$ are supplied as given constants [^2] Note that $w_{0}$ could be equally well included in $\lambda$ but here we keep it distinct for clarity.. Here the number of inputs $M$ is the dimension of the hyperparameter vector and there is a scalar ($N=1$) output. Forward and reverse calculations are presented in [franceschi2017forward] and repeated below.
Figure 7.1: Computational graph of the response function for forward and reverse mode hypergradient calculation.
In the case of forward mode, the local variable is $\dot{s}_{t}$ and is udpated by the recursion
$\dot{s}_{t}\doteq\frac{\partial s_{t}}{\partial\lambda}=\left.\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}\dot{s}_{t-1}\right|_{a=s_{t-1}}+\left.\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}\right|_{b=\lambda}$ (7.13)
where $\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}$ and $\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}$ are $d\times d$ and $d\times m$ Jacobian matrices, respectively. The base step is $\dot{s}_{0}=\frac{\partial s_{0}}{\partial\lambda}$ which is zero unless the hyperparameters directly affect the initial weights (e.g., their range, if randomly initialized, or even their values as in maclaurin2015gradient). The desired gradient is finally obtained at the last node:
$\nabla f(\lambda)=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}\dot{s}_{T}$
where $w_{T}$ is the proper slice of $s_{T}$.
In the case of reverse mode, the adjoint $\bar{s}_{t}$ is initialized as
$\bar{s}_{T}=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}$
and updated as
$\bar{s}_{t}\doteq\frac{\partial f}{\partial s_{t}}=\left.\bar{s}_{t+1}\frac{\partial\Phi(a,\lambda,\mathcal{D}_{t})}{\partial a}\right|_{a=s_{t-1}}.$
Since the variable $\lambda$ is a parent of all $s_{t}$, Eq. (7.12) requires us to sum all the contributions:
$\nabla f(\lambda)=\sum_{t=1}^{T}\bar{s}_{t}\left.\frac{\partial\Phi(s_{t-1},b,\mathcal{D}_{t})}{\partial b}\right|_{b=\lambda}.$ (7.14)
It must be remarked the striking similarity between hypergradient calculations and gradient calculations in recurrent neural networks (RNN). Indeed, the computational graphs are identical, where $s_{t}$ matches the “state” of an RNN, $\lambda$ the weights, and $\Phi$ the state update function. Equations (7.13) and (7.14) closely resemble real-time recurrent learning [williams_learning_1989] and backpropagation through time [werbos1990backpropagation], respectively.
7.4 Fixed-point methods
Another possible approach that sits in between the iterative and the implicit differentiation schema is given by the so-called fixed-point method. The method is based on the observation that one can rewrite the condition (7.4) as a fixed point equation of an appropriate contractive dynamics
$w(\lambda)=\Phi(w(\lambda),\lambda),$ (7.15)
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla\mathcal{J}$.  [^3] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to (7.15) yields to
$\frac{dw(\lambda)}{d\lambda}=\left(I-\frac{\partial\Phi(w(\lambda),\lambda)}{\partial w}\right)^{-1}\frac{\partial\Phi(w(\lambda),\lambda)}{\partial w\partial\lambda}$
and, repeating the procedure described above, Equation (7.7) becomes
$\left(I-\partial_{w}\Phi(w_{T},\lambda)\right)^{\intercal}q=\nabla\mathscr{M}(w_{T}).$ (7.16)
Starting from any point $q_{T,0}\in\mathds{R}^{d}$, one can iterate
$q_{T,k}=\partial_{w}\Phi(w_{T},\lambda)^{\intercal}q_{T,k-1}+\nabla\mathscr{M}(w_{T})\qquad k\in[K]$ (7.17)
to find an approximate fixed-point of (7.16) and then plug $q_{T,K}$ into Equation (7.8) to obtain an hypergradient.
This computational scheme emerges in more general cases where $\Phi$ is not necessarily an optimization dynamics: $\Phi$ may represent a stable recurrent neural network [miller2019stable], a graph neural network [scarselli2009graph] or an equilibrium model [bai2019deep]. In these contexts, the fixed-point algorithm is usually known as recurrent backpropagation [almeida1987learning, liao2018reviving].
Warm-starting $q_{T,0}$ with the vector obtained at the previous optimization iteration (of $\lambda$) may speed-up the convergence of (7.17).
7.5 Use cases
Gradient-based approaches have mostly found application in academic contexts. Their flexibility, however, allows us to tune very high-dimensional hyperparameter vectors, as showcased in several papers.
One possibility is to attach a hyperparameter to every weight in a deep network. We have already mentioned the idea proposed by maclaurin2015gradient of regarding the initial values for the weights as hyperparameters. In the same paper, the authors also suggest a closely related opportunity where a regularization penalty is associated to every individual weights. The meta-learning technique presented in [franceschi2018bilevel] can be interpreted as treating all weights in the first layers of a deep network as hyperparameters, and optimize the loss on new tasks with respect to the weights in the classification head.
Another direction that has been studied is to attach hyperparameters to data. maclaurin2015gradient demonstrated an extreme experiments where image pixels themselves are hyperparameters. In a more realistic scenario, hypergradients have been used to tune the hyperparameters of differentiable data augmentation procedures, yielding significantly speedups without performance loss compared to other data augmentation searching methods [lin:2019:online_hyperparameter_learning, hataya:2020:faster_autoaugment_learning]. A different opportunity is to assign a cost weight to each example in a supervised learning setting so that the overall loss for a training set of $n$ examples is $\sum_{i=1}^{n}\lambda^{(i)}\mathcal{J}(h(x^{(i)}),y^{(i)})$. Here we have $n$ hyperparameters $\lambda^{(i)}$. franceschi2017forward show how this idea can be exploited to filter out from the training set examples with noisy (or perhaps poisoned) labels. In this regards, it is interesting to note that (although not directly related to HPO) the problem of building training set attacks on linear models can be formulated in terms of bilevel optimization [mei:2015:using_machine_teaching] and has commonalities with gradient-based HPO in its mathematical structure.
We finally note here that tuning the whole learning rate schedule has also been tackled with hypergradient approaches [baydin_2017_online, donini2019opt, micaelli2021gradient, sharifnassab:2024:metaoptimize_framework_optimizing]. This specific problem is covered in more detail in Section 8.2.
7.6 Probabilistic reparameterization of discrete hyperparameters
The presence of discrete hyperparameters into the response function – and therefore into the training objective or dynamics – prevents a straightforward application of gradient-based HPO techniques. For some such hyperparameters, however, it might make sense to reformulate them as random variable that follow some discrete distribution. For instance, consider a discrete binary choice $\lambda\in\{0,1\}$, e.g. indicating whether to use an $L^{1}$ or $L^{2}$ regularizer in (2.3). Then, one may define a discrete binary random variable $\iota$ as follows:
$\iota\sim\mathrm{Ber}(\theta),\qquad\theta\in[0,1]$ (7.18)
and reinterpret $\lambda$ as a drawing of $\iota$, where $\mathrm{Ber}(\theta)$ is the Bernoulli distribution with success parameter $\theta$. 444 Similarly, one may reformulate hyperparmeters encoding multiple choices via categorical distributions. Crucially, the distribution parameter $\theta$ is now continuous and one may consider optimizing the response function
$g(\theta)=\mathbb{E}_{\iota\sim\mathrm{Ber}(\theta)}[f(\iota)]=\theta f(1)+(1-\theta)f(0)$ (7.19)
instead of $f$, having that $\min_{\theta}g(\theta)=\min_{\lambda}f(\lambda)$. Although (7.19) might look like a circular argument, the value of such reformulation is unlocked when one computes (approximate) hypergradients via appropriate continuous relaxation or discrete gradient estimators [niculae2023discrete]. These techniques may be combined with the procedures described in this chapter, making it possible to avoid the exponential cost of computing $f$ on all the configuration. This is still an active area of research: successful examples include the optimization of graph structures in transductive learning [franceschi2019learning] and topological order of nodes in DAG learning [zantedeschi2023dag], deep learning architectures [liu2018darts] and dropout coefficients [gal2017concrete].
7.7 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g. for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems. A comparative theoretical study between the three main approaches covered in this chapter is presented in [grazzi2020iteration].
Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to completely avoid the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to $(j+1)C-1$ to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered in [shaban2019truncated], together with a more radical truncation strategy similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90], which however introduces further bias in the computation of hypergradients.
Chapter 8 Further Topics
In the last five chapters we covered the main classes of algorithms for hyperparamter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization (Section 8.1), online HPO and hyperparameter schedules (Section 8.2) and neural architecture search (Section 8.3). Next, we dedicate Section 8.4 to meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting. We conclude the chapter with a brief discussion on transfer of hyperparamters across model sizes, a novel topic that is relevant especially for the tuning of large foundational models (Section 8.5).
8.1 Optimizing for multiple objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example, when working with hardware-constrained systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective, it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front—the set of all non-dominated solutions—to help users make an informed trade-off decision. More formally, consider a function $f:\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be an $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda_{1}$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination. Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally:
$\mathcal{P}_{f}=\{\lambda\in\Lambda\,|\,\nexists\lambda^{\prime}\in\Lambda:\lambda^{\prime}\succ\lambda\}.$
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $B\subset\mathcal{Y}$ of objective vectors near the Pareto front. The quality of $B$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $\mathbf{r}$, $\mathcal{H}(B)$ equals the hypervolume dominated by $B$:
$\mathcal{H}(B)=\text{Vol}\left(\{\mathbf{y}\in\mathcal{Y}=\mathbb{R}^{n}\,|\,\exists\mathbf{y}^{\prime}\in B:\mathbf{y}^{\prime}\succeq\mathbf{y}\wedge\mathbf{y}\succeq\mathbf{r}\}\right).$
Computing $\mathcal{H}(B)$ involves partitioning the space into hypercubes, an operation that scales exponentially with the number of objectives $n$. This exponential scaling poses a key bottleneck for hypervolume-based multi-objective optimizers.
Examples
A popular example of an MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to as acceptance-based and error-based, can conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some extent and expose empirical trade-offs between various definitions can still be important from societal perspectives. Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a surrogate model as discussed in Chapter 4. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.
8.1.1 Multi-objective optimization
Initial efforts towards MO optimization for HPO [Jin2006:Multi] were inspired by evolutionary techniques [Yao1999:Evolving, Nolfi2002:Evolution, Igel2005:Evolutionary, friedrichs2005evolutionary]. To mitigate the large computational cost of evaluating a single hyperparameter configuration (e.g., training a DNN), sample-efficient Bayesian optimization (BO) techniques have been introduced [hutter_sequential_2011, snoek_practical_2012, bergstra2011algorithms].
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective BO techniques. The scalarization is defined by using a weight vector $\mathbf{w}\in\mathbb{R}^{n}$.
Some common choices for $V$ are:
  • Inner product between objective and weight vectors (also known as random weights): $V_{\text{RW}}(f(\lambda),\mathbf{w})=f(\lambda)^{\top}\mathbf{w}$;
  • The scalarization used in [Knowles2006:Parego], called ParEGO:
    $V_{\text{ParEGO}}(f(\lambda),\mathbf{w})=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}\mathbf{w});$
  • The scalarization used in [Golovin2020]:
    $V_{\text{Golovin}}(f(\lambda),\mathbf{w})=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}.$
Note that by optimizing different scalarizations, solutions along the Pareto front can be obtained. The choice of $V$ influences the subset of the Pareto front that is found, and the right selection might change depending on the task at hand.
A second class of MBO techniques builds on a performance measure of Pareto front approximations, namely the dominated hypervolume [Zitzler1998:Multiobjective]. Both the expected hypervolume improvement (EHI) [Emmerich2011:Hypervolume] and probability hyper-improvement (PHI) [Keane2006:Statistical] operate by extending their single-objective BO counterparts—expected improvement [Mockus1978:Application, Jones1998:Efficient] and probability of improvement [Kushner1964:New], respectively. Other methods include step-wise uncertainty reduction (SUR) [Picheny2015:Multiobjective], smsEGO [Ponweiser2008:Multiobjective, Wagner2010:Expected] and expected maximin improvement (EMMI) [Svenson2010:Multiobjective].
Finally, information-theoretic MBO approaches aim to select points that reduce uncertainty about the location of the Pareto front. Methods in this class tend to be more sample efficient and scale better with the number of objectives. PAL [Zuluaga2013:Active] iteratively reduces the size of a discrete uncertainty set. PESMO [Hernandez2016:Predictive] adapts the predictive entropy search (PES) [Hernandez2014:Predictive] criterion. Pareto-frontier entropy search (PFES) [Suzuki2019:Multi] is suitable when dealing with decoupled objectives. MESMO [Belakaria2019:Max] builds on the max-value entropy-search criterion [Wang2017:Max] and enjoys an asymptotic regret bound. Building on MESMO, two recent works have proposed MF-OSEMO [Belakaria2020:Multi] and iMOCA [Belakaria2020:Information], two multi-fidelity based information-theoretic MBO techniques which internally use multi-fidelity Gaussian processes.
8.1.2 Constrained hyperparameter optimization
When optimizing for multiple targets, some objectives may have predefined constraints on their values. For example, one might seek to optimize a given objective function while enforcing a strict memory constraint. In this section, we focus on this scenario.
Formally, the goal in a constrained optimization scenario is to minimize the target response function $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda):\Lambda\rightarrow\mathbb{R}$ and $c(x):\Lambda\rightarrow\mathbb{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$\min_{\{\lambda\in\Lambda\,:\,c(\lambda)\leq\delta\}}f(\lambda),$ (8.1)
where $\delta\in\mathbb{R}$ is a threshold on the constraint. For instance, in the case of BO, the latent functions $f(\lambda)$ and $c(\lambda)$ can be conditionally independent in our surrogate model, with different GP priors placed on them.
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assumes that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent function can be represented as GPs with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a DNN fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash. Rather, we only know that an out-of-memory error occurred, which is a binary outcome. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.1) as follows:
$y_{\star}=\min_{\lambda\in\Lambda}\left\{y(\lambda)\;\|\;P(z_{c}=+1|\lambda)=\sigma(c(\lambda))\leq\sigma(\delta)\right\}.$
This formulation is similar to the one proposed in [Gelbart14]. The parameter $\sigma(\delta)\in(0,1)$ controls the size of the (random) feasible region for defining $y_{\star}$. Finally, note that in the example of out of memory errors training failures, the criterion observation $z_{y}$ is obtained only if $z_{c}=-1$: if a training run crashes, a validation error is not obtained for the queried configuration. Apart from an experiment in [Gelbart14], the case of contrained BO work covering the binary feedback case is covered in [perrone_cmes].
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. A separate regression model is used to learn the constraint function $c(\lambda)$ (typically a GP), and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=\mathbb{P}\{c(\lambda)\leq\delta\}EI(\lambda)$, where $\mathbb{P}\{c(\lambda)\leq\delta\}$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.3.
Several issues with cEI are detailed in [Lobato15a]. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter features in [Gelbart14], but is only used to recommend the final hyperparameter configuration. Another approach was proposed in [Lobato15a], where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case are developed in [Picheny2016] and [Ariafar2019]. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. An alternative option that does not require expensive computations is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
8.2 Hyperparameter schedules and online HPO
The vast majority of learning algorithms do not produce an hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapters we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.3). When this is the case, it may be sensible to consider hyperparameter schedules rather than fixed values. One of the most prominent example in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A:\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}:\mathcal{H}\times\Lambda\times\rightarrow\mathcal{H}$. 111 We disregard the dependence of the maps on data, for notational simplicity. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\lambda)=\Phi_{T}(\cdot,\lambda_{T})\circ\Phi_{T-1}(\cdot,\lambda_{T-1})\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1}).$ (8.2)
for some $T>0$, where we generally view $\lambda\in\Lambda^{T}$ as an hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other sequence of operations.
Hyperparameter iterates may be left “free to float”, in which case one can formulate a related HPO problem following the general setup of Section 2.2, but this time searching for configurations in $\Lambda^{T}$ rather than in $\Lambda$. This is seldom done in practice due to the complications arising from an expanded search space, few exceptions being [maclaurin2015gradient, wu2018learning]. More commonly, the iterates are expressed as a function of the step $t$ and a set of other fixed hyperparameters $\lambda^{\prime}\in\Lambda^{\prime}$ Learning rate schedules are a primary example. For instance, a triangular schedule (i.e. warmup followed by linear decay) is given by:
$\eta_{t}=\left\{\begin{array}[]{ll}\eta_{\mathrm{max}}(t/T_{\mathrm{max}})&\text{ for }\,t\geq T_{\mathrm{max}}\\ \eta_{\mathrm{max}}(1-(t-T_{\mathrm{max}})/(T-T_{\mathrm{max}}))&\text{ for }\,T_{\mathrm{max}}<t\leq T\end{array}\right.$
where $\lambda^{\prime}=(\eta_{\mathrm{max}},T_{\mathrm{max}})$. Once again, when defining a related HPO problem, one may fall back to the setting of Section 2.2, optimizing over $\Lambda^{\prime}$ rather than $\Lambda$. Along these lines, one may even let the hyperparameter iterates be a function of a state (e.g. the weights of the trainee model at step $t$), an approach that has been explored e.g. by hansen2016using using reinforcement learning to fit a learning rate adjustment policy or lorraine2018stochastic using hyper-networks and gradient-based HPO.
Online HPO.
However, perhaps one main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparamter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors.
Another stream of recent research focuses on tuning hyperparameters of online learning algorithms, such as contextual bandit [li2010contextual] methods like UCB [auer2002finite] or Thomson Sampling [agrawal2013thompson]. In these algorithms, hyperparameters affect the agent’s exploration-exploitation trade-off and must adapt in real time. bouneffouf2020hyper treat this as a nested bandit problem, tuning exploration parameters online via meta-bandits. ding2022syndicated generalize this idea with Syndicated Bandits, a multi-layer optimization scheme that tunes several hyperparameters simultaneously, with favorable regret guarantees. kang2024online further extend this line by framing tuning as a continuum-armed bandit problem, enabling efficient search in continuous spaces. These methods depart from gradient-based HPO discussed above by tightly coupling hyperparameter adaptation with decision-time feedback, offering principled and scalable alternatives tailored to the online nature of bandit learning.
Online HPO remains an area of active research.
8.3 Neural architecture search
Deep learning has revolutionized the traditional approach of hand-crafting features by enabling the automatic learning of features during training. However, the process of creating neural network architectures still depends significantly on manual design, which involves a substantial amount of trial and error. We can extend the idea of hyperparameter optimization to automatically design neural network architectures, a process known as neural architecture search (NAS).
NAS shares similarities with hyperparameter optimization (HPO), but focuses solely on the hyperparameters that define the architecture. However, NAS faces challenges distinct from most HPO problems, as its search spaces are often significantly higher dimensional, exclusively discrete, and of highly conditional structure (e.g., the hyperparameter controlling the number of units in a given layer is only applicable if that layer exists, as controlled by another hyperparameter).
8.3.1 Search Space
A key question in NAS is how to define a search space to parameterize a neural network architecture. An expressive search space might lead to entirely new and innovate architecture but might be also prohibitively expensive to explore. On the other hand, a too restrictive space will lower the chance to improve upon manually design architectures. We discuss two popular approaches to define NAS search spaces [white-arxiv23]: Marco Search Spaces and Cell Search Spaces.
Macro search spaces.
We assume a pre-defined architecture topology, such as for example residual neural network or a transformer architecture. Macro search spaces consider design decisions to configure the architecture on a macro level, for example whether to use pre- or post-layer normalization or how many heads to use for all multi-head attention layers.
For example, MNASNet [tan-cvpr19] separates an architecture into blocks and for each block optimizes different hyperparameters, such as the convolution type or kernel size. NASBOT [kandasamy-neurips18] defines the entire network as a DAG where nodes perform a certain operation and edges represent direct dependencies between nodes, or in other words, the output of one node is the input of an adjacent node. For each node, the NAS algorithm has to select the operation type and the number of convolutional filters or fully connected nodes.
The advantage of macro search spaces is that they allow parameterizing each block or layer of the architecture individually. However, while this leads to a very expressive space, it will be quite expensive to search.
Cell search space.
Compared to the Macro Search Space, the Cell Search Space considers a block or cell which is a smaller computational unit that is stacked multiple times on top of each other to form the final architecture. zoph:2018:learning_transferable defined each cell as a directed acyclic graph where for each node they select a certain operation to be performed, for example 1x1 convolution or 3x3 convolution and how to place the edges in the graph. To reduce the overall complexity, zoph2016neural distinguished between two different types of cell, which share the same structure but different strides to reduce the spatial dimension of the input.
Instead of assigning operations to each node, liu2018darts assigns operations to every edge of the graph. Nodes that accumulate the feature vectors from all incoming edges [did not understand]. Their proposed method DARTS then uses gradient descent to select an operation for each edge.
ying-icml19 proposed NASBench-101 which consists of a cell search space with a maximum of 7 nodes and 3 operation choices per node. Furthermore, they limited the total number of edges to 9, which led to a total of 400K different architectures. In an expensive offline step, all architectures were evaluated on CIFAR-10 to form a table. This allows running NAS methods by simply looking up the performance of an architecture instead of training it, accelerating the search process by multiple orders of magnitude.
In follow-up work,  dong-iclr20 proposed NASBench-201 which reduced the number of nodes but allowed for more operations per node. While the total number of architectures in this search space is substantially smaller than in NASBench-101, there are no constraints on the total number of edges, which makes it more accessible for different search algorithms. Furthermore, NASBench-201 evaluated each architecture on CIFAR-10, CIFAR-100 and Imagenet-16.
8.3.2 Search strategies
How do we explore the search space? This is particularly challenging in the case of NAS because of the large number of possible architectures in the search space and the expensive nature of training neural network models.
In one of the first works on NAS, zoph2016neural used a controller based on a recurrent neural network to suggest new architectures. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. real2019regularized proposed an evolutionary approach that maintains a population of architectures (see Chapter 6). In each iteration, a random architecture is drawn from the population and mutated along a single dimension. While these approaches led to architectures outperforming state-of-the-art models at that point, the original methods required substantial computational resources since each network was trained from scratch.
To accelerate the search process, we can utilize recent advancements from HPO, such as multi-fidelity optimization. For example, there is a body of work [white-neurips21] that tries to predict learning curves of neural networks to stop the training process early.
A more radically different approach to accelerate NAS is so-called weight-sharing NAS. The idea is to define a single super-network that contains all possible architectures in the finite search space. Each path through the super-network represents a single architecture. In the original work by pham-icml18, this method accelerates the search by more than a factor of 1000.
In a similar vein, liu2018darts proposed DARTS, which adds a continuous architecture parameter for every operation at each edge in the super-network. These architecture parameters are jointly optimized with the weights of the neural network in a bi-level fashion using variations of algorithms we described in Chapter 7. The inner loop fits the weights of the super-network to minimize the training loss, and the outer loop optimizes the architecture parameters to minimize the validation loss. After training the super-network, the final architecture is obtained by selecting the $k$-most important operations per edge. This architecture is then usually re-trained from scratch.
Another approach is so-called two-stage NAS [cai-iclr20, yu-eccv20], which first trains the super-network and then runs gradient-free optimization methods, such as random search or evolutionary algorithms, to select the optimal sub-network without any further training. For example, cai-iclr20 proposed the once-for-all model, which provides a single super-network that allows selecting the optimal architecture with weights for a given hardware, minimizing both latency on the target device and validation error.
8.4 Meta-learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to HPO. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta learning, see e.g., [hospedales:2021:metalearning_neural_networks] and references therein.
8.4.1 Definition
In meta-learning (also called learning-to-learn), we have a sequence of tasks $\mathcal{T}^{1},\mathcal{T}^{2},\dots,\mathcal{T}^{N}$, where each task $\mathcal{T}^{j}=(p^{j},\mathcal{L}^{j})$ consists of a distribution $p^{j}(x,y)$, $x\in\mathcal{X}$, $y\in\mathcal{Y}^{j}$ over data points and a loss function $\mathcal{L}^{j}$. Note that all tasks share the same input space but their label spaces may differ. Tasks are sampled from a meta distribution $\mathcal{P}$ on tasks $(\mathcal{T})$. Unlike a traditional supervised algorithm (which, given a dataset, will produce a prediction function $h$), a meta-learning algorithm outputs a learning algorithm that is hoped to learn a good prediction function when presented with a new (unseen) task $\mathcal{T}=(p,\mathcal{L})$ sampled from the same meta distribution $\mathcal{P}$. The meta-training phase consists in computing the learning algorithm $A=ML(\mathcal{D}^{1},\dots,\mathcal{D}^{N})$ where $\mathcal{D}^{j}\sim p^{j}$ and $ML$ is the meta-learning algorithm. The meta-test phase, in turn, consists in computing the prediction function using the new task: $h=A(\mathcal{D})$, with $\mathcal{D}\sim p$.
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multi-task learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several “source” tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\leavevmode\nobreak\ \sim p$ will be available for the new task (few-shot learning) or even a single example per class (one-shot learning). The main insight is that when the tasks are related, as captured by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit]. Note that while HPO tipycally involves the optimization of a “black-box” function, both the meta-training and the meta-test phases in meta-learning more usually involve the optimization of functions with a know structure, enabling the application of gradient-based techniques, as discussed below.
8.4.2 Meta-learning of common representations
One effective approach to meta-learning consists of using past tasks to build strong representations, which can be later used for few-shot learning on the new task. We briefly discuss this approach here, as it is the most closely related to the HPO problem [franceschi2018bilevel]. As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}:\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}:\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}:\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathbb{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multi-task network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define
$f(\lambda)=\sum_{j=1}^{N}\mathcal{L}^{j}\left(w^{j},\lambda,\mathcal{D}^{j}_{\text{val}}\right).$
Figure 8.1: Tasks and datasets for meta-learning on Omniglot [lake:2019:omniglot_challenge_3year]. At the end of the meta-training phase, a small training set for the the new task $\mathcal{T}$ is provided and performance is evaluated on test data $\mathcal{D}_{\text{test}}$. A similar figure was shown in [ravi_optimization_2016] for tasks derived from Imagenet.
We are now ready to formulate the following bilevel program
$\min_{\lambda}$ $f(\lambda)$ (8.3)
s.t. $w^{j}\in\text{arg}\min_{u}\mathcal{L}^{j}\left(w^{j},\lambda,\mathcal{D}^{j}_{\text{tr}}\right)$ (8.4)
where the lower level (or inner) objectives $\mathcal{L}^{j}$ are empirical risk minimization on each training tasks, and the upper (or outer) objective is the validation error on the same tasks. If $\Phi$ and $\gamma$ are realized by deep learning modules, the inner problem would be solved by some iterative gradient-based algorithm, implying that the solution would only be found approximately. Moreover, in general, the above problem is not guaranteed to have a solution. In [franceschi2018bilevel] it is shown that a solution is guaranteed to exist if $\gamma$ is linear (i.e., the classification head consists of a single dense layer) and an algorithm based on hypergradients (using the method described in  Section 7.3) is presented to solve the meta-learning problem. A closely related algorithm is proposed in [finn_model-agnostic_2017] although it is not explicitly formulated as a bilevel program and the emphasis is on fine-tuning to the target task an architecture pretrained on the sequence of previous tasks. A solution based on implicit gradients (see Section 7.2) is offered in [rajeswaran2019meta].
Lastly, while we presented meta-learning as shared feature learning, we note that other tasks’ sharing knowledge mechanisms are possible. Notable examples are biased regularization [denevi2018learning, denevi2019learning, balcan2019provable], in which the meta-parameter $\lambda$ describes a common mean among the tasks, hyper-network approaches, in which $\lambda$ is associated to a function that outputs the parameters of a task-specific models [zhao2020meta, zakerinia2024more], and conditional meta-learning [denevi2020advantage, denevi2022conditional, rusu2018meta, wang2020structured], in which $\lambda$ is a conditioning function of side information specific to the task at hand.
8.5 Transfer of hyperparameters across model scale
A foundation model [brown2020language], such as a large language model, is a large-scale, pre-trained machine learning model, typically based on deep learning architectures like transformers [vaswani_attention_2017], that serves as a general-purpose backbone for various downstream tasks through fine-tuning or prompting. These models mark a cornerstone in artificial intelligence and have led to several breakthroughs in code generation, natural language processing, and common-sense reasoning. However, the expensive pre-training phase requires vast amounts of data and tremendous GPU resources.
While this also depends on a careful selection of hyperparameters, the hyperparameter optimization methods presented here are not directly applicable. Even efficient state-of-the-art multi-fidelity approaches (see Chapter 5) require at least a few training runs on the full budget, but for contemporary LLMs, this is infeasible.
Inspired by a theoretical perspective from tensor programs, a recent line of work [yang-arxiv22, everett-icml24a] approaches this problem from a different angle. While hyperparameters of neural networks, such as the learning rate, change when we increase the width or depth of the network and need to be adjusted accordingly, the idea here is to change the parameterization of neural networks, such that the optimal hyperparameters stay the same when scaling from smaller models to larger models. This allows for finding the optimal hyperparameters on a small scale and directly transferring them to the largest scale.
The original parameterization proposed by [yang-arxiv22], called $\mu$-parameterization, initially provided only a theoretical foundation for scaling model width, however they also provided empirical results confirming its effectiveness for depth scaling. Recent work by [everett-icml24a] has shown that even hyperparameters such as the learning rate can be transferred across different model widths using the standard parameterization of neural networks by adapting them on a per-layer basis.
While this line of research has shown promising results, it has so far been mostly theoretical and primarily focused on width scaling. Nevertheless, it has the potential to lead to more practical hyperparameter optimization methods that scale efficiently to large foundation models.
Chapter 9 Hyperparameter Optimization Systems
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO software landscape changes frequently as new problems, methods and application domain become relevant: the provided list here is by no means exhaustive and only represents a snapshot.
9.1 Open Source Libraries and Frameworks
The Python package Spearmint [snoek_practical_2012] is one of the first libraries that provides Bayesian optimization for hyperparameter optimization. It implements Bayesian optimization using Gaussian processes (see Section 4.2). At the same time, bergstra2011algorithms provided HyperOpt, offering Bayesian hyperparameter optimization by using the tree Parzen estimator (Section 4.2.5) instead of Gaussian processes (Section 4.2.2). Hutter2011:Sequential introduced SMAC, which implemented Bayesian optimization strategies using Random Forests. Initially targeting at more generic algorithm configuration problems, such as configuring SAT solvers, it also offered functionality for HPO of machine learning methods. The current version [smac3] is written in Python and also provides more advanced functionalities, such as multi-fidelity optimization (see Chapter 5).
Among the first Python libraries that implemented multi-fidelity HPO, such as Fabolas [klein2017fast] or Hyperband [li2016hyperband], was the RoBO library by klein-bayesopt17 and Dragonfly by kandasamy-jmlr19.
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on the large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distribute the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
Recent PyTorch implementations of gradient-based hyperparameter optimization algorithms include the libraries HyperTorch [grazzi2020iteration], higher [grefenstette2019generalized] and Betty [choe2023betty]. JaxOpt [blondel2022efficient] is a Jax implementation that focuses on implicit differentiation.
Most HPO libraries or frameworks expect the objective function either as function handle, for example in Python, or as a single training script. The search space is usually defined outside of the actual training script and hyperparameters are passed as input arguments. Optuna [akiba-sigkdd19] follows a different approach by a so-called define-by-run API that allows users to dynamically define the search space instead of providing it in advance.
9.2 Commercial Systems
Since hyperparameter optimization plays such an integral role in the machine learning workflow, most machine learning platforms, for example Weights and Biases, provide some functionality to select hyperparameters. There are also several commercial systems specific for automated hyperparameter optimization. For example, Google Vizier [google_vizier] is an internal service that contributed to several Google products and research efforts. oss_vizier also introduces a open-source version of Google Vizier as a stand-alone Python package. AWS SageMaker provides a fully automated service called Automated Model Tuner [perrone2021sagemaker] for gradient-free optimization problems, which allows for large-scale HPO leveraging built-in SageMaker functionalities. Another example is Sigopt, which provides optimization-as-a-service for hyperparameter optimization utilizing Bayesian optimization techniques.
9.3 Benchmarking
A key challenge for an empirical driven research field such as HPO to make progress is the thorough evaluation of new approaches and comparison to baseline methods. However, due to the expensive nature of the underlying optimization problem, it is often prohibitively expensive to obtain statistically reliable results including multiple repetitions, which are necessary to characterize HPO algorithms that often come with intrinsic randomness.
To reduce the computational requirements and enable large-scale experimentation, recent work tries to emulate expensive HPO problems by using cheap-to-evaluate approximations of the underlying objective function. eggensperger-aaai15 introduced the idea of surrogate benchmarks that replace the original benchmark with the predictions of a machine learning model. This model is first trained in an offline step on data collected from the original benchmark. Unlike the probabilistic model used in Bayesian optimization, where data is generated iteratively, all data in this approach is generated simultaneously, and the model is trained only once. While using a machine learning model to approximate the objective function introduces potentially some bias, single function evaluations are now orders of magnitude cheaper, since they only require a single prediction.
Instead of using a machine learning model to approximate the true objective, klein-arxiv19a first discretzied the search space of a fully-connected neural network leading to a finite number of hyperparameter configurations and then performed an exhaustive search to collect the performance and runtime of each configuration. This allows to replace the actual training and validation of a hyperparameter configuration by simple table look-ups during the HPO process. To mimic the observation noise that comes with the HPO of neural networks, each HPO configuration was evaluated multiple times. During the HPO process, random entries for each hyperparameter configuration are returned to simulate the observation noise. The idea of tabularized benchmarks has since been extended to neural architecture search [ying-icml19, dong-iclr20] (see Section 8.3) leading to several new benchmarks. While tabularized benchmarks return for each query truly observed values of the objective function instead of prediction of the model, pfisterer-automl22 argued that the discretization of the search space eliminates the local structure of the objective function, leading actually to a higher bias than using surrogate models. The idea of surrogate benchmarks has been extended to define a generative model by [klein-nips19], which allows sampling an arbitrary set of new benchmarks.
To standardize the interface to different benchmarks, eggensperger_towards_2013 introduced HPOLib, which consolidates different benchmarks from the literature. In follow up work [eggensperger-neurips21], HPOBench containerized these benchmarks to increase reproducibility of benchmarking results. HPO-B [pineda-neurips21] provides tablularized benchmarks based on OpenML datasets [vanschoren-sigkdd13a]. In a similar vein, zimmer-ieee21 introduced LCBench, which provides full learning curves of neural network models trained on OpenML datasets. salinas-automl22 also provided a unified interface to tabular and surrogate benchmarks in Syne-Tune, using a fast data format to accelerate the table look-ups. Another suite for surrogate-based benchmarking is YAHPO [pfisterer-automl22], consisting of different scenarios, including combined hyperparameter optimization and algorithm selection.
Chapter 10 Research Directions
From some perspective, hyperparameter optimization may be considered by now a mature field with well-established algorithmic solutions and with industry-level software frameworks. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (random and quasi-random search, model-based, multi-fidelity, population-based, and gradient-based) and highlighted connections to closely related areas such neural architecture search and meta-learning.
Yet, in all classes of algorithms there remain open issues and interesting directions for improvement. In addition, the development of hybrid algorithms has so far been under-explored, but could potentially bring notable benefits in practical applications. We do not discuss these particular aspects here, but rather touch upon some general topics which we feel may become (or rather, are already becoming) broad and important research directions in HPO.
10.1 Response functions for unsupervised learning
HPO requires a well defined response function to be minimized (see Section 2.2), which in turn requires a well defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning.
A notable example is anomaly detection, an area rich of methods, results, and applications [pang:2021:deep_learning_anomaly]. The anomaly or novelty detection problem essentially consists of identifying the high probability region of the distribution behind the data, so that points outside this region can be detected as outliers. For this problem, [ma:2023:need_unsupervised_outlier] survey a number of internal (i.e., not using any ground truth labels) approaches for assessing outlier detection models and conclude that “none of the existing and adapted strategies would be practically useful.”
A second important example is generative models. In this case the goal is to construct a model of the data distribution such that sampling from that model produces fake data points that are indistinguishable from real data. Again, the literature is rich in terms of models, ranging from early generative adversarial networks [DBLP:journals/corr/GoodfellowPMXWOCB14] to normalizing flows [pmlr-v37-rezende15] and more recently probabilistic diffusion models [ho:2020:denoising_diffusion_probabilistic]. In the particular case of images, popular metrics used to compare different generative models are the Inception Score [NIPS2016_8a3363ab] and the Fréchet inception distance (FID) [DBLP:conf/nips/HeuselRUNH17]. These metrics only capture one aspect of the generative process, namely fidelity. However, the problem of assessing the quality of a model is inherently more complicated: it is not sufficient to generate faithful points, we also need to cover all regions where the data distribution is high.
There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports (of real and fake data) are estimated and compared (fake points off the support of real distribution are interpreted as false positives, and real points outside the support of the model’s distribution are interpreted as false negatives). The idea has sparkled a number of practically computable metrics [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to guide HPO. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions requires reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionalty issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced in [yang:2020:evaluation_generative_models] for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data. A closely related metric based on maximum mean discrepancy [obray:2021:evaluation_metrics_graph], popular in the context of graph generative models, has been shown to be complicated due to its strong sensitivity to hyperparameters in the definition of the metric itself (leading to a sort of hyper-hyperparameters).
A notable exception in the realm of unsupervised learning is the case of obtaining disentangled representations, which is the problem of identifying a small number of explanatory factors behind observed data [bengio:2013:representation_learning_reviewa, locatello:2019:challenging_common_assumptions], such as shape, color and position for image classification problems. For this particular case, evaluation metrics have been developed based on informativeness, separability and inference, and interpretability [do:2019:theory_evaluation_metrics]. For this particular problem, a hyperparameter tuning algorithm was developed in [duan:2019:unsupervised_model_selection] using ranking scores. Still, the development of general purpose HPO algorithms for unsupervised learning remains an open problem.
10.2 Learning to (hyperparameter) optimize
In section 8.4 we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning an hyperparamter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and response functions of one or several standard learning algorithms. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g. validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on data collected while optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modelling and learning to optimize, a sub-field of meta-learning, [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
The usage of rich meta-data, past experience, and the capacity to leverage similarities across standard learning algorithms may hold the promise of discovering HPO algorithm that are more efficient than classic approaches. However, significant challenges remain in developing truly general and robust meta-learning for HPO, including handling the diversity of problem types, scalability to high-dimensional spaces, and transferability across vastly different domains and algorithms.
10.3 HPO and foundation models
The emergence of foundation models [bommasani2021opportunities] as a new frontier for machine learning applications and research opens up a series of novel scenarios and challenges for hyperparameter optimization.
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become standard paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices. These choices are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparamter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (cf. Chapter 7) that optimizes for discrete “trigger” tokens, Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; cf. Section 5.2.1); HbBoPs [schneider2024hyperband], which leverages a structure-aware Gaussian Process surrogate combined with Hyperband to efficiently select prompts in a black-box LLM setting; Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches (with discrete gradient estimators – see Section 7.6) may prove useful in these directions,
In addition, the emergence of agents and agent networks where various LLMs interact and act by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs queried through APIs. These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the the communication protocols and the interactions graph between agents – see e.g. GPTSwarm [zhuge2024language]. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.
Acknowledgements.
The work of PF was partially supported by Italian national project CAI4DSA, CUP B13C23005640006.
[…]